### v3

# Jaime Angulo Pava  and  Alexander Muñoz

Abstract
The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a graph consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex. Characterizations of extensions generating unitary and contractile dynamics are provided for the Airy operator in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex. Employing the same abstract techniques used in the former case, we characterize the self-adjoint extensions of the Schrödinger operator on $\mathcal{T}$-shaped graphs, i.e., to graphs that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$. The theory established in this manuscript can be considered as a first step towards a better understanding of the dynamics for the remarkable dispersive models NLS and KdV on metric graphs.
Mathematics Subject Classification (2020). Primary 35Q51, 35Q55, 81Q35, 35R02; Secondary 47E05.
Key words. Schrödinger equation, Airy equation, quantum graphs, Krein spaces, extension theory of symmetric operators.
1. Introduction
Nonlinear evolution models on metric graphs arise as models in wave propagation, for instance, in quasi-one-dimensional systems (e.g., meso- or nanoscale) that resemble a thin neighborhood of a graph. These models provide a convenient means to study various physical effects in the real world, both from theoretical and experimental perspectives. The flexibility in setting the geometry of the graph configuration allows for the creation of different dynamics. To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [17, 19, 20, 24, 25, 27, 30, 35, 37, 40, 45] and reference therein).
Recently, nonlinear models on graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries model, have been studied extensively (see [2, 11, 12, 8, 9, 13, 14, 15, 3, 30, 40] and references therein). From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the ”topology of the graph”), which governs the dynamics on the network via groups. This area of study has attracted significant attention, particularly in the context of soliton transport. Solitons and other nonlinear waves in branched systems provide valuable insights into the dynamics of these models.
Addressing these issues is challenging because both the equations of motion and the graph topology can be complex. Additionally, a critical aspect that complicates the analysis is the presence of one or more vertices where a soliton profile, arriving at the vertex along one of the edges, undergoes complicated behaviors such as reflection and the emergence of radiation. This makes it difficult to observe how energy travels across the network. As a result, the study of soliton propagation through networks presents significant challenges. The mechanisms for the existence and stability (or instability) of soliton profiles remain unclear for many types of graphs and models.
We recall that a metric graph $\mathcal{G}$ is a structure represented by a finite number of vertices $V=\{\nu_{i}\}$ and a set of adjacent edges at the vertices $E=\{e_{j}\}$ (for further details, see [17]). Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$. Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical ”wires” or ”networks” connecting them. The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$. This identification introduces the coordinate $x_{e}$ along the edge $e$.
An example of a metric graph, which will be the focus of our study here, is a looping-edge graph. This is a graph composed of a circle and a finite number $N$ of infinite half-lines attached to a common vertex $\nu=L$. If the circle is identified with the interval $[-L,L]$ and the half-lines with $[L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1). We use the subscripts 00 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively. The special case $N=1$ is known as a tadpole graph (see also Figure 1).
$-L$
$L$
$-L$
$L$
Tadpole graph
Looping-edge graph
Figure 1. Tadpole and looping-edge-type graphs
A wave function U defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\textbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
Consider the case of nonlinear Schrödinger-type equations (NLS)
$i\textbf{U}_{t}\ +\Delta\textbf{U}+(p+1)|\textbf{U}|^{2p}\textbf{U}=\textbf{0},\quad p>0,$ (1.1)
where the action of the Laplacian operator $\Delta$ on a general graph $\mathbb{G}$ is given by
$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
The NLS in (1.1) has been extensively studied in recent literature for various types of metric graphs $\mathbb{G}$ and with specific domains for the Laplacian, which make it a self-adjoint operator on $L^{2}(\mathbb{G})$ (see the review manuscript [34]). For instance, research has focused on star graphs ([2, 1, 11, 12, 32, 40] and references therein), looping-edge graphs ([6, 7, 21, 42, 41] and references therein), as well as flower graphs, dumbbell graphs, double-bridge graphs, and periodic ring graphs ([21, 18, 33, 31, 36, 41, 43] and references therein).
One of the objectives of this work is to obtain a characterization of all self-adjoint extensions of the symmetric operator $\mathcal{H}_{0}\equiv-\Delta$ on the metric graph $\mathcal{G}$ with domain $D(\mathcal{H}_{0})$ defined by
$D(\mathcal{H}_{0})=C_{0}^{\infty}(-L,L)\oplus\bigoplus_{j=1}^{N}C_{0}^{\infty}(L,+\infty).$ (1.3)
Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $-\Delta_{ext}$ is given by $-\Delta_{ext}=-\Delta$. Each member of this family defines a unitary dynamic for the linear evolution problem
$\left\{\begin{array}[]{ll}i\textbf{U}_{t}=-\Delta_{ext}\textbf{U}\\ \textbf{U}(0)=\textbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right.$ (1.4)
We point out that our approach, based on Krein spaces [44], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point $-L$ as a terminal vertex through a relatively simple matrix process (see Theorems 5.2-5.3 below). At this level, our approach can be seen as a complement to the works of Exner et al. [27, 28, 29], which, to our knowledge, has not formally appeared in the literature on $\mathcal{G}$. Furthermore, in general, the extensions $(-\Delta_{ext},-D(\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of $[-L,L]$. (see, for instance, Example 8 associated to $\delta^{\prime}$- type interactions on looping-edge graphs and Example 10 below by considering $-L$ as a terminal vertex). Moreover, our approach to obtaining these extensions will be based on those associated with the Airy operator (1.7) which is the main focus of our study. Both extensions theories to be established can be seen as a first step for future studies on the existence and stability of soliton profiles for both NLS-(1.1) and KdV-(1.8) models, for instance, on a looping-edge graph or on a $\mathcal{T}$-shaped graph, i.e., a graph that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$ (see Figure 4). In the case of looping-edge graphs, the extensions must ensure that the wave functions are continuous at least on the loop (i.e., $\phi(L)=\phi(-L)$). For example, in the study of electron motion in thin metallic ”wires,” it is required that the wave function remains continuous at the junction $\nu=L$. These boundary conditions have a clear physical interpretation, representing the conservation of probability flow at the junction (see [28]).
This physical requirement can be formally expressed through the following family of self-adjoint operator extensions $(-\Delta,D_{Z,N})_{Z\in\mathbb{R}}$ for $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ with
$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdot\cdot\cdot=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
where for any $n\geqq 0$,
$H^{n}(\mathcal{G})=H^{n}(-L,L)\oplus\bigoplus_{j=1}^{N}H^{n}(L,+\infty).$
The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann-Kirchhoff type if $Z=0$. The parameter $Z$ is a coupling constant between the loop and the several half-lines (see Example 7 and Remark 1). The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [27], [29], and references therein). It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 6, 7, 16, 21, 22, 34, 42, 41], and references therein).
Our second focus in this manuscript is associated to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t),\quad x\neq 0,t\in\mathbb{R},\textbf{e}\in E,$ (1.6)
with real coefficients given by the sequences $(\alpha_{\textbf{e}})_{e\in E}$, $(\beta_{\textbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping edge graph (see Theorems 4.2, 4.3 and 4.7). Our approach is based on the abstract results in [44] applied to the case of the Airy-operator
$A_{0}:\{u_{e}\}_{e\in E}\mapsto\{\alpha_{e}\partial_{x}^{3}u_{e}+\beta_{e}\partial_{x}u_{e}\}_{e\in E},$ (1.7)
when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and treating the point $-L$ as a terminal vertex (as in the case of a $\mathcal{T}$-graph). Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary. Thus, we get in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. For instance, in Example 1 below we get $\delta$-interactions type boundary condictions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]) under conditions on $\alpha_{\textbf{e}}$ and $\beta_{\textbf{e}}$. Also, we establish conditions for obtaining extensions of $A_{0}$ generating contractile dynamics.
We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg-de Vries (KdV) equation
$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t)+2u_{\textbf{e}}(x,t)\partial_{x}u_{\textbf{e}}(x,t),$ (1.8)
on the metric graph $\mathcal{G}$. Recently, Mugnolo, Noja, and Seifert [38] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively. The half-lines are connected at a unique vertex $\nu=0$. In this case the graph is sometimes called a star-shaped metric graph. Several arguments used in [44] and [38] in star-shaped graphs are extended in this work for looping-edge graphs. We recall that when the star-shaped metric graph is balanced ($|{E}_{+}|=|{E}_{-}|$), the extension theory implies that there is a $9|{E}_{+}|^{2}$-parameter family of skew-self-adjoint operators for $A_{0}$. Now, on looping-edge graph when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we get that there must be $9(k+1)^{2}$ family of unitary groups, each one having its own representation.
As far as we know, the study of the KdV equation on metric graphs remains relatively underdeveloped. Cavalcante [23] studied the local well-posedness of the Cauchy problem associated with (1.8) in Sobolev spaces $H^{s}(\mathcal{Y})$ with low regularity on a $\mathcal{Y}$-junction graph, $\mathcal{Y}=(-\infty,0)\cup(0,+\infty)\cup(0,+\infty)$. Subsequently, Angulo and Cavalcante [8] established a criterion for the linear instability of stationary solutions for the KdV on arbitrary star-shaped metric graphs. This criterion was applied to tail and bump profiles on balanced graphs with specific boundary conditions at the vertex $\nu=0$, ensuring they remain linearly unstable (see Theorem 6.2 in [8]). More recently, Mugnolo, Noja, and Seifert [39] investigated the existence of solitary waves on a star-shaped metric graph in terms of the coefficients of the KdV equation $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$ on each edge, with coupling conditions at the vertex $\nu=0$, which may or may not involve continuity, and examined the speeds of the traveling waves. Lastly, Angulo and Cavalcante [10] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$- interaction type. They also extended the linear instability result from [8] to establish nonlinear instability.
Results of the dynamics of the KdV model on looping edge graphs or on $\mathcal{T}$-shaped graphs, such as local well-posedness of the Cauchy problem, existence and stability of solitary waves, etc., are not known in the literature to the best of the authors knowledge. We believe that the theory established in this manuscript can be considered as a first step towards a better understanding of this notable dispersive model on metric graphs. Organization of the article.
  1. 2.
    Preliminaries
    1. 2.1.
      Notation
    2. 2.2.
      Stone and Lummer-Phillips Theorems
    3. 2.3.
      Operators on Krein Spaces
  2. 3.
    The free Airy operator and deficiency indices
  3. 4.
    The Airy operator on a looping-edge graph
    1. 4.1.
      Boundary system for even amount of half-lines
      1. 4.1.1.
        Extensions generating unitary dynamics
      2. 4.1.2.
        Examples: $\delta$-type interactions on a looping-edge graph
    2. 4.2.
      Boundary system for arbitrary amount of half-lines
      1. 4.2.1
        Extensions generating contractile dynamics
      2. 4.2.2
        Examples: $\delta$-type interactions on a tadpole graph
    3. 4.3.
      Boundary system separating derivatives
      1. 4.3.1
        Extensions generating contractile dynamics via splitting derivatives
      2. 4.3.2
        Examples
  4. 5.
    The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped metric graphs.
    1. 5.1.
      Deficiency indices for the free Schrödinger operator
    2. 5.2.
      Boundary system
    3. 5.3.
      Extensions generating unitary dynamics
    4. 5.4.
      Extensions generating unitary dynamics parameterized by linear subspaces
    5. 5.5.
      Extensions to $\mathcal{T}$-shaped metric graphs
2. Preliminaries
2.1. Notation
Let us introduce some notation related to the boundary values of a function posed in $\mathcal{G}$. For $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$, we denote the vector of values of U and its derivatives at the vertex with
$\vec{\textbf{U}}:=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\ \vec{\textbf{U}}^{\prime}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mbox{ and }\vec{\textbf{U}}^{\prime\prime}:=\begin{pmatrix}\phi^{\prime\prime}(-L)\\ \phi^{\prime\prime}(L)\\ \psi_{1}^{\prime\prime}(L)\\ \vdots\\ \psi_{N}^{\prime\prime}(L)\end{pmatrix}.$
For the components of U, we denote the derivatives values at the vertex with
$\partial\phi(-L):=\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\end{pmatrix},\ \partial\phi(L):=\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}\mbox{ and }\partial\psi_{j}(L)=\begin{pmatrix}\psi_{j}(L)\\ \psi_{j}^{\prime}(L)\\ \psi_{j}^{\prime\prime}(L)\end{pmatrix}.$
We adopt concatenation of vectors in vertical stack. For instance,
$\begin{pmatrix}\partial\phi(-L)\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\\ \phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}.$
For convenience we denote for $\psi=(\psi_{1},\dots,\psi_{N})$
$\partial\psi(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{2}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix},\ \partial\psi_{even}(L):=\begin{pmatrix}\partial\psi_{2}(L)\\ \partial\psi_{4}(L)\\ \vdots\\ \partial\psi_{2k}(L)\end{pmatrix}\mbox{ and }\partial\psi_{odd}(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{3}(L)\\ \vdots\\ \partial\psi_{2k-1}(L)\end{pmatrix}.$ (2.1)
2.2. Stone and Lumer-Phillips Theorems
In this subsections we recall some well-known features of group theory.
Let $X$ be a Banack space and $X^{\prime}$ its dual space. Consider $\mathcal{J}(x):=\{x^{\prime}\in X^{\prime}\mid\langle x,x^{\prime}\rangle=\|x\|^{2}=\|x^{\prime}\|^{2}\}$. Note $\mathcal{J}(x)$ is non-empty by Hahn-Banach Theorem. Moreover, if $X$ is a Hilbert space then $\mathcal{J}(x)$ consists of a single element.
Recall that a linear operator $(A,D(A))$ on a Banach space $X$ is called dissipative if $\|(\lambda-A)x\|\geqslant\lambda\|x\|$.
Proposition 2.1.
An operator $(A,D(A))$ is dissipative if and only if for every $x\in D(A)$ there exists $j(x)\in\mathcal{J}(x)$ such that
$\mbox{Re}\langle Ax,j(x)\rangle\leqslant 0.$ (2.2)
If $A$ is the generator of a strongly continuous contraction semigroup, then (2.2) holds for all $x\in D(A)$ and arbitrary $x^{\prime}\in\mathcal{J}(x)$.
Proof.
See Proposition 3.23 in [26]
Note in the case $X$ is a Hilbert space, condition (2.2) reduces to $\mbox{Re}(Ax\mid x)\leqslant 0$.
Theorem 2.2.
(Lummer-Phillips) Let $(A,D(A))$ be a densely defined operator on a Banach space $X$. If both, $A$ and $A^{*}$ are dissipative, then the closure $\overline{A}$ of $A$ generates a contraction semigroup on $X$.
Proof.
See Corollary 3.17 in [26]. ∎
We also recall the Stone’s theorem for unitary groups.
Theorem 2.3 (Stone).
Let $(A,D(A))$ be a densely defined operator on a Hilbert space $H$. Then $A$ generates a unitary group $(T(t))_{t\in\mathbb{R}}$ on $H$ if and only if $A$ is skew-self-adjoint, i.e., $A^{*}=-A$ and $D(A^{*})=D(A)$.
2.3. Operators on Krein Spaces
Following [44] we introduce the abstract notions of self-orthogonal subspaces and operators based on indefinite inner product spaces.
Definition 2.1.
Let $Y$ be a linear space, let $X\subset Y$ be a subspace and $w$ a sesquilinear form defined on $(Y\times Y)^{2}$. The space $X$ is called $w$-self-orthogonal if
$X=X^{\perp_{w}}\equiv\{(x,y)\in Y\times Y\mid w((x,y),(\tilde{x},\tilde{y}))=0\mbox{ for all }(\tilde{x},\tilde{y})\in X\}.$ (2.3)
Definition 2.2.
A Krein space $(\mathcal{K},(\cdot\mid\cdot)_{\mathcal{K}})$ is a complex linear space $\mathcal{K}$ on which a indefinite inner product $(\cdot\mid\cdot)_{\mathcal{K}}$ is defined, namely, the function $(x,y)\in\mathcal{K}\times\mathcal{K}\to(x\mid y)_{\mathcal{K}}\in\mathbb{C}$ satisfies the following axioms:
  1. i)
    Linearity in the first argument:
    $(\alpha x_{1}+\beta x_{2}\mid y)_{\mathcal{K}}=\alpha(x_{1}\mid y)_{\mathcal{K}}+\beta(x_{2}\mid y)_{\mathcal{K}}$
    for all $x_{1},x_{2},y\in\mathcal{K}$ and all complex numbers $\alpha,\beta$;
  2. ii)
    antisymmetry:
    $(x\mid y)_{\mathcal{K}}=\overline{(y\mid x)_{\mathcal{K}}}$
    for all $x,y\in\mathcal{K}$;
  3. iii)
    nondegeneracy: if $(x\mid y)_{\mathcal{K}}=0$ for all $y\in\mathcal{K}$, then $x=0$.
Thus, the function $(\cdot\mid\cdot)_{\mathcal{K}}$ satisfies all the properties of a standard inner product with the possible exception that $(x\mid x)_{\mathcal{K}}$ may be nonpositive for $x\neq 0$.
We recall that if $dim(\mathcal{K})=n$ ($\mathcal{K}\cong\mathbb{C}^{n}$) every indefinite inner product $(\cdot\mid\cdot)_{\mathcal{K}}$ on $\mathcal{K}$ is determined by an $n\times n$ invertible and hermitian matrix $H$ such that $(x\mid y)_{\mathcal{K}}=\langle Hx,y\rangle_{\mathbb{C}^{n}}$ holds for all $x,y\in\mathcal{K}$.
Definition 2.3.
Let $\mathcal{K}_{1},\mathcal{K}_{2}$ be Krein spaces and a linear operator $\mathcal{L}:\mathcal{K}_{1}\to\mathcal{K}_{2}$.
  1. (1)
    $\mathcal{L}$ is called a $(\mathcal{K}_{1},\mathcal{K}_{2})$-contraction if
    $\left(\mathcal{L}x\mid\mathcal{L}x\right)_{\mathcal{K}_{2}}\leqslant\left(x\mid x\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L}).$ (2.4)
  2. (2)
    If $\mathcal{L}$ is densely defined, the $(\mathcal{K}_{2},\mathcal{K}_{1})$-adjoint of $\mathcal{L}$ is defined by
    $D(\mathcal{L}^{\#}):=\{y\in\mathcal{K}_{2}\mid\exists z\in\mathcal{K}_{1}:\left(\mathcal{L}x\mid y\right)_{\mathcal{K}_{2}}=\left(x\mid z\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L})\}\mbox{ and }\mathcal{L}^{\#}y=z.$ (2.5)
  3. (3)
    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and #$\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
3. The free Airy operator and deficiency indices
In this section we fix notation and specify the functional spaces in which the Airy operator will be studied. We also provide an analysis on how the deficiency indices for the free Airy operator behave.
We consider the Airy operator
$A_{0}:\{u_{e}\}_{e\in E}\mapsto\{\alpha_{e}\partial_{x}^{3}u_{e}+\beta_{e}\partial_{x}u_{e}\}_{e\in E},$ (3.1)
where $\alpha_{j}$ and $\beta_{j}$ are real numbers for any $j$. For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
$A_{0}\textbf{U}=\left(\alpha_{0}\partial_{x}^{3}\phi+\beta_{0}\partial_{x}\phi,\ \alpha_{1}\partial_{x}^{3}\psi_{1}+\beta_{1}\partial_{x}\psi_{1},\ \dots\ ,\ \alpha_{N}\partial_{x}^{3}\psi_{N}+\beta_{N}\partial_{x}\psi_{N}\right).$ (3.2)
We consider $A_{0}$ as a densely defined operator in
$L^{2}(\mathcal{G}):=L^{2}(-L,L)\oplus\bigoplus_{j=1}^{N}L^{2}(L,\infty)$ (3.3)
with domain
$D(A_{0})=C_{0}^{\infty}(-L,L)\oplus\bigoplus_{j=1}^{N}C_{0}^{\infty}(L,\infty).$ (3.4)
For $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$, we denote the inner product on $L^{2}(\mathcal{G})$ with
[…]
We consider analogously the Sobolev spaces
$H^{s}(\mathcal{G}):=H^{s}(-L,L)\oplus\bigoplus_{j=1}^{N}H^{s}(L,\infty).$
Note in particular $D(A_{0}^{*})=H^{3}(\mathcal{G})$, where $A_{0}^{*}=-A_{0}$ is the adjoint of the minimal operator defined as the closure of $A_{0}$ with domain $D(A_{0})$ in (3.4).
For the densely defined symmetric operator $iA_{0}$ we consider the deficiency spaces
$\mathcal{D}_{\mp}(iA_{0}^{*}):=\mbox{Ker}(iA_{0}^{*}\pm iI),$ (3.6)
with deficiency indices $d_{\mp}=\mbox{dim}\mathcal{D}_{\mp}(iA_{0}^{*})$.
From the definition of $A_{0}$ in (3.1) we have that solutions of $-i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
$u_{e}=c_{1}e^{r_{1}x_{e}}+c_{2}e^{r_{2}x_{e}}+c_{3}e^{r_{3}x_{e}}$
where $r_{1},r_{2}$ and $r_{3}$ are the roots of $p(x)=\alpha_{e}x^{3}+\beta_{e}x\mp 1$. We refer the reader to the proof of Lemma 2.1 in [38] for an explicit description.
We note that $e^{rx}$ belong to $L^{2}(-L,L)$ independent on $r$, which means $d_{-_{0}}=d_{+_{0}}=3$ along $e_{0}$. On the other hand, on a half-line $[L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$ and $\beta_{e}$. We summarize the behaviour (on a single edge $[L,\infty)$) in the following table.
Signs of $\alpha_{e}$ and $\beta_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
$\alpha_{e}>0,\beta_{e}\geqslant 0$ 2 1
$\alpha_{e}>0,\beta_{e}<0$ 2 1
$\alpha_{e}<0,\beta_{e}\geqslant 0$ 1 2
$\alpha_{e}<0,\beta_{e}<0$ 1 2
From the Von Neumann and Krein theory, one can expect self-adjoint extensions of $iA_{0}$ only if the deficiency indices agree, that is, $d_{+}=d_{-}$. In such case, according to Theorem 2.3, there exist extensions that generate unitary dynamics. Several combinations of signs of $\alpha_{e}$ and $\beta_{e}$ along the edges of $\mathcal{G}$ may lead to equal deficiency indices. For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $33k+$ 3. Thus, in this case, we know that there must be a $9(k+1)^{2}$ family of unitary groups for the case of the looping-edge graph, each one having its own representation.
4. The Airy operator on a looping-edge graph
As mentioned in the previous section, unitary dynamics are expected only in particular cases where the deficiency indices match for the free Airy operator. In the following subsections, we will introduce different approaches on a looping-edge graph $\mathcal{G}$ based on whether the deficiency indices coincide or not. We start by considering a ”pair-wise” reference frame on $\mathcal{G}$ that generates unitary dynamics, (see Theorem 4.3 below).
4.1. Boundary system for even amount of half-lines
Let us first consider $\mathcal{G}$ consisting of an even amount $2k$ of half-lines attached to the vertex in which exactly half of them have a leading coefficient $\alpha_{e}>0$. More precisely, we consider $\mathcal{G}$ defined trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if ±$\pm\alpha_{e}>0$.
Our objective in this section is to provide a first description of the boundary values of $A_{0}^{*}=-A_{0}$ ($D(A_{0}^{*})=H^{3}(\mathcal{G})$), we proceed as in Section 2 of [44]. For an alternative description see also Lemma 4.8 in Section 4.3.
Definition 4.1.
We define in the graph of $A_{0}^{*}=-A_{0}$,
$G(A_{0}^{*})=\{(\textbf{U},A_{0}^{*}\textbf{U})\mid\textbf{U}\in D(A_{0}^{*})\}$
the standard anti-symmetric form $\Omega:G(A_{0}^{*})\times G(A_{0}^{*})\to\mathbb{C},$ defined by
[…]
with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
Proposition 4.1.
For any $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$ we have
$\begin{split}&[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\\ &\hskip 14.22636pt\left(\begin{pmatrix}\beta_{0}&0&\alpha_{0}\\ 0&-\alpha_{0}&0\\ \alpha_{0}&0&0\end{pmatrix}\begin{pmatrix}\phi(-L)\\ \phi^{\prime}(-L)\\ \phi^{\prime\prime}(-L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}(-L)\\ \tilde{\phi}^{\prime}(-L)\\ \tilde{\phi}^{\prime\prime}(-L)\end{pmatrix}\right)-\left(\begin{pmatrix}\beta_{0}&0&\alpha_{0}\\ 0&-\alpha_{0}&0\\ \alpha_{0}&0&0\end{pmatrix}\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\\ \phi^{\prime\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}(L)\\ \tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime\prime}(L)\end{pmatrix}\right)\\ &\hskip 14.22636pt+\sum_{j=1}^{N}\left(\begin{pmatrix}\beta_{j}&0&\alpha_{j}\\ 0&-\alpha_{j}&0\\ \alpha_{j}&0&0\end{pmatrix}\begin{pmatrix}\psi_{j}(L)\\ \psi_{j}^{\prime}(L)\\ \psi_{j}^{\prime\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\psi}_{j}(L)\\ \tilde{\psi}_{j}^{\prime}(L)\\ \tilde{\psi}_{j}^{\prime\prime}(L)\end{pmatrix}\right)\\ &\hskip 14.22636pt=\left(B_{0}\partial\phi(-L)\mid\partial\tilde{\phi}(-L)\right)-\left(B_{0}\partial\phi(L)\mid\partial\tilde{\phi}(L)\right)+\sum_{j=1}^{N}\left(B_{j}\partial\psi_{j}(L)\mid\partial\tilde{\psi}_{j}(L)\right).\end{split}$ (4.2)
where
$B_{j}:=\begin{pmatrix}\beta_{j}&0&\alpha_{j}\\ 0&-\alpha_{j}&0\\ \alpha_{j}&0&0\end{pmatrix},\ j=0,\dots,N,$ (4.3)
[…]
Proof.
The proof follows by integration by parts. ∎
Next, we define the boundary-subspaces $\mathcal{B}_{\pm}$ as
+=$\mathcal{B}_{+}=\Big{\{}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}$ (4.4)
and
=$\mathcal{B}_{-}=\Big{\{}\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}.$ (4.5)
In this form, we have the following definition.
Definition 4.2.
Consider the spaces $\mathcal{B}_{+}\cong\mathcal{B}_{-}\cong\mathbb{C}^{3(k+1)}$ in (4.4)-(4.5) and endow these spaces with the non-degenerated indefinite inner products
$\left(x\mid y\right)_{\mathcal{B}_{+}}:=\left(x\mid y\right)_{+}=\left(B_{+}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{\mathcal{B}_{-}}:=\left(x\mid y\right)_{-}:=\left(B_{-}x\mid y\right),$ (4.6)
[…]
$B_{+}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{3}\\ &&&\ddots\\ &&&&B_{2k-1}\end{pmatrix}\mbox{ and }B_{-}:=\begin{pmatrix}B_{0}\\ &-B_{2}\\ &&-B_{4}\\ &&&\ddots\\ &&&&-B_{2k}\end{pmatrix}.$
$(\mathcal{B}_{\pm},\left(\cdot\mid\cdot\right)_{\pm})$ are Krein spaces.
Under these new notations, Proposition 4.1 can be rewritten in a compact form as
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}.$ (4.7)
Definition 4.3.
We define the function
$\begin{split}F:&\ G(A_{0}^{*})\to\mathcal{B}_{+}\oplus\mathcal{B}_{-}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\right)\end{split}$ (4.8)
and the sesquilinear form
$\begin{split}w:&\ (\mathcal{B}_{+}\oplus\mathcal{B}_{-})\times(\mathcal{B}_{+}\oplus\mathcal{B}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\left(x\mid\tilde{x}\right)_{+}-\left(y\mid\tilde{y}\right)_{-}.\end{split}$ (4.9)
Note from (4.1) and (4.2) that
$\Omega((\textbf{U},A_{0}^{*}\textbf{U}),(\textbf{V},A_{0}^{*}\textbf{V}))=w(F(\textbf{U},A_{0}^{*}\textbf{U}),F(\textbf{V},A_{0}^{*}\textbf{V})).$ (4.10)
By [44] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$. Note that obviously $G(A)=\{(x,Ax):x\in D(A)\}\subset G(A_{0}^{*})$
Proposition 4.2.
An extension $A$ of $A_{0}$ is skew-self-adjoint if and only if there exists a $w$-self-orthogonal subspace $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$ such that $G(A)=F^{-1}(X)$.
Proof.
It follows immediately from Definitions 2.1 and 4.3, relations (4.7) and (4.10), and from [44]. ∎
Hence, $w$-self-orthogonal subspaces $X$ parametrize the skew-self-adjoint extensions $A$ of $A_{0}$. A more explicit description of these objects is given next.
4.1.1. Extensions generating unitary dynamics
Theorem 4.3.
Let $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$. $X$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
Proof.
We follow the same ideas of [38]. For sake of completeness we carry on with the details. Let $X$ be $w$-self-orthogonal and let $(x,y)\in X$. By definition we have $w((x,y),(x,y))=0$, that is,
$(x\mid x)_{+}=(y\mid y)_{-}.$ (4.11)
Since $X$ is a subspace, by setting $x=0$ we get $0=(y\mid y)_{-}$. Since $(\cdot\mid\cdot)_{-}$ is non-degenerated it must be the case $y=0$. By linearity we can conclude that $X$ is the graph of a linear operator $\mathcal{L}:\mathcal{B}_{+}\to\mathcal{B}_{-}$. Also note that if $y=0$ it also implies $x=0$; which means $\mathcal{L}$ is injective.
To see $\mathcal{L}$ is densely defined, let $z\in D(\mathcal{L})^{\perp}$ and $x\in D(\mathcal{L})$. We have
$0=(z\mid x)=(B_{+}B_{+}^{-1}z\mid x)=(B_{+}^{-1}z\mid x)_{+}=w((B_{+}^{-1}z,0),(x,\mathcal{L}x)).$ (4.12)
Since $x\in D(\mathcal{L})$ is arbitrary, we conclude $(B_{+}^{-1}z,0)\in X^{\perp_{w}}=X=G(\mathcal{L})$. Since $\mathcal{L}$ is injective, it must be the case $B^{-1}_{+}z=0$. The latter implies $z=0$. A symmetric argument shows the image $R(\mathcal{L})$ is dense.
Let us now see #$\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
$(\mathcal{L}x\mid y)_{-}=(x\mid\mathcal{L}^{-1}y)_{+}.$
The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$. On the other hand, if $y\in D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=x$ we have
$(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}\ \ \mbox{for all }z\in D(\mathcal{L}).$
Note the latter is equivalent to say
$w((z,\mathcal{L}z),(x,y))=0\ \mbox{ for all }z\in D(\mathcal{L}),$
which implies $(x,y)\in X^{\perp_{w}}=X=G(\mathcal{L})$. Thus, $(y,x)\in G(\mathcal{L}^{-1})$. We conclude $\mathcal{L}^{-1}=\mathcal{L}^{\#}$; proving $\mathcal{L}$ is unitary.
Conversely, let $\mathcal{L}$ be a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator with $X=G(\mathcal{L})$. Let $x\in D(\mathcal{L})$. For any $z\in D(\mathcal{L})$ we have $(\mathcal{L}x\mid\mathcal{L}z)_{-}=(x\mid z)_{+}$. That is, $w((x,\mathcal{L}x),(z,\mathcal{L}z))=0$ for any $z\in D(\mathcal{L})$. We conclude $(x,\mathcal{L}x)\in X^{\perp_{w}}$, i.e., $X\subset X^{\perp_{w}}$. On the other hand, if $(x,y)\in X^{\perp_{w}}$ we have $w((x,y),(z,\mathcal{L}z))=0$ for all $z\in D(\mathcal{L})$, i.e., $(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}$ for all $z\in D(\mathcal{L})$. By definition of $\mathcal{L}^{\#}$ the latter implies $y\in D(\mathcal{L}^{\#})=D(\mathcal{L}^{-1})$ with $\mathcal{L}^{-1}y=\mathcal{L}^{\#}y=x$. We conclude $(x,y)\in G(\mathcal{L})=X$. Hence $X=X^{\perp_{w}}$ as desired. ∎
4.1.2. Examples
With the first example below we intend to resemble the so-called $\delta$ interactions on a looping-edge graph $\mathcal{G}$ in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]).
Example 1 ($\delta$-type interactions on a looping-edge graph).
Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
$A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right),$ (4.13)
and $B_{+}=B_{-}$ in Definition 4.2.
For parameters $z,m\in\mathbb{R}$ consider the operator
$\mathcal{L}:=\begin{pmatrix}1&0&0&0&0&0\\ z&1&0&0&0&0\\ \frac{z^{2}}{2}&z&1&m&0&0\\ 0&0&0&1&0&0\\ 0&0&0&z&1&0\\ -m&0&0&\frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.14)
Since #$\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. According to Proposition 4.2 and Theorem 4.3, $\mathcal{L}$ defines a skew-self-adjoint extension $A_{\mathcal{L}}$ of $A_{0}$ defined via the relation $F(G(A_{\mathcal{L}}))=G(\mathcal{L})$. That is, an extension with domain $D(A_{\mathcal{L}})$ consisting of the functions $\textbf{V}=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=z\phi(-L),\ \\ &\psi^{\prime}_{2}(L)-\psi^{\prime}_{1}(L)=z\psi_{1}(L),\ \phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{z^{2}}{2}\phi(-L)+z\phi^{\prime}(-L)+m\psi_{1}(L)\\ &\mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=\frac{z^{2}}{2}\psi_{1}(L)+z\psi_{1}^{\prime}(L)-m\phi(-L).\end{split}$ (4.15)
Note in particular that if $\textbf{V}\in D(A_{\mathcal{L}})$ is so that $\phi(-L)=\psi_{1}(L)$ we may conclude from (4.15) that
$\begin{split}&\phi(-L)=\phi(L)=\psi_{1}(L)=\psi_{2}(L),\\ &\left(\phi^{\prime}(L)+\psi^{\prime}_{2}(L)\right)-\left(\phi^{\prime}(-L)+\psi^{\prime}_{1}(L)\right)=2z\phi(-L).\end{split}$ (4.16)
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
$-L$
$L$
$\phi$
$\psi_{1}$
$\psi_{2}$
Figure 2. Examples of profiles satisfying (4.15) (on the left) and (4.16) (on the right).
More generally, consider $k\in\mathbb{Z}^{+}$ arbitrary (i.e. $2k$ half-lines attached to the common vertex $\nu=L$) and a parameter $z\in\mathbb{R}$. Under the analogous assumption $-\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $-\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
$A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right).$ (4.17)
Define the following operator with diagonal matrix of blocks
$\mathcal{L_{\delta_{Z}}}:=\begin{pmatrix}\delta_{z}\\ &\delta_{z}\\ &&\ddots\\ &&&\delta_{z}\end{pmatrix}\mbox{ where }\delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.18)
Note that each block $\delta_{z}$ is such that $B_{0}^{-1}\delta_{z}^{*}B_{0}=\delta_{z}^{-1}$ for any $z\in\mathbb{R}$ ($B_{0}$ in (4.2)) and since $\mathcal{L}_{\delta_{z}}^{\#}$ is a diagonal matrix of blocks of the form $B_{0}^{-1}\delta_{z}^{*}B_{0}$ it can be seen that $\mathcal{L}_{\delta_{z}}^{\#}=\mathcal{L}_{\delta_{z}}^{-1}$. From there it is easy to see that $\mathcal{L}_{\delta_{z}}$ is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator. From Proposition 4.2 and Theorem 4.3, $\mathcal{L}_{\delta_{z}}$ defines a skew-self-adjoint extension $A_{\mathcal{L}_{\delta_{z}}}$ of $A_{0}$ with domain
$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k.\right\}$ (4.19)
Note from (4.19) we obtained ”pair-wise” the conditions in (4.15). More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at $-L$ and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
$\begin{split}&\psi_{2j}(L)=\psi_{2j-1}(L),\ \psi_{2j}^{\prime}(L)-\psi_{2j-1}^{\prime}(L)=z\psi_{2j}(L)\mbox{ and}\\ &\psi_{2j}^{\prime\prime}(L)-\psi_{2j-1}^{\prime\prime}(L)=z\psi_{2j-1}^{\prime}(L)+\frac{z^{2}}{2}\psi_{2j-1}(L).\end{split}$ (4.20)
Moreover, if in particular $\textbf{V}\in D_{\delta_{z}}$ is such that $\phi(-L)=\psi_{2j-1}(L)$ for all $j=1,\dots,k$ then we conclude from (4.19) that
$\phi(-L)=\phi(L)=\psi_{1}(L)=\cdots=\psi_{2k}(L)\mbox{ and }\sum_{j=0}^{k}\psi^{\prime}_{2j-1}(L)-\sum_{j=0}^{k}\psi^{\prime}_{2j}(L)=(1+k)z\phi(-L).$ (4.21)
For the following examples we assume $|E_{+}|=|E_{-}|=1$. From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $366^{2}=$ 36 parameters. We exhibit a $2$ and a $4$-parameter independent families of unitary operators that maintain the looping-edge structure (i.e. the loop condition $\phi(-L)=\phi(L)$) and are not necessarily interacting pair-wise.
Example 2.
Consider again $k=1$ and $B_{+}=B_{-}$. It can be seen for $m_{1},m_{2}\in\mathbb{R}$ that
$\mathcal{L}=\begin{pmatrix}1&0&0&0&0&0\\ 0&0&0&m_{1}&1&0\\ 0&0&1&m_{2}&0&0\\ 0&0&0&1&0&0\\ 0&1&0&0&0&0\\ -m_{2}&0&0&\frac{m_{1}^{2}}{2}&m_{1}&1\end{pmatrix}$ (4.22)
is a unitary operator and therefore that it defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of the functions $v=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(-L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(-L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
Example 3.
Consider again $k=1$, $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
$\mathcal{L}=\begin{pmatrix}1&0&0&0&0&0\\ m_{1}&0&0&m_{2}&1&0\\ \frac{m_{1}^{2}+1}{2}&1&1&m_{3}&m_{1}&0\\ 0&0&0&1&0&0\\ 1&1&0&m_{4}&0&0\\ m_{4}-m_{3}+m_{1}m_{2}&m_{4}&0&\frac{m_{2}^{2}+m_{4}^{2}}{2}&m_{2}&1\end{pmatrix}$ (4.23)
defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of $v\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$
4.2. Boundary system for arbitrary amount of half-lines
For this second part we no longer seek for unitary dynamics. We feel confortable allowing the dynamics to be generated by a $C_{0}$-semigroup of contractions. To do so, we consider $\mathcal{G}$ consisting of finite amount $N$ of half-lines attached to the vertex $L$ with no restrictions on the signs of the coefficients of the Hamiltonian. We still treat $-L$ as a terminal vertex but we will be emphatic in whether the examples to be presented maintain or not the looping-edge structure. It is worth mentioning that most of the following results also apply to tadpole-like graphs.
Definition 4.4.
Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ and endow these spaces with the non-degenerated indefinite inner products
$\left(x\mid y\right)_{1}:=\left(C_{1}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{0}:=\left(C_{0}x\mid y\right),$ (4.24)
where
$C_{1}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{2}\\ &&&\ddots\\ &&&&B_{N}\end{pmatrix}\mbox{ and }C_{0}:=\frac{1}{N+1}\begin{pmatrix}B_{0}\\ &B_{0}\\ &&B_{0}\\ &&&\ddots\\ &&&&B_{0}\end{pmatrix}.$
$(\mathcal{C}_{1},\left(\cdot\mid\cdot\right)_{1})$ and $(\mathcal{C}_{0},\left(\cdot\mid\cdot\right)_{0})$ are Krein spaces.
Under these new notations, Proposition 4.1 can be rewritten as
$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}.$ (4.25)
Definition 4.5.
We consider the function
$\begin{split}\mathcal{F}:&\ G(A_{0}^{*})\to\mathcal{C}_{1}\oplus\mathcal{C}_{0}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (4.26)
4.2.1. Extensions generating contractile dynamics
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be a densely defined operator. Consider an extension $A_{\mathcal{L}}$ of $A_{0}$ defined by the relation $G(A_{\mathcal{L}})=\mathcal{F}^{-1}(G(\mathcal{L}))$. In other words,
$D(A_{\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\in D(\mathcal{L})\mbox{ and }\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right\}.$ (4.27)
acting as $A_{\mathcal{L}}\textbf{U}:=-A_{0}^{*}\textbf{U}$.
Let us determine the adjoint of operator of $A_{\mathcal{L}}$.
Proposition 4.4.
For $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ densely defined we have
$D(A_{\mathcal{L}}^{*}):=\left\{(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ and }\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right\}$ (4.28)
acting as $A_{\mathcal{L}}^{*}\textbf{U}=A_{0}^{*}\textbf{U}.$
Proof.
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$, $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})\in D(A_{0}^{*})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}[A_{\mathcal{L}}\textbf{U},\textbf{V}]&=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.29)
Therefore, $\textbf{V}\in D(A_{\mathcal{L}}^{*})$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.$ (4.30)
Note from the definition of $A_{\mathcal{L}}$ that $\mathcal{F}(G(A_{\mathcal{L}}))=G(\mathcal{L})$; which means
$D(\mathcal{L})=\left\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\textbf{U}\in D(A_{\mathcal{L}})\right\}.$ (4.31)
We conclude (4.30) is equivalent to
#$\mathcal{L}^{\#}\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}=\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}.$
From Proposition 4.4 we see that knowing the particular expression defining $\mathcal{L}^{\#}$ is important to determine explicitly the domain of $A_{\mathcal{L}}^{*}$. Note in case $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ is given by a $3(N+1)\times 3(N+1)$ matrix, the $(\mathcal{C}_{0},\mathcal{C}_{1})$-adjoint $\mathcal{L}^{\#}$ is given by the expression
#$\mathcal{L}^{\#}=C_{1}^{-1}\mathcal{L}^{*}C_{0}.$ (4.32)
In fact, for $y\in\mathcal{C}_{0}$ the definition of $\mathcal{L}^{\#}$ imply that for all $x\in\mathcal{C}_{1}$ we have
$\left(C_{0}\mathcal{L}x\mid y\right)=\left(\mathcal{L}x\mid y\right)_{0}=\left(x\mid\mathcal{L}^{\#}y\right)_{1}=\left(C_{1}x\mid\mathcal{L}^{\#}y\right)_{0}.$ (4.33)
Since $C_{1}$ is real symmetric we obtain that for all $x\in\mathcal{C}_{1}$
$\left(x\mid\ \left(\mathcal{L}^{*}C_{0}-C_{1}\mathcal{L}^{\#}\right)y\right)=0.$ (4.34)
Since $y$ is arbitrary, it follows
$\mathcal{L}^{*}C_{0}=C_{1}\mathcal{L}^{\#},$ (4.35)
which implies (4.32)
In the following two propositions we characterize when both, $A_{\mathcal{L}}$ and $A_{\mathcal{L}}^{*}$, are dissipative operators. We recall that dissipative operators on Hilbert spaces were briefly described on subsection 2.2.
Proposition 4.5.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$. $A_{\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
Proof.
Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+\overline{[A_{\mathcal{L}}\textbf{U},\textbf{U}]}=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.36)
Hence, $\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}\leqslant\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.$ (4.37)
From (4.31) we conclude that (4.37) is equivalent to $\mathcal{L}$ being a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction. ∎
Proposition 4.6.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction.
Proof.
Let $\textbf{V}\in D(A_{\mathcal{L}}^{*})$. From (4.7) we have
$\begin{split}2\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]\\ &=\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}.\end{split}$ (4.38)
Hence, $\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]\leqslant 0$ if and only if
(#$\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}\leqslant\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}\ \mbox{ for all }\textbf{V}\in D(A_{\mathcal{L}}^{*}).$ (4.39)
From the definition of $D(A_{\mathcal{L}}^{*})$ we conclude (4.39) is equivalent to $\mathcal{L}^{\#}$ being a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction. ∎
From the previous two propositions above and the Lummer-Phillips Theorem (see Theorem 2.2) we can obtain the following characterization of extensions generating $C_{0}$-semigroups of contractions.
Theorem 4.7.
Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}$ is the generator of a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39).
Proof.
Suppose $A_{\mathcal{L}}$ generates a semigroup $\{T(t)\}_{t\geqslant 0}$ of contractions. By Proposition 2.1, $A_{\mathcal{L}}$ is dissipative. Since $A_{\mathcal{L}}^{*}$ generates the semigroup $\{T(t)^{*}\}_{t\geqslant 0}$ which is also a contraction semigroup, from Proposition 2.1 we conclude $A_{\mathcal{L}}^{*}$ is dissipative. From Propositions 4.5 and 4.6 we conclude that $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39). The converse follows form Propositions 4.5 and 4.6 and Theorem 2.2. ∎
4.2.2. Examples
Let us show an example an extension $A_{\mathcal{L}}$ generating semigroups of contractions. We intent to present an operator whose first row preserves the looping-edge structure ($\phi(L)=\phi(-L)$) with continuity at the vertex and still keeping several parameters. For simplicity we consider a tadpole graph ($N=1$) with $\alpha_{0}=\alpha_{1}=\beta_{0}=\beta_{1}$.
Example 4 ($\delta$ type interactions on a tadpole graph).
Let $m_{1},\dots,m_{4}\in\mathbb{R}$ and let
$M:=\begin{pmatrix}1&0&0&0&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{1}&m_{2}+m_{3}&2&m_{2}&0&0\\ 0&0&0&1&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{3}&2&0&m_{4}&0&2\end{pmatrix}.$ (4.40)
According to (2.4), $M$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if for all $x\in\mathbb{C}^{6}$
$\left(C_{0}Mx\mid Mx\right)-\left(C_{1}x\mid x\right)\leqslant 0,$ (4.41)
i.e., if the matrix $M^{*}C_{0}M-C_{1}$ is negative semidefinite.
It can be seen that it is enough to consider
$\alpha_{0}<0,\ m_{4}\leqslant\frac{3}{2}\mbox{ and }4m_{1}-(m_{2}^{2}+m_{3}^{2})\leqslant 2.$ (4.42)
Now, according to (4.32) we have
$M^{\#}=\begin{pmatrix}1&0&0&0&0&0\\ \frac{-m_{2}-m_{3}}{2}&\frac{1}{2}&0&-1&\frac{1}{2}&0\\ \frac{m_{1}-1}{2}&\frac{-m_{2}-m_{3}}{4}&\frac{1}{2}&\frac{m_{3}}{2}&\frac{-m_{2}-m_{3}}{4}&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&0\\ \frac{m_{2}}{2}&\frac{-1}{2}&0&\frac{m_{4}-1}{2}&\frac{-1}{2}&\frac{1}{2}\end{pmatrix}.$ (4.43)
In order to $M_{z}^{\#}$ be a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction we require for all $y\in\mathbb{C}^{6}$ that
$\left(C_{1}M^{\#}y\mid M^{\#}y\right)\leqslant\left(C_{0}y\mid y\right),$ (4.44)
i.e., we require the matrix $(M^{\#})^{*}C_{1}M^{\#}-C_{0}$ to be negative semidefinite The latter holds under the exact same condition (4.42).
The operator $M$ defines an operator $A_{M}$ acting as $A_{M}\textbf{U}=-A_{0}^{*}\textbf{U}$ with domain consisting of functions $(\phi,\psi)\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=\left(\frac{m_{2}+m_{3}+2}{2}\right)\phi(-L),\\ &\phi^{\prime\prime}(L)-2\psi^{\prime\prime}(L)=(m_{3}+m_{4})\psi(L)+2\phi^{\prime}(-L)\ \mbox{ and }\\ &\phi^{\prime\prime}(L)-2\phi^{\prime\prime}(-L)=(m_{1}+m_{2})\phi(-L)+(m_{2}+m_{3})\phi^{\prime}(L).\end{split}$ (4.45)
Note if you take the parameters satisfying $m_{4}=m_{1}=0$, $m_{3}=-2$ and denoting $z=m_{2}/2\in\mathbb{R}$ it follows from (4.45)
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=z\phi(-L)\mbox{ and }\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)-\psi^{\prime\prime}(-L)=z\phi^{\prime}(-L)+(z-1)\phi(-L).\end{split}$ (4.46)
We conclude from Proposition 4.5, Proposition 4.6 and Theorem 4.7 that $A_{M}$ generates a semigroup of contractions.
4.3. Boundary system separating derivatives
A better description of the boundary behavior can be performed by separating the relation on the first derivatives exhibited by the matrices $B_{j}$ in (4.2). It can be attained by rewriting Proposition 4.1 in the following way. We recall the notation associated to U, $\vec{\textbf{{U}}}$, $\vec{\textbf{{U}}}^{\prime}$ and $\vec{\textbf{{U}}}^{\prime\prime}$ given in subsection 2.1.
Lemma 4.8.
For any $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$ we have
$\begin{split}&[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(D_{\beta}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime\prime}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right),\end{split}$ (4.47)
where
$\begin{split}&D_{\alpha}:=\begin{pmatrix}\alpha_{0}&&&&\\ &-\alpha_{0}&&&&\\ &&\alpha_{1}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\beta}:=\begin{pmatrix}\beta_{0}&&&&\\ &-\beta_{0}&&&&\\ &&\beta_{1}&&\\ &&&\ddots&\\ &&&&\beta_{N}\end{pmatrix}.\end{split}$ (4.48)
Definition 4.6.
Consider the spaces $\mathcal{K}_{+}\cong\mathcal{K}_{-}\cong\mathbb{C}^{N+1}$ and endow them with the weighted products
$\left(x\mid y\right)_{\mathcal{K}_{+}}:=\left(D^{+}_{\alpha}x\mid y\right)\mbox{ and }\left(x\mid y\right)_{\mathcal{K}_{-}}:=\left(D_{\alpha}^{-}x\mid y\right),$ (4.49)
where
$\begin{split}&D_{\alpha}^{+}:=\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{1}&&&&\\ &&\alpha_{2}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\alpha}^{-}:=\frac{1}{N+1}\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{0}&&&&\\ &&\alpha_{0}&&\\ &&&\ddots&\\ &&&&\alpha_{0}\end{pmatrix}.\end{split}$
Then $(\mathcal{K}_{\pm},(\cdot\mid\cdot)_{\mathcal{K}_{\pm}})$ are Krein spaces.
Note the matrices $D_{\alpha}^{-}$ and $D_{\alpha}^{+}$ are simply splitting the matrix $D_{\alpha}$ into two matrices, one dealing with the coefficient of $\phi(L)$ and another dealing with the rest of the coefficients. In fact,
$\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.50)
From (4.47) and (4.50) we can rewrite
$\begin{split}[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]&=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)\\ &=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\hskip 28.45274pt-\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}+\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.51)
4.3.1. Extensions generating contractile dynamics via splitting derivatives
Let us introduce extensions of $A_{0}$ that are parameterized by two items: a linear subspace $Y$ that characterize the interaction between the function and its second derivatives and a linear operation $\mathcal{L}$ that governs how the first derivatives relate each other. Different than in the case of subsection 4.2.1, we cannot obtain mixed expression involving the function and its first derivatives with these extensions $A_{Y,\mathcal{L}}$ defined next. One possible application of the results below, can be to investigate the existence and stability of solitary waves for the KdV model (1.6) on looping edge graphs in terms of the coefficients of the equations on each edge (see [39]-section 3 and Theorem 4.2).
Definition 4.7.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$. We define $A_{Y,\mathcal{L}}$ on the set
$\begin{split}D(A_{Y,\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\right.&\vec{\textbf{U}}\in Y,\ D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right\},\end{split}$ (4.52)
acting as $A_{Y,\mathcal{L}}\textbf{U}=-A_{0}^{*}\textbf{U}$.
The following proposition describe explicitly the form of the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$.
Proposition 4.9.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$. We have that the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$ is defined on the domain
$\begin{split}D(A_{Y,\mathcal{L}}^{*})=\{\textbf{V}\in D(A_{0}^{*})\mid&\vec{\textbf{V}}\in Y,\ \frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\right\}\end{split}$ (4.53)
acting as $A_{Y,\mathcal{L}}^{*}\textbf{V}=A_{0}^{*}\textbf{V}$.
Proof.
Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ and $\textbf{V}\in D(A_{0}^{*})$. We have
$\begin{split}[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]&=[-A_{0}^{*}\textbf{U},\textbf{V}]\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]-\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\ +\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.54)
In order to V be in $D(A_{Y,\mathcal{L}}^{*})$ it must satisfy $[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]=[\textbf{U},A_{Y,\mathcal{L}}^{*}\textbf{V}]$ for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$. Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that ${\textbf{U}}=c{\textbf{U}}^{\prime}=\vec{\textbf{0}}$. It follows from (4.54) that $\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\mid\vec{\textbf{V}}\right)=0$. Since by definition of $A_{Y,\mathcal{L}}$ we have $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\in Y^{\bot}$, we conclude
$\vec{\textbf{V}}\in Y.$ (4.55)
Now, Let $x\in Y$ be arbitrary. There exists $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that $\vec{\textbf{U}}=x$ with $\vec{\textbf{U}}^{\prime}=0$. Since $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}$ is in $Y^{\bot}$, from (4.54) we get
$\left(x,\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(\vec{\textbf{U}},\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}},\vec{\textbf{V}}\right)=0.$ (4.56)
We conclude
$\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}$ (4.57)
From (4.54), (4.55) and (4.57) we obtain that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$ we must have
$\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}=\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.58)
From definition of $D(A_{Y,\mathcal{L}})$, (4.58) is equivalent to have that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$
$\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.$ (4.59)
It is clear that
$\left\{\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}})\right\}\subset D(\mathcal{L}).$
Moreover, the equality holds since for any $x\in D(\mathcal{L})$ there exists $\textbf{V}\in D(A_{Y,\mathcal{L}})$ such that $x=\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}$. Therefore, (4.59) is equivalent to say
$\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}.$ (4.60)
Proposition 4.10.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
Proof.
Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$. We have
$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{Y,\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.61)
Since $\textbf{U}\in D(A_{Y,\mathcal{L}})$ is arbitrary we conclude (arguing as in the proof of Proposition 4.9) that $\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction. ∎
The following two propositions characterize when the extensions $A_{Y,\mathcal{L}}$ are dissipative.
Proposition 4.11.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
Proof.
Let $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$. We have
$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]=\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}-\left(\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.\end{split}$ (4.62)
It is clear that
$\left\{\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})\right\}\subset D(\mathcal{L}^{\#}).$
Moreover, the equality holds since for any $y\in D(\mathcal{L}^{\#})$ there is $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$ such that
$y=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}.$
From (4.62) we conclude $A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
The following theorem is a direct consequence of the previous two propositions and can be proved as done in the proof of Theorem 4.7.
Theorem 4.12.
Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. $A_{Y,\mathcal{L}}$ generates a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction and $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
4.3.2. Examples
Let us first describe the operators $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ that define contractions so that its respective adjoint $\mathcal{L}^{\#}$ is also a contraction. Let us assume for simplicity $N=1$ (i.e. a tadpole graph) and $\alpha_{0}=\alpha_{1}$.
Example 5.
For $m_{1},m_{2},m_{3}\in\mathbb{R}$ with $m_{3}\neq 0$, consider the operator
$\mathcal{L}:=\begin{pmatrix}m_{1}&m_{2}\\ m_{3}&\frac{-m_{1}m_{2}}{m_{3}}\end{pmatrix}.$
Note we have
#$\mathcal{L}^{\#}=\begin{pmatrix}\frac{m_{1}}{2}&\frac{m_{3}}{2}\\ \frac{m_{2}}{2}&\frac{-m_{1}m_{2}}{2m_{3}}\end{pmatrix}.$
It can be seen that both, $\mathcal{L}$ and $\mathcal{L}^{\#}$ are contractions under the conditions
$m_{3}^{2}+m_{1}^{2}\leqslant 2,\ m_{2}^{2}(m_{1}^{2}+1)\leqslant 2,\ m_{1}(m_{3}^{2}-m_{2}^{2})=0,\ m_{1}^{2}+m_{2}^{2}\leqslant 1\ \mbox{and }\frac{m_{1}^{2}m_{2}^{2}}{m_{3}^{2}}+m_{3}^{2}\leqslant 2.$ (4.63)
In case $m_{3}=m_{2}\neq 0$, (4.63) reduces to
$m_{2}^{2}+m_{1}^{2}\leqslant 2\mbox{ and }m_{2}^{2}(m_{1}^{2}+1)\leqslant 2.$
In particular, if we take $m_{2}=m_{1}=1$ you get that
$\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$
defines a contraction with $(\mathcal{K}_{-},\mathcal{K}_{+})$-adjoint also a contraction.
The next example pretends to characterize the extensions obtained via separating derivatives that maintain the looping-edge structure ($\phi(-L)=\phi(L)$) and include continuity at the vertex (when $z=1$ below).
Example 6.
Let $z\in\mathbb{R}$ be fixed and $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a linear operator. Consider the subspace $Y_{z}:=\mbox{span}\{(1,1,z)^{T}\}$. Let us first assume $z\neq 0$. We have that
$Y_{z}^{\perp}=\left\{(x,y,w)^{T}\in\mathbb{C}^{3}\mid w=\frac{-x-y}{z}\right\}.$
From the Definition 4.7 we have that the domain of $A_{Y_{z},\mathcal{L}}$ is made of functions $u\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=z\phi(L),\ z\left(\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right)\\ &\mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$ (4.64)
For instance, if we take $z=1$ and $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we have that $A_{Y_{1},\mathcal{L}}$ with domain
$\begin{split}D(A_{Y_{1},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L)=\psi(L),\ \phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\ \psi^{\prime}(L)=0\mbox{ and }\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)=\alpha_{0}(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L))\}.\end{split}$ (4.65)
defines a contractile dynamic on the tadpole graph.
$-L$
$L$
$\psi$
$\phi$
Figure 3. Example of profile in (4.66)
On the other hand, assume $z=0$. In such case
$Y_{0}^{\perp}=\{(x,y,w)\in\mathbb{C}^{3}\mid x=-y\}.$
Therefore the domain of $A_{Y_{0},\mathcal{L}}$ is made of functions $\textbf{U}\in D(A_{0}^{*})$ such that
$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=0,\ \phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\ \mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$
In particular if we take $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we get that $A_{Y_{0},\mathcal{L}}$ with domain
$\begin{split}D(A_{Y_{0},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L),\ \psi(L)=0,\ \phi^{\prime}(L)=\phi^{\prime}(-L)\mbox{ and }\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\}.\end{split}$ (4.66)
defines an extension of $A_{0}$ generating a contractive dynamic on the tadpole.
5. The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped graphs
In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex. This strategy will allow us in the future to study the dynamics of NLS model (1.1) specially on $\mathcal{T}$-graphs.
Let us now consider the Schrödinger operator
$\mathcal{H}_{0}:\{u_{e}\}_{e\in E}\mapsto\left\{-\frac{d^{2}}{dx_{e}^{2}}u_{e}\right\}_{e\in E},$ (5.1)
seen as a densely defined operator on $L^{2}(\mathcal{G})$ with domain $D(\mathcal{H}_{0})$ defined on (1.3). Note we would have $D(\mathcal{H}_{0}^{*})=H^{2}(\mathcal{G})$.
5.1. Deficiency indices of the free Schrödinger operator
For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $-\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
$u_{e}=c_{1}e^{\frac{(1\pm i)}{\sqrt{2}}x_{e}}+c_{2}e^{-\frac{(1\pm i)}{\sqrt{2}}x_{e}}.$
We note that $e^{\pm\frac{(1\pm i)}{\sqrt{2}}x}$ belong to $L^{2}(-L,L)$. On the other hand, on a half-line $[L,\infty)$, it would be square integrable only if the real part of ±$\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative. Along each edge $e=[L,\infty)$ we would therefore always have that the deficiency indices are ${d_{+}}_{e}={d_{-}}_{e}=1$.
In a looping-edge graph with $N$ half-lines attached at the vertex we would have for $\mathcal{H}_{0}$ that the deficiency indices $d_{\mp}=dim(Ker(\mathcal{H}_{0}\pm iI))$ satisfy $d_{+}=d_{-}=2+N$. From the von Neumann and Krein theory, one can always expect self-adjoint extensions of $\mathcal{H}_{0}$ and therefore unitary dynamics.
5.2. Boundary systems
In this subsection we adapt the ideas developed in the case of the Airy operator, for the Schrödinger operator posed on $\mathcal{G}$.
Definition 5.1.
We define in the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by
[…]
Proposition 5.1.
For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(P_{+}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)-\left(P_{-}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)$ (5.3)
where
$P_{+}:=\begin{pmatrix}P_{0}\\ &P_{0}\\ &&\ddots\\ &&&P_{0}\end{pmatrix}\ \mbox{with }P_{0}:=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\ \mbox{and }P_{-}:=\frac{1}{N+1}P_{+}.$
Proof.
The proof follows integrating by parts. ∎
Definition 5.2.
Consider the spaces $\mathcal{J}_{+}\cong\mathcal{J}_{-}\cong\mathbb{C}^{2(N+1)}$ edowed with the indefinite non-degenerated products
$\langle x,y\rangle_{+}:=\left(P_{+}x\mid y\right)\ \mbox{and }\langle x,y\rangle_{-}=\left(P_{-}x\mid y\right).$
$(\mathcal{J}_{\pm},\langle\cdot,\cdot\rangle_{\pm})$ are Krein spaces.
Definition 5.3.
We define the function
$\begin{split}\mathbb{F}:&\ G(\mathcal{H}_{0}^{*})\to\mathcal{J}_{+}\oplus\mathcal{J}_{-}\\ &(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (5.4)
and the sesquilinear form
$\begin{split}\rho:&\ (\mathcal{J}_{+}\oplus\mathcal{J}_{-})\times(\mathcal{J}_{+}\oplus\mathcal{J}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\langle x,\tilde{x}\rangle_{+}-\langle y,\tilde{y}\rangle_{-}.\end{split}$ (5.5)
From (5.2), (5.3) and the previous definition we conclude for $u\in D(\mathcal{H}_{0}^{*})$ that
$\Lambda((\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}))=\rho(\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})).$ (5.6)
It follows directly from (5.6) and [44] the following theorem.
Theorem 5.2.
Let $\mathcal{H}$ be an extension of $\mathcal{H}_{0}$. $\mathcal{H}$ is self-adjoint if and only if there exists a subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$, $X$ $\rho$-self-orthogonal with $\mathbb{F}(G(\mathcal{H}))=X$.
5.3. Extensions generating unitary dynamics
In a similar fashion as done to obtain Theorem 4.3 in the case of the Airy operator, we can characterize the $\rho$-self-orthogonal subspaces in terms of the graph of a linear operator that, according to Theorem 5.2, characterize the self-adjoint extensions of $\mathcal{H}_{0}$.
Theorem 5.3.
A subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$ is $\rho$-self-orthogonal if and only if there is $\mathcal{L}:\mathcal{J}_{+}\to\mathcal{J}_{-}$ unitary such that $X=G(\mathcal{L})$.
Proof.
The proof follows in a similar fashion as in the proof of Theorem 4.3. ∎
With our first example we show that the extensions inducing the so-called $\delta$ interactions can be seen as a member of a $\frac{(N+2)(N+1)}{2}$-parameters family of extensions generating unitary dynamics.
Example 7 ($\delta$ type interactions on looping edge graphs).
Let us first set $N=1$ and let $\mathcal{L}_{1}$ be defined by
$\mathcal{L}_{1}=\begin{pmatrix}1&0&0&0\\ m_{1}&2&m_{2}&0\\ 0&0&1&0\\ m_{2}&0&m_{3}&2\end{pmatrix}.$ (5.7)
It is not hard to see that $\mathcal{L}_{1}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From Theorems 5.2 and 5.3 $\mathcal{L}_{1}$ defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}_{1}}$ of $\mathcal{H}_{0}$ via the relation $\mathbb{F}(G(\mathcal{H}_{\mathcal{L}_{1}}))=G(\mathcal{L}_{1})$. That is, an extension $\mathcal{H}_{\mathcal{L}_{1}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=(m_{1}+m_{2})\phi(-L),\ \mbox{and }\\ &\phi^{\prime}(L)-2\psi^{\prime}(L)=(m_{2}+m_{3})\phi(-L).\end{split}$ (5.8)
Adding the equations in (5.8) we get
$\phi^{\prime}(L)-\phi^{\prime}(-L)=\psi^{\prime}(L)+\frac{m_{1}+2m_{2}+m_{3}}{2}\phi(-L).$ (5.9)
Of course by taking $m_{1}=-m_{3}$ we recover $D_{Z,1}$ defined in (1.5).
More generally, let us consider $N\in\mathbb{Z}^{+}$ arbitrary and parameters $m_{ij}\in\mathbb{R}$ such that $m_{ij}=m_{ji}$ for $i,j\in\{1,\dots,N+1\}$. We define $\mathcal{L}_{N}$ as a $2(N+1)\times 2(N+1)$ matrix given in terms of $2\times 2$ blocks $(A_{ij})_{ij}$ obeying the rule
$A_{ii}=\begin{pmatrix}1&0\\ m_{ii}&N+1\end{pmatrix}\mbox{ and }A_{ij}=\begin{pmatrix}0&0\\ m_{ij}&0\end{pmatrix}\mbox{ if }i\neq j.$ (5.10)
We abuse notation to denote with $\psi_{0}(L):=\phi(-L)$ and $\psi^{\prime}_{0}(L)=\phi^{\prime}(-L)$. Note that if $i=j$ we have that
$A_{ii}\partial\psi_{i-1}(L)=\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi^{\prime}_{i-1}(L)\end{pmatrix}$ (5.11)
and if $i\neq j$ we have
$A_{ij}\partial\psi_{j-1}(L)=\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}.$ (5.12)
From the structure of the rows of $\mathcal{L}_{N}$ it can be seen that it is a full-rank invertible matrix. In particular we conclude $\mathcal{L}_{N}$ is injecitve. Moreover, a tedious but elementary computation shows that $\mathcal{L}_{N}^{*}P_{-}\mathcal{L}_{N}=P_{+}$, which implies that for all $x,y\in\mathbb{C}^{2(N+1)}$
$\langle\mathcal{L}_{N}x,\mathcal{L}_{N}y\rangle_{-}=\langle x,y\rangle_{+}.$
The latter implies $\mathcal{L}_{N}^{\#}=\mathcal{L}_{N}^{-1}$, showing $\mathcal{L}_{N}$ defines a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From Theorems 5.2 and 5.3, $\mathcal{L}_{N}$ defines a self-adjoint extension $\mathcal{H_{\mathcal{L}_{N}}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions in $D(\mathcal{H}_{0}^{*})$ such that
$\mathcal{L}_{N}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{1}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}.$ (5.13)
Note from the even rows of $\mathcal{L}_{N}$ that we are a priori fixing $\phi(-L)=\phi(L)=\psi_{1}(L)=\cdots=\psi_{N}(L)$. Taking this into account, let us open the expression (5.13). We have
$\begin{split}&A_{11}\partial\psi_{0}(L)+A_{12}\partial\psi_{1}(L)+\cdots+A_{1(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &A_{21}\partial\psi_{0}(L)+A_{22}\partial\psi_{1}(L)+\cdots+A_{2(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &\hskip 142.26378pt\vdots\\ &A_{(N+1)1}\partial\psi_{0}(L)+A_{(N+1)2}\partial\psi_{1}(L)+\cdots+A_{(N+1)(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ \end{split}$ (5.14)
Adding the equations in (5.14) we get
$\sum_{i=1}^{N+1}A_{i1}\partial\psi_{0}(L)+A_{i2}\partial\psi_{1}(L)+\cdots+A_{i(N+1)}\partial\psi_{N}(L)=(N+1)\partial\phi(L).$ (5.15)
According to (5.11) and (5.12), we can rewrite (5.15) as
$\frac{1}{N+1}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N+1}\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}+\frac{1}{N+1}\sum_{i=1}^{N+1}\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi_{i-1}^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\end{pmatrix}.$ (5.16)
From the continuity at the vertex and the second entries in (5.16) we conclude
$\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{i=1}^{N}\psi^{\prime}_{i}(L)+\phi(-L)\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}.$ (5.17)
Any choice of the parameters such that $\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}=Z\in\mathbb{R}$ makes $D(\mathcal{H}_{\mathcal{L}_{N}})=D_{Z,N}$, where $D_{Z,N}$ is defined on (1.5).
Remark 1.
It is worth mentioning that the operator $\mathcal{L}_{N}$ was built by fixing continuity at the vertex (see the even rows of $\mathcal{L}_{N}$) and removing the problematic entries that won’t allow $\mathcal{L}_{N}$ to be a real unitary matrix in its boundary system. Therefore, we can infer from the previous example that all the self-adjoint extensions of the Schrödinger operator coming from a real matrix that include continuity at the vertex and are posed on a looping-edge graph are of the form $(\mathcal{H},D(\mathcal{H}))=(\mathcal{H}_{0},D_{Z,N})$ with $Z$ varying on $\mathbb{R}$.
The following example pretend to explore families of operators that induce unitary dynamics and are compatible with continuity of the derivatives at the vertex.
Example 8 ($\delta^{\prime}$ type interactions on looping edge and $\mathcal{T}$-graphs).
Consider $N=1$ (tadpole case). At first, we begin considering $-L$ as an independent vertex, i.e., we won’t fix the condition $\phi(-L)=\phi(L)$ directly from the operator. Let $m_{1},\dots,m_{4}\in\mathbb{R}$ be parameters. Define
$\mathcal{L}=\begin{pmatrix}m_{1}&m_{2}&0&0\\ 0&\frac{2}{m_{1}}&0&m_{3}\\ -m_{1}m_{3}&-m_{2}m_{3}&2&m_{4}\\ 0&0&0&1\end{pmatrix}.$ (5.18)
It can be seen that $\mathcal{L}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator. From theorems 5.2 and 5.3 we have a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ with domain made of the functions in $D(\mathcal{H}_{0}^{*})$ such that
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)-m_{1}\phi(-L)=m_{2}\phi^{\prime}(-L),\ \phi^{\prime}(L)-2m_{1}^{-1}\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{1}m_{3}\phi(-L)=-m_{2}m_{3}\phi^{\prime}(-L)+2\psi(L)+m_{4}\psi^{\prime}(L).\end{split}$ (5.19)
In particular if we take $m_{1}=1$ and $m_{2}=0$ we recover the looping structure $\phi(-L)=\phi(L)$ since (5.19) can be rewritten as
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)=\phi(-L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{3}\phi(-L)-2\psi(L)=m_{4}\psi^{\prime}(L).\end{split}$ (5.20)
Moreover, if we additionally let $m_{3}=-1$ in (5.20) we get
$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L)=\phi^{\prime}(-L),\ \phi(L)=\phi(-L),\\ &\mbox{ and }\psi(L)=\frac{-m_{4}}{2}\psi^{\prime}(L).\end{split}$ (5.21)
Remark 2.
The existence and stability of standing waves profiles for the NLS model (1.1) on a tadpole graph, ${\bf{U}}(x,t)=e^{-i\omega t}(\phi,\psi)$, with $(\phi,\psi)$ satisfying the $\delta^{\prime}$-type interactions in (5.21) is the aim of a work in progress by the authors.
5.4. Extensions generating unitary dynamics parameterized by linear subspaces
We now emulate the strategy of separating derivatives used for the Airy operator. We note in this case there is no actual separation of derivatives but another reference frame to build the extensions. We begin by following the same approach as done in Section 4.3 to rewrite (5.3) in a simpler way.
Proposition 5.4.
For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(Q\vec{\textbf{U}}\mid\vec{\textbf{V}}^{\prime}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right),$ (5.22)
where
$Q=\begin{pmatrix}1\\ &-1\\ &&1\\ &&&\ddots\\ &&&&1\end{pmatrix}.$
Definition 5.4.
Given a subspace $Y\subset\mathbb{C}^{2+N}$ consider the operator $\mathcal{H}_{Y}$, acting as $\mathcal{H}_{Y}\textbf{U}=\mathcal{H}_{0}\textbf{U}$, with domain
$D(\mathcal{H}_{Y}):=\left\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{U}}\in Y\ \mbox{and }Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}\right\}.$ (5.23)
The next proposition determine explicitly the adjoint operator of $\mathcal{H}_{Y}$.
Proposition 5.5.
For any closed subspace $Y\subset\mathbb{C}^{2+N}$ the operator $\mathcal{H}_{Y}$ is a self-adjoint extension of $\mathcal{H}_{0}$.
Proof.
It is enough to prove that the adjoint of $\mathcal{H}_{Y}$ is given by $\mathcal{H}_{Y}^{*}\textbf{V}=\mathcal{H}_{0}^{*}\textbf{V}$ with domain
$D(\mathcal{H}_{Y}^{*}):=\left\{\textbf{V}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{V}}\in Y\ \mbox{and }Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}\right\}=D(\mathcal{H}_{Y}).$ (5.24)
It is obvious from the definition of $D(\mathcal{H}_{Y})$ that if $Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}$ and $\vec{\textbf{V}}\in Y$ then $\textbf{V}\in D(\mathcal{H}_{Y}^{*})$.
Let us prove the converse. Let $\textbf{V}\in D(\mathcal{H}_{0}^{*})$. By definition V is in $D(\mathcal{H}_{Y}^{*})$ if and only if $[\mathcal{H}_{Y}\textbf{U},\textbf{V}]=[\textbf{U},\mathcal{H}_{Y}^{*}\textbf{V}]$ for all $\textbf{U}\in D(\mathcal{H}_{Y})$.
Let $x\in Y^{\perp}$ be arbitrary. There exists $\textbf{U}\in D(\mathcal{H}_{Y})$ such that $\vec{\textbf{U}}=0$ and $Q\vec{\textbf{U}}^{\prime}=x$. From (5.22), we must have
$\left(x\mid\vec{\textbf{V}}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)=0.$
Since $x$ was arbitrary, it must be the case $\vec{\textbf{V}}\in(Y^{\perp})^{\perp}=Y$.
Similarly, for $x\in Y$ there exists $\textbf{U}\in D(\mathcal{H}_{Y})$ such that $\vec{\textbf{U}}^{\prime}=0$ and $\vec{\textbf{U}}=x$. From (5.22) we conclude it must be the case
$\left(x\mid Q\vec{\textbf{V}}^{\prime}\right)=\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right)=0.$
We therefore have $Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}$. ∎
Let us give some examples on the tadpole graph. Assume $N=1$.
Example 9.
Consider the space $Y_{0}=\mbox{span}\{(1,1,0)^{T}\}$. Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$. Also, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\phi^{\prime}(L)$. We therefore have that
$D(\mathcal{H}_{Y_{0}})=\{u\in D(\mathcal{H}_{0}^{*})\mid\ \phi(-L)=\phi(L)\ \mbox{and }\phi^{\prime}(L)=\phi^{\prime}(-L)\}.$
Example 10.
Consider the space $Y=\mbox{span}\{(1,0,-1)\}$. Consequently, $Y^{\perp}=\{(x,y,w)\mid x=w\}$. Note $\vec{\textbf{U}}\in Y$ if and only if $\phi(L)=0$ and $\phi(-L)=-\psi(L)$. Similarly, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\psi^{\prime}(L)$. According to Proposition 5.5, ${\mathcal{H}}_{Y}$ defines a self-adjoint extension of $\mathcal{H}_{0}$ with domain
$D(\mathcal{H}_{Y})=\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\phi(-L)=-\psi(L),\ \phi(L)=0,\ \mbox{and}\ \phi^{\prime}(-L)=\psi^{\prime}(L)\}.$ (5.25)
5.5. The Schrödinger operator on $\mathcal{T}$-shaped metric graphs
As highlighted in the introduction of the section, the analysis performed in the previous subsections for the Schrödinger equation on looping-edge graphs extends to $\mathcal{T}$-shaped graphs, i.e., to graphs $\mathcal{G}$ that consists on an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with 00 (see Figure 4). A similar frame-work can be employed to obtain self-adjoint extensions of $\mathcal{H}_{0}$ posed on these kind of metric graphs. We proceed to exhibit some examples of them.
$L$
00
Figure 4. $\mathcal{T}$-shaped metric graph.
Example 11.
Let $N=1$ (exactly one half-line attached to 00) and consider again the operator $\mathcal{L}$ defined in (5.18) with $m_{1}=1$ and $m_{2}=0$ (see (5.20)). Instead of $m_{3}=-1$, let us now consider $m_{3}=1$. It can be obtained analogously to (5.20) an extension of $\mathcal{H}_{0}$ with domain of functions in $H^{2}(\mathcal{G})$ such that
$\begin{split}&\psi^{\prime}(0+)=\phi^{\prime}(0-),\ \phi^{\prime}(-L)=0,\ \phi(0-)=\phi(-L)\ \\ &\mbox{ and }\phi(0-)-\psi(0+)=\frac{m_{4}}{2}\psi^{\prime}(0+),\end{split}$ (5.26)
where $0+$ and $0-$ refer to the right and left-handed limits respectively.
Thus, (5.26) is an example of the kind of domain of extensions that might be related to interactions of $\delta^{\prime}$-type in a $\mathcal{T}$-graph.
With the following example we intend to present, under this frame of reference, all the extensions generating unitary dynamics with continuity in the derivatives at the vertex. This family includes the so-called $\delta^{\prime}$ interactions. For simplicity we work with $N=1$ and $N=2$.
Example 12.
Let us first set $N=1$ and let $\mathcal{L}$ be defined by
$\mathcal{L}=\begin{pmatrix}2&m_{1}&0&m_{2}\\ 0&1&0&0\\ 0&m_{2}&2&m_{3}\\ 0&0&0&1\end{pmatrix}.$ (5.27)
It can be seen that in a respective boundary system $\mathcal{L}$ is a unitary operator which defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ via the relation $\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(0+)\end{pmatrix}=\begin{pmatrix}\partial\phi(0-)\\ \partial\phi(0-)\end{pmatrix}$. That is, an extension $\mathcal{H}_{\mathcal{L}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}(0+),\ \phi(0+)-2\phi(-L)=(m_{1}+m_{2})\psi^{\prime}(0+)\mbox{ and }\\ &\phi(0-)-2\psi(0+)=(m_{2}+m_{3})\psi^{\prime}(0+).\end{split}$ (5.28)
Adding the last two equations in (5.28) we get
$\phi(0-)-\phi(-L)=\psi(0+)+\frac{m_{1}+2m_{2}+m_{3}}{2}\psi^{\prime}(0+).$ (5.29)
In particular, if $m_{1}=-m_{3}$ it follows
$\phi(0-)-\phi(-L)=\psi(0+)+m_{2}\psi^{\prime}(0+).$ (5.30)
Similarly, let us assume $N=2$ and consider the operator
$\mathcal{L}_{2}=\begin{pmatrix}3&m_{1}&0&m_{2}&0&m_{3}\\ 0&1&0&0&0&0\\ 0&m_{2}&3&m_{4}&0&m_{5}\\ 0&0&0&1&0&0\\ 0&m_{3}&0&m_{5}&3&m_{6}\\ 0&0&0&0&0&1\end{pmatrix}$ (5.31)
It can be seen that $\mathcal{L}_{2}$ defines a unitary extension $H_{\mathcal{L}_{2}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions $(\phi,\psi_{1},\psi_{2})\in H^{2}(\mathcal{G})$ such that
$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}_{1}(0+)=\psi^{\prime}_{2}(0+),\ \phi(0-)-3\phi(-L)=(m_{1}+m_{2}+m_{3})\psi^{\prime}(0+)\\ &\phi(0-)-3\psi_{1}(0+)=(m_{2}+m_{4}+m_{5})\psi^{\prime}(0+)\mbox{ and }\phi(0-)-3\psi_{2}(0+)=(m_{3}+m_{5}+m_{6})\psi^{\prime}(0+).\end{split}$ (5.32)
Adding the last three equations we get
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+(m_{1}+m_{4}+m_{6}+2m_{2}+2m_{3}+2m_{5})\psi^{\prime}(0+).$ (5.33)
Of course there is a choice of the parameters that turn (5.33) into
$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.34)
Moreover, the operator $\mathcal{L}_{2}$ can be generalized to any amount of half-lines $N\in\mathbb{Z}^{+}$ by following a similar procedure to the one done in Example 7. That way it can be obtained an extension with domain made of functions satisfying continuity in the derivatives and
$\phi(0-)-\phi(-L)=\sum_{j=1}^{N}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.35)
Acknowledgments
A. Muñoz is partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil. J. Angulo was partially funded by CNPq/Brazil Grant.
Data availability statement. The data that support the findings of this study are available upon request from the authors.

### v4

@@
 Abstract
-The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a graph consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex.
-Characterizations of extensions generating unitary and contractile dynamics are provided for the Airy operator in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex.
-Employing the same abstract techniques used in the former case, we characterize the self-adjoint extensions of the Schrödinger operator on $\mathcal{T}$-shaped graphs, i.e., to graphs that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$.
-The theory established in this manuscript can be considered as a first step towards a better understanding of the dynamics for the remarkable dispersive models NLS and KdV on metric graphs.
+The aim of this work is to study the Airy and Schrödinger operators on looping-edge graphs, a class of metric graphs consisting of a circle and a finite number $N$ of infinite half-lines attached to a common vertex.
+For the Airy operator, we characterize all extensions generating unitary and contractive dynamics in terms of self-orthogonal subspaces and linear operators acting on indefinite inner product spaces (Krein spaces) associated to the boundary values at the vertex.
+Employing similar abstract techniques, we then describe a systematic way to produce self-adjoint extensions of the Schrödinger operator that are compatible with prescribed boundary relations on looping-edge and $\mathcal{T}$-shaped graphs.
 Mathematics Subject Classification (2020). Primary 35Q51, 35Q55, 81Q35, 35R02; Secondary 47E05.
@@
 The flexibility in setting the geometry of the graph configuration allows for the creation of different dynamics.
-To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [17, 19, 20, 24, 25, 27, 30, 35, 37, 40, 45] and reference therein).
+To mention a few examples of bifurcated systems: Josephson junction structures, networks of planar optical waveguides and fibers, branched structures associated with DNA, blood pressure waves in large arteries, nerve impulses in complex arrays of neurons, conducting polymers, and Bose-Einstein condensation (see [19, 21, 22, 26, 27, 29, 32, 38, 40, 43, 50] and references therein).
 Recently, nonlinear models on graphs, such as the nonlinear Schrödinger equation, the sine-Gordon model, and the Korteweg-de Vries model, have been studied extensively (see [2, 13, 14, 10, 11, 15, 16, 17, 3, 32, 43] and references therein).
-From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the ”topology of the graph”), which governs the dynamics on the network via groups.
+From a mathematical viewpoint, an evolution model on a graph is equivalent to a system of PDEs defined on the edges (intervals), where the coupling is determined exclusively by the boundary conditions at the vertices (known as the “topology of the graph”), which governs the dynamics on the network via groups.
 This area of study has attracted significant attention, particularly in the context of soliton transport.
@@
 Each edge $e_{j}$ can be identified with a finite or infinite interval of the real line, $I_{e}$.
-Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical ”wires” or ”networks” connecting them.
+Thus, the edges of $\mathcal{G}$ are not merely abstract relations between vertices but can be viewed as physical “wires” or “networks” connecting them.
 The notation $e\in E$ will be used to indicate that $e$ is an edge of $\mathcal{G}$.
 This identification introduces the coordinate $x_{e}$ along the edge $e$.
-An example of a metric graph, which will be the focus of our study here, is a looping-edge graph.
+An example of a metric graph, which will be the main focus of our study here, is a looping-edge graph.
 This is a graph composed of a circle and a finite number $N$ of infinite half-lines attached to a common vertex $\nu=L$.
-If the circle is identified with the interval $[-L,L]$ and the half-lines with $[L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1).
-We use the subscripts 00 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively.
+If the circle is identified with the interval [L,L][-L,L] and the half-lines with $L,\infty)$, we obtain a particular metric graph structure that we denote with $\mathcal{G}$, represented by $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$ and $V=\{L\}$ (see Figure 1).
+We use the subscripts 0 and $j$ to refer to the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$, $j=1,\dots,N$, respectively.
 The special case $N=1$ is known as a tadpole graph (see also Figure 1).
-$-L$
+L-L
 $L$
-$-L$
+L-L
 $L$
@@
 Figure 1. Tadpole and looping-edge-type graphs
-A wave function U defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\textbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
+A wave function $\mathbf{U}$ defined on $\mathcal{G}$ will be understood as an $(N+1)$-tuple of functions $\mathbf{U}=(\phi,(\psi_{j})_{j=1}^{N})=(\phi,\psi_{1},\dots,\psi_{N})$, where $\phi$ is defined on $e_{0}=[-L,L]$ and $\psi_{j}$ is defined on $e_{j}=[L,\infty)$ for $j=1,\dots,N$.
+The primary aim of this manuscript is to develop a self-contained and systematic theory of linear dynamics for both the Airy and the Schrödinger operators on looping-edge graphs, intended to serve as a reference for future nonlinear applications.
+We emphasize that this abstract theory is presented as a standalone contribution: the dynamical part is established here independently, so that subsequent studies on existence, stability, controllability, or stabilization of nonlinear waves on these graphs can build upon it without having to reconstruct the linear theory from scratch.
+For the Airy operator, to the best of our knowledge, no results on the characterization of linear dynamics on looping-edge graphs appear in the existing literature.
+We fill this gap by providing, in terms of Krein space techniques, a complete description of all boundary conditions at the vertex that make the Airy operator the generator of a unitary group or a contraction semigroup.
+For the Schrödinger operator, our approach offers an additional structural advantage compared to the classical methods: it allows one to prescribe a priori the type of vertex interactions desired (for instance, continuity of derivatives at the vertex) and then to systematically identify all self-adjoint extensions compatible with those conditions.
+In this way the theory does not merely produce a parameter family of extensions, but organizes them according to physically or geometrically motivated constraints.
 Consider the case of nonlinear Schrödinger-type equations (NLS)
-$i\textbf{U}_{t}\ +\Delta\textbf{U}+(p+1)|\textbf{U}|^{2p}\textbf{U}=\textbf{0},\quad p>0,$ (1.1)
+$i\mathbf{U}_{t}+\Delta\mathbf{U}+(p+1)|\mathbf{U}|^{2p}\mathbf{U}=\mathbf{0},\quad p>0,$ (1.1)
 where the action of the Laplacian operator $\Delta$ on a general graph $\mathbb{G}$ is given by
-$-\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
+$\Delta:(u_{e})_{e\in E}\to(-u^{\prime\prime}_{e})_{e\in E}.$ (1.2)
 The NLS in (1.1) has been extensively studied in recent literature for various types of metric graphs $\mathbb{G}$ and with specific domains for the Laplacian, which make it a self-adjoint operator on $L^{2}(\mathbb{G})$ (see the review manuscript [37]).
@@
 $D(\mathcal{H}_{0})=C_{0}^{\infty}(-L,L)\oplus\bigoplus_{j=1}^{N}C_{0}^{\infty}(L,+\infty).$ (1.3)
-Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $-\Delta_{ext}$ is given by $-\Delta_{ext}=-\Delta$.
+Since the deficiency indices of $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ are equal to $2+N$, the extension theory of Krein and von Neumann provides a $(2+N)^{2}$-parameter family of self-adjoint extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ of $\mathcal{H}_{0}$ on $L^{2}(\mathcal{G})$, where the action of $\Delta_{ext}$ is given by $\Delta_{ext}=-\Delta$.
 Each member of this family defines a unitary dynamic for the linear evolution problem
-$\left\{\begin{array}[]{ll}i\textbf{U}_{t}=-\Delta_{ext}\textbf{U}\\ \textbf{U}(0)=\textbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right.$ (1.4)
-We point out that our approach, based on Krein spaces [44], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point $-L$ as a terminal vertex through a relatively simple matrix process (see Theorems 5.2-5.3 below).
-At this level, our approach can be seen as a complement to the works of Exner et al.
+$\left\{\begin{array}[]{ll}i\mathbf{U}_{t}=-\Delta_{ext}\mathbf{U}\\ \mathbf{U}(0)=\mathbf{u}_{0}\in D(-\Delta_{ext}).\end{array}\right$. (1.4)
+We point out that this approach, based on boundary systems [49], demonstrates that every extension operator $(-\Delta_{ext},D(-\Delta_{ext}))$ can be characterized by the boundary conditions at the vertices $\nu=\pm L$, treating the point L-L as a terminal vertex through a relatively simple matrix process (see Theorem 5.2 below).
+At this level, this part of the work can be seen as a complement to the works of Exner et al.
 [29, 30, 31], which, to our knowledge, has not formally appeared in the literature on $\mathcal{G}$.
-Furthermore, in general, the extensions $(-\Delta_{ext},-D(\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of $[-L,L]$.
-(see, for instance, Example 8 associated to $\delta^{\prime}$- type interactions on looping-edge graphs and Example 10 below by considering $-L$ as a terminal vertex).
-Moreover, our approach to obtaining these extensions will be based on those associated with the Airy operator (1.7) which is the main focus of our study.
-Both extensions theories to be established can be seen as a first step for future studies on the existence and stability of soliton profiles for both NLS-(1.1) and KdV-(1.8) models, for instance, on a looping-edge graph or on a $\mathcal{T}$-shaped graph, i.e., a graph that consists of an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with $\nu=0$ (see Figure 4).
+Furthermore, in general, the extensions $(-\Delta_{ext},D(-\Delta_{ext}))$ do not necessarily satisfy continuity conditions at the boundary of [L,L][-L,L].
 In the case of looping-edge graphs, the extensions must ensure that the wave functions are continuous at least on the loop (i.e., $\phi(L)=\phi(-L)$).
-For example, in the study of electron motion in thin metallic ”wires,” it is required that the wave function remains continuous at the junction $\nu=L$.
+For example, in the study of electron motion in thin metallic “wires,” it is required that the wave function remains continuous at the junction $\nu=L$.
 These boundary conditions have a clear physical interpretation, representing the conservation of probability flow at the junction (see [30]).
 This physical requirement can be formally expressed through the following family of self-adjoint operator extensions $(-\Delta,D_{Z,N})_{Z\in\mathbb{R}}$ for $(\mathcal{H}_{0},D(\mathcal{H}_{0}))$ with
-$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdot\cdot\cdot=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
-where for any $n\geqq 0$,
+$\begin{array}[]{ll}D_{Z,N}=\{\mathbf{U}\in H^{2}(\mathcal{G}):&\phi(L)=\phi(-L)=\psi_{1}(L)=\cdots=\psi_{N}(L)\;\text{and}\\ &\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{j=1}^{N}\psi_{j}^{\prime}(L+)+Z\psi_{1}(L)\},\end{array}$ (1.5)
+where for any $n\geqslant 0$,
 $H^{n}(\mathcal{G})=H^{n}(-L,L)\oplus\bigoplus_{j=1}^{N}H^{n}(L,+\infty).$
-The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann-Kirchhoff type if $Z=0$.
-The parameter $Z$ is a coupling constant between the loop and the several half-lines (see Example 7 and Remark 1).
-The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [27], [29], and references therein).
-It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 6, 7, 16, 21, 22, 34, 42, 41], and references therein).
-Our second focus in this manuscript is associated to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
-$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t),\quad x\neq 0,t\in\mathbb{R},\textbf{e}\in E,$ (1.6)
-with real coefficients given by the sequences $(\alpha_{\textbf{e}})_{e\in E}$, $(\beta_{\textbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping edge graph (see Theorems 4.2, 4.3 and 4.7).
-Our approach is based on the abstract results in [44] applied to the case of the Airy-operator
+The boundary conditions in (1.5) are called $\delta$-type if $Z\neq 0$, and Neumann–Kirchhoff type if $Z=0$.
+The parameter $Z$ is a coupling constant between the loop and the several half-lines.
+The choice of coupling at the vertex $\nu=L$ corresponds to a feasible quantum-wire experiment (see [29, 31] and references therein).
+It should be noted that, based on the domain $D_{Z,N}$, several recent studies have been conducted, particularly on the existence and stability of standing-wave solutions for the model in (1.1) (see [3, 4, 5, 7, 8, 18, 23, 24, 37, 46, 44] and references therein).
+A distinctive feature of the approach developed here is that it allows one to impose vertex conditions a priori and then identify all extensions compatible with them, rather than parametrizing all extensions at once and then selecting the relevant subfamily.
+Concrete illustrations are provided in Examples 8 and 10 below.
+The existence and (in)stability of standing wave profiles for the NLS model (1.1) on looping-edge and $\mathcal{T}$-shaped graphs under dynamics induced by some of these extensions is an active research direction pursued by the authors.
+As a further illustration of this framework, in Section 5.7 we discuss the orbital instability of elementary (yet novel) standing waves for the cubic NLS on $\mathcal{G}$ under pure periodic plus Neumann type of interactions.
+Another important goal of this work is to obtain a characterization of all boundary conditions under which the Airy-type evolution equation
+$\partial_{t}u_{\mathbf{e}}(x,t)=\alpha_{\mathbf{e}}\partial_{x}^{3}u_{\mathbf{e}}(x,t)+\beta_{\mathbf{e}}\partial_{x}u_{\mathbf{e}}(x,t),\quad x\neq 0,\ t\in\mathbb{R},\ \mathbf{e}\in E,$ (1.6)
+with real coefficients given by the sequences $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$, generates either a contraction semigroup or a unitary group on a looping-edge graph (see Theorems 4.2, 4.3 and 4.6).
+Our approach is based on the abstract results in [49] applied to the case of the Airy operator
 $A_{0}:\{u_{e}\}_{e\in E}\mapsto\{\alpha_{e}\partial_{x}^{3}u_{e}+\beta_{e}\partial_{x}u_{e}\}_{e\in E},$ (1.7)
-when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and treating the point $-L$ as a terminal vertex (as in the case of a $\mathcal{T}$-graph).
-Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary.
-Thus, we get in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$.
-For instance, in Example 1 below we get $\delta$-interactions type boundary condictions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [8]) under conditions on $\alpha_{\textbf{e}}$ and $\beta_{\textbf{e}}$.
-Also, we establish conditions for obtaining extensions of $A_{0}$ generating contractile dynamics.
-We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg-de Vries (KdV) equation
-$\partial_{t}u_{\textbf{e}}(x,t)=\alpha_{\textbf{e}}\partial_{x}^{3}u_{\textbf{e}}(x,t)+\beta_{\textbf{e}}\partial_{x}u_{\textbf{e}}(x,t)+2u_{\textbf{e}}(x,t)\partial_{x}u_{\textbf{e}}(x,t),$ (1.8)
+when we consider the metric graph $\mathcal{G}$ represented by $E$ defined above and virtually treating the point L-L as a terminal vertex.
+In contrast with the symmetric operator case $\mathcal{H}_{0}$ the skew-symmetry of $A_{0}$ induce a slightly different boundary system and therefore different structure of the characterizations (compare Theorems 5.2 and 4.3 below) Therefore, notions of self-orthogonal subspaces and operators based on indefinite inner product spaces (Krein spaces) will be necessary.
+Thus, we obtain in particular a description of all (skew-)self-adjoint realizations of $A_{0}$, i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$.
+For instance, in Example 1 below we obtain $\delta$-interaction type boundary conditions on a looping-edge graph in a similar way as done in the case of metric star-shaped graphs (see, for instance, Theorem 3.6 in [10]) under conditions on $\alpha_{\mathbf{e}}$ and $\beta_{\mathbf{e}}$.
+We also establish conditions for obtaining extensions of $A_{0}$ generating contractive dynamics.
+We recall that the evolution linear model in (1.6) is associated to the following vectorial Korteweg–de Vries (KdV) equation
+$\partial_{t}u_{\mathbf{e}}(x,t)=\alpha_{\mathbf{e}}\partial_{x}^{3}u_{\mathbf{e}}(x,t)+\beta_{\mathbf{e}}\partial_{x}u_{\mathbf{e}}(x,t)+2u_{\mathbf{e}}(x,t)\partial_{x}u_{\mathbf{e}}(x,t)$ (1.8)
 on the metric graph $\mathcal{G}$.
-Recently, Mugnolo, Noja, and Seifert [38] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively.
-The half-lines are connected at a unique vertex $\nu=0$.
+Recently, Mugnolo, Noja, and Seifert [41] characterized all boundary conditions under which the Airy equation in (1.6) generates either a semigroup or a group on a metric graph that has a structure represented by the set $E\equiv E_{-}\cup E_{+}$, where ${E}_{+}$ and ${E}_{-}$ are finite or countable collections of semi-infinite edges $e$ parametrized by $(-\infty,0)$ and $(0,+\infty)$, respectively, connected at a unique vertex $\nu=0$.
 In this case the graph is sometimes called a star-shaped metric graph.
@@
 We recall that when the star-shaped metric graph is balanced ($|{E}_{+}|=|{E}_{-}|$), the extension theory implies that there is a $9|{E}_{+}|^{2}$-parameter family of skew-self-adjoint operators for $A_{0}$.
-Now, on looping-edge graph when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we get that there must be $9(k+1)^{2}$ family of unitary groups, each one having its own representation.
+On a looping-edge graph, when the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, there must be a $9(k+1)^{2}$-parameter family of unitary groups, each one having its own representation.
 As far as we know, the study of the KdV equation on metric graphs remains relatively underdeveloped.
@@
 More recently, Mugnolo, Noja, and Seifert [42] investigated the existence of solitary waves on a star-shaped metric graph in terms of the coefficients of the KdV equation $(\alpha_{\mathbf{e}})_{e\in E}$, $(\beta_{\mathbf{e}})_{e\in E}$ on each edge, with coupling conditions at the vertex $\nu=0$, which may or may not involve continuity, and examined the speeds of the traveling waves.
-Lastly, Angulo and Cavalcante [10] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$- interaction type.
-They also extended the linear instability result from [8] to establish nonlinear instability.
-Results of the dynamics of the KdV model on looping edge graphs or on $\mathcal{T}$-shaped graphs, such as local well-posedness of the Cauchy problem, existence and stability of solitary waves, etc., are not known in the literature to the best of the authors knowledge.
-We believe that the theory established in this manuscript can be considered as a first step towards a better understanding of this notable dispersive model on metric graphs.
+Lastly, Angulo and Cavalcante [12] developed a local well-posedness theory for the KdV on a balanced star graph, considering boundary conditions at the vertex $\nu=0$ of $\delta$-interaction type, and extended the linear instability result from [10] to establish nonlinear instability.
+Results on the dynamics of the KdV model on looping-edge graphs, such as local well-posedness of the Cauchy problem and existence and stability of solitary waves, are not known in the literature to the best of the authors’ knowledge.
+The theory developed in this manuscript is intended to fill that gap at the linear level and to lay the groundwork for future nonlinear investigations.
+Several applications are already emerging from the framework presented here.
+On the one hand, the boundary system based on separation of first derivatives, developed in Section 4.3, has already proven effective in the analysis of dispersive equations on metric graphs: the technique was used in [42] to construct explicit solitary-wave solutions for the KdV equation on star graphs, where the first-derivative coupling at the vertex plays a distinguished role.
+Moreover, the domains generated via separation of derivatives in Example 7 directly recover the looping-edge analogue to the vertex conditions employed in [6, 48], works on the stabilization and controllability of the KdV on networks with bounded and unbounded domains.
+In general these abstract theory can be used to yield the class of boundary conditions needed for controllability problems on metric graphs, in fact the class of extensions $A_{Y,R_{1}}$ constitute the functional-analytic backbone of an ongoing program on exponential stabilization and controllability of the KdV equation on $\mathcal{G}$, currently in an advanced stage of research.
+As a concrete illustration of the use of contractive dynamics, in Section 4.4 we show that adding a damping term $\gamma\mathbf{U}$ to the Airy evolution under the boundary conditions of Example 6 yields exponential decay of the $L^{2}(\mathcal{G})$-energy at rate $\gamma$, with the dissipation arising entirely from the boundary structure rather than from the dispersive coefficients.
 Organization of the article.
-  1. 2.
-    Preliminaries
-    1. 2.1.
-      Notation
-    2. 2.2.
-      Stone and Lummer-Phillips Theorems
-    3. 2.3.
-      Operators on Krein Spaces
-  2. 3.
-    The free Airy operator and deficiency indices
-  3. 4.
-    The Airy operator on a looping-edge graph
-    1. 4.1.
-      Boundary system for even amount of half-lines
-      1. 4.1.1.
-        Extensions generating unitary dynamics
-      2. 4.1.2.
-        Examples: $\delta$-type interactions on a looping-edge graph
-    2. 4.2.
-      Boundary system for arbitrary amount of half-lines
-      1. 4.2.1
-        Extensions generating contractile dynamics
-      2. 4.2.2
-        Examples: $\delta$-type interactions on a tadpole graph
-    3. 4.3.
-      Boundary system separating derivatives
-      1. 4.3.1
-        Extensions generating contractile dynamics via splitting derivatives
-      2. 4.3.2
-        Examples
-  4. 5.
-    The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped metric graphs.
-    1. 5.1.
-      Deficiency indices for the free Schrödinger operator
-    2. 5.2.
-      Boundary system
-    3. 5.3.
-      Extensions generating unitary dynamics
-    4. 5.4.
-      Extensions generating unitary dynamics parameterized by linear subspaces
-    5. 5.5.
-      Extensions to $\mathcal{T}$-shaped metric graphs
+Section 2 collects the necessary preliminaries, including notation, the Stone and Lummer–Phillips theorems, and the relevant theory of operators on Krein spaces.
+Section 3 analyzes the free Airy operator and computes its deficiency indices on the looping-edge graph.
+Section 4 is the core of the paper and is devoted to the Airy operator on looping-edge graphs.
+It is divided into three parts: Section 4.1 treats the case of an even number of half-lines and characterizes extensions generating unitary dynamics; Section 4.2 handles an arbitrary number of half-lines and characterizes extensions generating contractive dynamics; Section 4.3 develops an alternative boundary system based on the separation of derivatives, providing a further class of extensions with contractive dynamics.
+Finally, Section 5 applies the same abstract framework to the Schrödinger operator on looping-edge graphs, characterizing its self-adjoint extensions and the associated unitary dynamics.
 2. Preliminaries
@@
 For $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$, we denote the vector of values of U and its derivatives at the vertex with
-$\vec{\textbf{U}}:=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\ \vec{\textbf{U}}^{\prime}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mbox{ and }\vec{\textbf{U}}^{\prime\prime}:=\begin{pmatrix}\phi^{\prime\prime}(-L)\\ \phi^{\prime\prime}(L)\\ \psi_{1}^{\prime\prime}(L)\\ \vdots\\ \psi_{N}^{\prime\prime}(L)\end{pmatrix}.$
+$\begin{split}\ &\vec{\mathbf{U}}:=\vec{\mathbf{U}}(L)=(\phi(-L),\ \phi(L),\ \psi_{1}(L),\ \dots,\psi_{N}(L))^{T},\\ \ &\vec{\mathbf{U}}^{\prime}:=\vec{\mathbf{U}}^{\prime}(L):=(\phi^{\prime}(-L),\ \phi^{\prime}(L),\ \psi_{1}^{\prime}(L),\ \dots,\psi_{N}^{\prime}(L))^{T},\\ \ &\vec{\mathbf{U}}^{\prime\prime}:=\vec{\mathbf{U}}^{\prime\prime}(L):=(\phi^{\prime\prime}(-L),\ \phi^{\prime\prime}(L),\ \psi_{1}^{\prime\prime}(L),\ \dots,\psi_{N}^{\prime\prime}(L))^{T}.\end{split}$
 For the components of U, we denote the derivatives values at the vertex with
@@
 For convenience we denote for $\psi=(\psi_{1},\dots,\psi_{N})$
-$\partial\psi(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{2}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix},\ \partial\psi_{even}(L):=\begin{pmatrix}\partial\psi_{2}(L)\\ \partial\psi_{4}(L)\\ \vdots\\ \partial\psi_{2k}(L)\end{pmatrix}\mbox{ and }\partial\psi_{odd}(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{3}(L)\\ \vdots\\ \partial\psi_{2k-1}(L)\end{pmatrix}.$ (2.1)
+$\partial\psi(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{2}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix},\ \partial\psi_{even}(L):=\begin{pmatrix}\partial\psi_{2}(L)\\ \partial\psi_{4}(L)\\ \vdots\\ \partial\psi_{2k}(L)\end{pmatrix}\mbox{ and }\partial\psi_{odd}(L):=\begin{pmatrix}\partial\psi_{1}(L)\\ \partial\psi_{3}(L)\\ \vdots\\ \partial\psi_{2k-1}(L)\end{pmatrix}$.
+(2.1)
 2.2. Stone and Lumer-Phillips Theorems
@@
     If $\mathcal{L}$ is densely defined, the $(\mathcal{K}_{2},\mathcal{K}_{1})$-adjoint of $\mathcal{L}$ is defined by
-    $D(\mathcal{L}^{\#}):=\{y\in\mathcal{K}_{2}\mid\exists z\in\mathcal{K}_{1}:\left(\mathcal{L}x\mid y\right)_{\mathcal{K}_{2}}=\left(x\mid z\right)_{\mathcal{K}_{1}}\mbox{ for all }x\in D(\mathcal{L})\}\mbox{ and }\mathcal{L}^{\#}y=z.$ (2.5)
+[…]
   3. (3)
-    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and #$\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
+    $\mathcal{L}$ is called $(\mathcal{K}_{1},\mathcal{K}_{2})$-unitary if $D(\mathcal{L})$ and $R(\mathcal{L})$ are dense, $\mathcal{L}$ is injective and $\mathcal{L}^{\#}=\mathcal{L}^{-1}$.
 3. The free Airy operator and deficiency indices
@@
 where $\alpha_{j}$ and $\beta_{j}$ are real numbers for any $j$.
-For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
+For $\mathcal{G}$ determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering L-L as a terminal vertex, we understand the action of $A_{0}$ over the edges in a component-wise manner, that is, for $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ posed on $\mathcal{G}$ we denote
 $A_{0}\textbf{U}=\left(\alpha_{0}\partial_{x}^{3}\phi+\beta_{0}\partial_{x}\phi,\ \alpha_{1}\partial_{x}^{3}\psi_{1}+\beta_{1}\partial_{x}\psi_{1},\ \dots\ ,\ \alpha_{N}\partial_{x}^{3}\psi_{N}+\beta_{N}\partial_{x}\psi_{N}\right).$ (3.2)
@@
 For $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$, we denote the inner product on $L^{2}(\mathcal{G})$ with
-[…]
+$\textbf{U},\textbf{V}]:=[(\phi,\psi_{1},\dots,\psi_{N}),(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})]_{L^{2}(\mathcal{G})}=\int_{-L}^{L}\phi\overline{\tilde{\phi}}dx+\sum_{j=1}^{N}\int_{L}^{\infty}\psi_{j}\overline{\tilde{\psi}_{j}}dx.$ (3.5)
 We consider analogously the Sobolev spaces
@@
 with deficiency indices $d_{\mp}=\mbox{dim}\mathcal{D}_{\mp}(iA_{0}^{*})$.
-From the definition of $A_{0}$ in (3.1) we have that solutions of $-i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
+From the definition of $A_{0}$ in (3.1) we have that solutions of $i\alpha_{e}\partial_{x}^{3}u_{e}-i\beta_{e}\partial_{x}u_{e}\pm iu_{e}=0$ without any boundary condition are obtained as linear combinations of complex exponentials of the form
 $u_{e}=c_{1}e^{r_{1}x_{e}}+c_{2}e^{r_{2}x_{e}}+c_{3}e^{r_{3}x_{e}}$
@@
 We note that $e^{rx}$ belong to $L^{2}(-L,L)$ independent on $r$, which means $d_{-_{0}}=d_{+_{0}}=3$ along $e_{0}$.
-On the other hand, on a half-line $[L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$ and $\beta_{e}$.
-We summarize the behaviour (on a single edge $[L,\infty)$) in the following table.
-Signs of $\alpha_{e}$ and $\beta_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
-$\alpha_{e}>0,\beta_{e}\geqslant 0$ 2 1
-$\alpha_{e}>0,\beta_{e}<0$ 2 1
-$\alpha_{e}<0,\beta_{e}\geqslant 0$ 1 2
-$\alpha_{e}<0,\beta_{e}<0$ 1 2
+On the other hand, on a half-line $L,\infty)$ we note $e^{rx}$ is square integrable only if the real part of $r$ is negative, which depends on the signs of $\alpha_{e}$.
+We summarize the behavior (on a single edge $L,\infty)$) in the following table.
+Sign of $\alpha_{e}$ $d_{-_{e}}$ $d_{+_{e}}$
+$\alpha_{e}>0$ 2 1
+$\alpha_{e}<0$ 1 2
 From the Von Neumann and Krein theory, one can expect self-adjoint extensions of $iA_{0}$ only if the deficiency indices agree, that is, $d_{+}=d_{-}$.
@@
 Several combinations of signs of $\alpha_{e}$ and $\beta_{e}$ along the edges of $\mathcal{G}$ may lead to equal deficiency indices.
-For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $33k+$ 3.
+For instance, in case the number of attached half-lines is $N=2k$ for some integer $k$ with the conditions $\alpha_{j},\beta_{j}>0$ if $j$ is odd and $\alpha_{j},\beta_{j}<0$ if $j$ is even, we would get deficiency indices $d_{+}$ and $d_{-}$ equal to $3k+3$.
 Thus, in this case, we know that there must be a $9(k+1)^{2}$ family of unitary groups for the case of the looping-edge graph, each one having its own representation.
@@
 Let us first consider $\mathcal{G}$ consisting of an even amount $2k$ of half-lines attached to the vertex in which exactly half of them have a leading coefficient $\alpha_{e}>0$.
-More precisely, we consider $\mathcal{G}$ defined trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if ±$\pm\alpha_{e}>0$.
+More precisely, we consider $\mathcal{G}$ identified trough $V=\{L,-L\}$ and $E=\{[-L,L]\}\cup E_{+}\cup E_{-}$ where $|E_{-}|=|E_{+}|=k\in\mathbb{Z}^{+}$ and $e\in E_{\pm}$ if and only if $\pm\alpha_{e}>0$.
 Our objective in this section is to provide a first description of the boundary values of $A_{0}^{*}=-A_{0}$ ($D(A_{0}^{*})=H^{3}(\mathcal{G})$), we proceed as in Section 2 of [49].
@@
 […]
-with [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
+with $\cdot$$\cdot$] being the $L^{2}(\mathcal{G})$-inner product in (3.5).
 Proposition 4.1.
@@
 $B_{j}:=\begin{pmatrix}\beta_{j}&0&\alpha_{j}\\ 0&-\alpha_{j}&0\\ \alpha_{j}&0&0\end{pmatrix},\ j=0,\dots,N,$ (4.3)
-[…]
+and $(\cdot\mid\cdot)$ represents the standard inner product in $\mathbb{C}^{3}$.
 Proof.
@@
 Next, we define the boundary-subspaces $\mathcal{B}_{\pm}$ as
-+=$\mathcal{B}_{+}=\Big{\{}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}$ (4.4)
+$\mathcal{B}_{+}=\Big\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big\}$ (4.4)
 and
-=$\mathcal{B}_{-}=\Big{\{}\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big{\}}.$ (4.5)
+$\mathcal{B}_{-}=\Big\{\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}:(\phi,\psi_{1}\dots,\psi_{N})\in H^{3}(\mathcal{G})\Big\}.$ (4.5)
 In this form, we have the following definition.
@@
 Consider the spaces $\mathcal{B}_{+}\cong\mathcal{B}_{-}\cong\mathbb{C}^{3(k+1)}$ in (4.4)-(4.5) and endow these spaces with the non-degenerated indefinite inner products
-$\left(x\mid y\right)_{\mathcal{B}_{+}}:=\left(x\mid y\right)_{+}=\left(B_{+}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{\mathcal{B}_{-}}:=\left(x\mid y\right)_{-}:=\left(B_{-}x\mid y\right),$ (4.6)
 […]
-$B_{+}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{3}\\ &&&\ddots\\ &&&&B_{2k-1}\end{pmatrix}\mbox{ and }B_{-}:=\begin{pmatrix}B_{0}\\ &-B_{2}\\ &&-B_{4}\\ &&&\ddots\\ &&&&-B_{2k}\end{pmatrix}.$
-$(\mathcal{B}_{\pm},\left(\cdot\mid\cdot\right)_{\pm})$ are Krein spaces.
+where $(\cdot\mid\cdot)$ represents the standard inner product in $\mathbb{C}^{3(k+1)}$ and
+$B_{+}:=\operatorname{diag}(B_{0},B_{1},B_{3},\dots,B_{2k-1})\ \mbox{ and }\ B_{-}:=\operatorname{diag}(B_{0},-B_{2},-B_{4},\dots,-B_{2k}).$
+(±,()±)($\mathcal{B}_{\pm}$$\left$($\cdot\mid\cdot\right$$_{\pm})$ are Krein spaces.
 Under these new notations, Proposition 4.1 can be rewritten in a compact form as
-$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}.$ (4.7)
+$A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{odd}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}_{odd}(L)\end{pmatrix}\right)_{+}-\left(\begin{pmatrix}\partial\phi(L)\\ \partial\psi_{even}(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \partial\tilde{\psi}_{even}(L)\end{pmatrix}\right)_{-}$.
+(4.7)
 Definition 4.3.
@@
 and the sesquilinear form
-$\begin{split}w:&\ (\mathcal{B}_{+}\oplus\mathcal{B}_{-})\times(\mathcal{B}_{+}\oplus\mathcal{B}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\left(x\mid\tilde{x}\right)_{+}-\left(y\mid\tilde{y}\right)_{-}.\end{split}$ (4.9)
+[…]
 Note from (4.1) and (4.2) that
 $\Omega((\textbf{U},A_{0}^{*}\textbf{U}),(\textbf{V},A_{0}^{*}\textbf{V}))=w(F(\textbf{U},A_{0}^{*}\textbf{U}),F(\textbf{V},A_{0}^{*}\textbf{V})).$ (4.10)
-By [44] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ and therefore self-adjoint extensions of $iA_{0}$.
+By [49] we can now characterize skew-self-adjoint extensions $A$ of $A_{0}$ - i.e., skew-self-adjoint restrictions of $A_{0}^{*}$ Ð and therefore self-adjoint extensions of $iA_{0}$.
 Note that obviously $G(A)=\{(x,Ax):x\in D(A)\}\subset G(A_{0}^{*})$
@@
 Theorem 4.3.
-Let $X\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}$.
-$X$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
+Let $\subset\mathcal{B}_{+}\oplus\mathcal{B}_{-}. XX$ is $w$-self-orthogonal if and only if there is a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator $\mathcal{L}$ such that $X=G(\mathcal{L})$.
 Proof.
@@
 A symmetric argument shows the image $R(\mathcal{L})$ is dense.
-Let us now see #$\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
+Let us now see $\mathcal{L}^{\#}=L^{-1}$. First note for $x\in D(\mathcal{L})$ and $y\in R(\mathcal{L})$, from (4.11), we have
 $(\mathcal{L}x\mid y)_{-}=(x\mid\mathcal{L}^{-1}y)_{+}.$
-The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$.
-On the other hand, if $y\in D(\mathcal{L}^{\#})$ with #$\mathcal{L}^{\#}y=x$ we have
-$(\mathcal{L}z\mid y)_{-}=(z\mid x)_{+}\ \ \mbox{for all }z\in D(\mathcal{L}).$
+The latter implies $R(\mathcal{L})\subset D(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}y=\mathcal{L}^{-1}y$; i.e., $\mathcal{L}^{-1}\subset\mathcal{L}^{\#}$.
+On the other hand, if $y\in D(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}y=x$ we have
+[…]
 Note the latter is equivalent to say
-$w((z,\mathcal{L}z),(x,y))=0\ \mbox{ for all }z\in D(\mathcal{L}),$
+[…]
 which implies $(x,y)\in X^{\perp_{w}}=X=G(\mathcal{L})$. Thus, $(y,x)\in G(\mathcal{L}^{-1})$. We conclude $\mathcal{L}^{-1}=\mathcal{L}^{\#}$; proving $\mathcal{L}$ is unitary.
@@
 Example 1 ($\delta$-type interactions on a looping-edge graph).
-Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
+Let us first consider $k=1$, i.e., two half-lines attached to a common vertex. Assume $\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $\beta_{2}=\beta_{1}=\beta_{0}$. In particular,
 $A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right),$ (4.13)
@@
 $\mathcal{L}:=\begin{pmatrix}1&0&0&0&0&0\\ z&1&0&0&0&0\\ \frac{z^{2}}{2}&z&1&m&0&0\\ 0&0&0&1&0&0\\ 0&0&0&z&1&0\\ -m&0&0&\frac{z^{2}}{2}&z&1\end{pmatrix}$. (4.14)
-Since #$\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator.
+Since $\mathcal{L}^{\#}=B_{+}^{-1}\mathcal{L}^{*}B_{-}=\mathcal{L}^{-1}$, it can be seen that $\mathcal{L}$ defines a $(\mathcal{B}_{+},\mathcal{B}_{-})$-unitary operator.
 According to Proposition 4.2 and Theorem 4.3, $\mathcal{L}$ defines a skew-self-adjoint extension $A_{\mathcal{L}}$ of $A_{0}$ defined via the relation $F(G(A_{\mathcal{L}}))=G(\mathcal{L})$.
@@
 $\begin{split}&\phi(-L)=\phi(L)=\psi_{1}(L)=\psi_{2}(L),\\ &\left(\phi^{\prime}(L)+\psi^{\prime}_{2}(L)\right)-\left(\phi^{\prime}(-L)+\psi^{\prime}_{1}(L)\right)=2z\phi(-L).\end{split}$ (4.16)
-$-L$
-$L$
-$\phi$
-$\psi_{1}$
-$\psi_{2}$
-$-L$
-$L$
-$\phi$
-$\psi_{1}$
-$\psi_{2}$
-Figure 2. Examples of profiles satisfying (4.15) (on the left) and (4.16) (on the right).
 More generally, consider $k\in\mathbb{Z}^{+}$ arbitrary (i.e.
 $2k$ half-lines attached to the common vertex $\nu=L$) and a parameter $z\in\mathbb{R}$.
-Under the analogous assumption $-\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $-\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
+Under the analogous assumption $\alpha_{2j}=\alpha_{2j-1}=\alpha_{0}$, $\beta_{2j}=\beta_{2j-1}=\beta_{0}$ for all $j\in\{1,\dots,k\}$, so we obtain in particular for $k=2$
 $A_{0}=\left(\alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x},\ \alpha_{0}\partial_{x}^{3}+\beta_{0}\partial_{x},\ -\alpha_{0}\partial_{x}^{3}-\beta_{0}\partial_{x}\right).$ (4.17)
 Define the following operator with diagonal matrix of blocks
-$\mathcal{L_{\delta_{Z}}}:=\begin{pmatrix}\delta_{z}\\ &\delta_{z}\\ &&\ddots\\ &&&\delta_{z}\end{pmatrix}\mbox{ where }\delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}.$ (4.18)
+$\mathcal{L_{\delta_{Z}}}:=\operatorname{diag}(\delta_{z},\delta_{z},\dots,\delta_{z})\ \mbox{ where }\ \delta_{z}:=\begin{pmatrix}1&0&0\\ z&1&0\\ \frac{z^{2}}{2}&z&1\end{pmatrix}$. (4.18)
 Note that each block $\delta_{z}$ is such that $B_{0}^{-1}\delta_{z}^{*}B_{0}=\delta_{z}^{-1}$ for any $z\in\mathbb{R}$ ($B_{0}$ in (4.2)) and since $\mathcal{L}_{\delta_{z}}^{\#}$ is a diagonal matrix of blocks of the form $B_{0}^{-1}\delta_{z}^{*}B_{0}$ it can be seen that $\mathcal{L}_{\delta_{z}}^{\#}=\mathcal{L}_{\delta_{z}}^{-1}$.
@@
 From Proposition 4.2 and Theorem 4.3, $\mathcal{L}_{\delta_{z}}$ defines a skew-self-adjoint extension $A_{\mathcal{L}_{\delta_{z}}}$ of $A_{0}$ with domain
-$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k.\right\}$ (4.19)
+$D_{\delta_{z}}:=\left\{\textbf{V}\in H^{3}(\mathcal{G})\ \mid\delta_{z}\partial\phi(-L)=\partial\phi(L)\mbox{ and }\delta_{z}\partial\psi_{2j-1}(L)=\partial\psi_{2j}(L),\ j=1,\dots,k\right\}.$ (4.19)
 Note from (4.19) we obtained ”pair-wise” the conditions in (4.15).
-More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at $-L$ and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
+More precisely, abusing notation to identify $\psi_{-1}$ with $\phi$ at L-L and $\psi_{0}$ with $\phi$ at $L$ we have for any $j=0,\dots,k$ that
 $\begin{split}&\psi_{2j}(L)=\psi_{2j-1}(L),\ \psi_{2j}^{\prime}(L)-\psi_{2j-1}^{\prime}(L)=z\psi_{2j}(L)\mbox{ and}\\ &\psi_{2j}^{\prime\prime}(L)-\psi_{2j-1}^{\prime\prime}(L)=z\psi_{2j-1}^{\prime}(L)+\frac{z^{2}}{2}\psi_{2j-1}(L).\end{split}$ (4.20)
@@
 For the following examples we assume $|E_{+}|=|E_{-}|=1$.
-From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $366^{2}=$ 36 parameters.
+From the deficiency indices in Section 3 and the Von Neumann and Krein theory we know the skew-self-adjoint extensions of $A_{0}$ are parametrized by $6^{2}=36$ parameters.
 We exhibit a $2$ and a $4$-parameter independent families of unitary operators that maintain the looping-edge structure (i.e.
@@
 is a unitary operator and therefore that it defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of the functions $v=(\phi,\psi_{1},\psi_{2})\in D(A_{0}^{*})$ such that
-$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(-L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(-L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
+$\begin{split}&\phi(-L)=\phi(L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(-L)=\psi^{\prime}_{2}(L),\ \phi^{\prime}(L)-\psi^{\prime}_{1}(L)=m_{1}\psi_{1}(L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=m_{2}\psi_{1}(L)\ \mbox{and }\psi_{2}^{\prime\prime}(L)-\psi_{2}^{\prime\prime}(L)=\frac{m_{1}^{2}}{2}\psi_{1}(L)+m_{1}\psi_{1}^{\prime}(L)-m_{2}\phi(-L).\end{split}$
 Example 3.
-Consider again $k=1$, $-\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $-\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
+Consider again $k=1$, $\alpha_{2}=\alpha_{1}=\alpha_{0}$ and $\beta_{2}=\beta_{1}=\beta_{0}$. It can be seen for $m_{1},\dots,m_{4}\in\mathbb{R}$ that
 $\mathcal{L}=\begin{pmatrix}1&0&0&0&0&0\\ m_{1}&0&0&m_{2}&1&0\\ \frac{m_{1}^{2}+1}{2}&1&1&m_{3}&m_{1}&0\\ 0&0&0&1&0&0\\ 1&1&0&m_{4}&0&0\\ m_{4}-m_{3}+m_{1}m_{2}&m_{4}&0&\frac{m_{2}^{2}+m_{4}^{2}}{2}&m_{2}&1\end{pmatrix}$ (4.23)
 defines a skew-self-adjoint extension $A_{\mathcal{L}}$ with domain $D(A_{\mathcal{L}})$ made of $v\in D(A_{0}^{*})$ such that
-$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$
+$\begin{split}&\phi(L)=\phi(-L),\ \psi_{1}(L)=\psi_{2}(L),\ \phi^{\prime}(L)-\psi_{1}^{\prime}(L)=m_{1}\phi(-L)+m_{2}\psi_{1}(L),\\ &\psi_{2}^{\prime}(L)-\phi^{\prime}(-L)=m_{4}\psi_{1}(L)+\phi(-L),\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)=\frac{m_{1}^{2}+1}{2}\phi(-L)+m_{3}\psi_{1}(L)+\phi^{\prime}(-L)+m_{1}\psi_{1}^{\prime}(L),\mbox{and}\\ &\psi_{2}^{\prime\prime}(L)-\psi_{1}^{\prime\prime}(L)=(m_{4}-m_{3}+m_{1}m_{2})\phi(-L)+\frac{m_{2}^{2}+m_{4}^{2}}{2}\psi_{1}(L)+m_{4}\phi^{\prime}(-L)+m_{2}\psi_{1}^{\prime}(L).\end{split}$
 4.2. Boundary system for arbitrary amount of half-lines
-For this second part we no longer seek for unitary dynamics.
-We feel confortable allowing the dynamics to be generated by a $C_{0}$-semigroup of contractions.
-To do so, we consider $\mathcal{G}$ consisting of finite amount $N$ of half-lines attached to the vertex $L$ with no restrictions on the signs of the coefficients of the Hamiltonian.
-We still treat $-L$ as a terminal vertex but we will be emphatic in whether the examples to be presented maintain or not the looping-edge structure.
-It is worth mentioning that most of the following results also apply to tadpole-like graphs.
+For this part we no longer seek unitary dynamics.
+We allow the dynamics to be generated by a $C_{0}$-semigroup of contractions.
+We consider $\mathcal{G}$ consisting of a finite amount $N$ of half-lines attached to the vertex $L$, with no restrictions on the signs of the coefficients of the Hamiltonian.
+We still virtually treat L-L as a terminal vertex.
 Definition 4.4.
-Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ and endow these spaces with the non-degenerated indefinite inner products
-$\left(x\mid y\right)_{1}:=\left(C_{1}x\mid y\right)\mbox{ and }\ \left(x\mid y\right)_{0}:=\left(C_{0}x\mid y\right),$ (4.24)
+Consider the spaces $\mathcal{C}_{1}\cong\mathcal{C}_{0}\cong\mathbb{C}^{3(N+1)}$ equipped with the non-degenerate indefinite inner products
+$(x\mid y)_{1}:=(C_{1}\,x\mid y),\qquad(x\mid y)_{0}:=(C_{0}\,x\mid y),$ (4.24)
 where
-$C_{1}:=\begin{pmatrix}B_{0}\\ &B_{1}\\ &&B_{2}\\ &&&\ddots\\ &&&&B_{N}\end{pmatrix}\mbox{ and }C_{0}:=\frac{1}{N+1}\begin{pmatrix}B_{0}\\ &B_{0}\\ &&B_{0}\\ &&&\ddots\\ &&&&B_{0}\end{pmatrix}.$
-$(\mathcal{C}_{1},\left(\cdot\mid\cdot\right)_{1})$ and $(\mathcal{C}_{0},\left(\cdot\mid\cdot\right)_{0})$ are Krein spaces.
-Under these new notations, Proposition 4.1 can be rewritten as
-$[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}.$ (4.25)
+$C_{1}:=\operatorname{diag}(B_{0},B_{1},B_{2},\dots,B_{N}),\qquad C_{0}:=\frac{1}{N+1}\operatorname{diag}(B_{0},B_{0},\dots,B_{0}).$
+Both $(\mathcal{C}_{1},(\cdot\mid\cdot)_{1})$ and $(\mathcal{C}_{0},(\cdot\mid\cdot)_{0})$ are Krein spaces.
+4.2.1. Boundary trace maps and the Green identity
+We introduce boundary trace maps $\Gamma_{1},\Gamma_{0}:D(A_{0}^{*})\to\mathbb{C}^{3(N+1)}$ by
+$\Gamma_{1}\mathbf{U}:=\begin{pmatrix}\partial\phi(-L)\\[2.0pt] \partial\psi_{1}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix}\in\mathcal{C}_{1},\qquad\Gamma_{0}\mathbf{U}:=\begin{pmatrix}\partial\phi(L)\\ \partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in\mathcal{C}_{0}.$ (4.25)
+In terms of these maps, the Green identity for $A_{0}^{*}$ reads
+$A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]=(\Gamma_{1}\mathbf{U}\mid\Gamma_{1}\mathbf{V})_{1}-(\Gamma_{0}\mathbf{U}\mid\Gamma_{0}\mathbf{V})_{0}.$ (4.26)
+4.2.2. Boundary relations, Krein adjoints, and contractive extensions
 Definition 4.5.
-We consider the function
-$\begin{split}\mathcal{F}:&\ G(A_{0}^{*})\to\mathcal{C}_{1}\oplus\mathcal{C}_{0}\\ &(\textbf{U},A_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (4.26)
-4.2.1. Extensions generating contractile dynamics
-Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be a densely defined operator.
-Consider an extension $A_{\mathcal{L}}$ of $A_{0}$ defined by the relation $G(A_{\mathcal{L}})=\mathcal{F}^{-1}(G(\mathcal{L}))$.
-In other words,
-$D(A_{\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\in D(\mathcal{L})\mbox{ and }\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right\}.$ (4.27)
-acting as $A_{\mathcal{L}}\textbf{U}:=-A_{0}^{*}\textbf{U}$.
-Let us determine the adjoint of operator of $A_{\mathcal{L}}$.
+A boundary relation is a linear subspace $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$. Its Krein adjoint $R^{[*]}\subset\mathcal{C}_{0}\times\mathcal{C}_{1}$ is defined by
+$(y,x)\in R^{[*]}\;\Longleftrightarrow\;(y_{1}\mid y)_{0}=(x_{1}\mid x)_{1}\quad\text{for all }(x_{1},y_{1})\in R.$ (4.27)
+The relation $R$ is called a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if
+[…]
+When $R=\operatorname{graph}(\mathcal{L})$ for a linear map $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$, the Krein adjoint satisfies $R^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with
+$\mathcal{L}^{\#}:=C_{1}^{-1}\mathcal{L}^{*}C_{0},$ (4.29)
+and contractivity of $R$ reduces to $\mathcal{L}^{*}C_{0}\mathcal{L}-C_{1}\preceq 0$.
+Remark 1.
+The multivalued part of $R$ is $\operatorname{mul}(R):=\{y:(0,y)\in R\}$.
+Whenever $\operatorname{mul}(R)\neq\{0\}$, the relation $R$ is not the graph of any operator, and no matrix $\mathcal{L}$ can encode the corresponding boundary conditions.
+Concretely, $(0,y)\in R$ imposes $\Gamma_{1}\mathbf{U}=0$, i.e.
+$\phi(-L)=\phi^{\prime}(-L)=\phi^{\prime\prime}(-L)=0,\quad\psi_{j}(L)=\psi_{j}^{\prime}(L)=\psi_{j}^{\prime\prime}(L)=0,\quad j=1,\dots,N,$
+while leaving $\Gamma_{0}\mathbf{U}$ unconstrained.
+These are Dirichlet-type conditions at the incoming edges; they are valid generators of contraction semigroups whenever the corresponding relation is contractive, but are inaccessible to the operator-based framework.
+Definition 4.6.
+Given a boundary relation $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$, define
+$A_{R}\mathbf{U}:=-A_{0}^{*}\mathbf{U},\qquad D(A_{R}):=\bigl\{\mathbf{U}\in D(A_{0}^{*}):(\Gamma_{1}\mathbf{U},\Gamma_{0}\mathbf{U})\in R\bigr\}.$ (4.30)
+Let us describe the adjoint of $A_{R}$.
 Proposition 4.4.
-For $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ densely defined we have
-$D(A_{\mathcal{L}}^{*}):=\left\{(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{0}^{*})\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ and }\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right\}$ (4.28)
-acting as $A_{\mathcal{L}}^{*}\textbf{U}=A_{0}^{*}\textbf{U}.$
+$A_{R}^{*}=-A_{R^{[*]}}$, where $A_{R^{[*]}}$ acts as $A_{0}^{*}$ with domain
+$D\!\left(A_{R^{[*]}}\right):=\bigl\{\mathbf{V}\in D(A_{0}^{*}):(\Gamma_{0}\mathbf{V},\Gamma_{1}\mathbf{V})\in R^{[*]}\bigr\}.$ (4.31)
 Proof.
-Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$, $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})\in D(A_{0}^{*})$.
-From (4.7) and the definition of $A_{\mathcal{L}}$ we have
-$\begin{split}[A_{\mathcal{L}}\textbf{U},\textbf{V}]&=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]+\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.29)
-Therefore, $\textbf{V}\in D(A_{\mathcal{L}}^{*})$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
-$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)_{0}=\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)_{1}.$ (4.30)
-Note from the definition of $A_{\mathcal{L}}$ that $\mathcal{F}(G(A_{\mathcal{L}}))=G(\mathcal{L})$; which means
-$D(\mathcal{L})=\left\{\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\textbf{U}\in D(A_{\mathcal{L}})\right\}.$ (4.31)
-We conclude (4.30) is equivalent to
-#$\mathcal{L}^{\#}\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}=\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}.$
-From Proposition 4.4 we see that knowing the particular expression defining $\mathcal{L}^{\#}$ is important to determine explicitly the domain of $A_{\mathcal{L}}^{*}$.
-Note in case $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ is given by a $3(N+1)\times 3(N+1)$ matrix, the $(\mathcal{C}_{0},\mathcal{C}_{1})$-adjoint $\mathcal{L}^{\#}$ is given by the expression
-#$\mathcal{L}^{\#}=C_{1}^{-1}\mathcal{L}^{*}C_{0}.$ (4.32)
-In fact, for $y\in\mathcal{C}_{0}$ the definition of $\mathcal{L}^{\#}$ imply that for all $x\in\mathcal{C}_{1}$ we have
-$\left(C_{0}\mathcal{L}x\mid y\right)=\left(\mathcal{L}x\mid y\right)_{0}=\left(x\mid\mathcal{L}^{\#}y\right)_{1}=\left(C_{1}x\mid\mathcal{L}^{\#}y\right)_{0}.$ (4.33)
-Since $C_{1}$ is real symmetric we obtain that for all $x\in\mathcal{C}_{1}$
-$\left(x\mid\ \left(\mathcal{L}^{*}C_{0}-C_{1}\mathcal{L}^{\#}\right)y\right)=0.$ (4.34)
-Since $y$ is arbitrary, it follows
-$\mathcal{L}^{*}C_{0}=C_{1}\mathcal{L}^{\#},$ (4.35)
-which implies (4.32)
-In the following two propositions we characterize when both, $A_{\mathcal{L}}$ and $A_{\mathcal{L}}^{*}$, are dissipative operators.
-We recall that dissipative operators on Hilbert spaces were briefly described on subsection 2.2.
+Let $\mathbf{U}\in D(A_{R})$ and $\mathbf{V}\in D(A_{0}^{*})$. Using (4.26) and $A_{R}\mathbf{U}=-A_{0}^{*}\mathbf{U}$,
+$A_{R}\mathbf{U},\mathbf{V}]=[\mathbf{U},A_{0}^{*}\mathbf{V}]+(\Gamma_{0}\mathbf{U}\mid\Gamma_{0}\mathbf{V})_{0}-(\Gamma_{1}\mathbf{U}\mid\Gamma_{1}\mathbf{V})_{1}.$ (4.32)
+Hence $\mathbf{V}\in D(A_{R}^{*})$ if and only if
+$(y_{1}\mid\Gamma_{0}\mathbf{V})_{0}=(x_{1}\mid\Gamma_{1}\mathbf{V})_{1}\quad\forall\,(x_{1},y_{1})\in R,$
+which by (4.27) is equivalent to $(\Gamma_{0}\mathbf{V},\Gamma_{1}\mathbf{V})\in R^{[*]}$. In this case $A_{R}^{*}\mathbf{V}=A_{0}^{*}\mathbf{V}=-(-A_{0}^{*}\mathbf{V})=-A_{R^{[*]}}\mathbf{V}$. ∎
+We now obtain dissipativity via contractivity in the following sense:
 Proposition 4.5.
-Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$. $A_{\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
+$A_{R}$ is dissipative if and only if $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
 Proof.
-Let $\textbf{U}=(\phi,\psi_{1},\dots,\psi_{N})\in D(A_{\mathcal{L}})$. From (4.7) and the definition of $A_{\mathcal{L}}$ we have
-$\begin{split}2\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+\overline{[A_{\mathcal{L}}\textbf{U},\textbf{U}]}=[A_{\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}-\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.\end{split}$ (4.36)
-Hence, $\mbox{Re}[A_{\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if for all $\textbf{U}\in D(A_{\mathcal{L}})$ it follows
-$\left(\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{0}\leqslant\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\right)_{1}.$ (4.37)
-From (4.31) we conclude that (4.37) is equivalent to $\mathcal{L}$ being a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction. ∎
+For $\mathbf{U}\in D(A_{R})$, set $(x,y):=(\Gamma_{1}\mathbf{U},\Gamma_{0}\mathbf{U})\in R$. The Green identity (4.26) applied to the pair $(\mathbf{U},\mathbf{U})$ gives
+$2\operatorname{Re}[A_{R}\mathbf{U},\mathbf{U}]=(y\mid y)_{0}-(x\mid x)_{1}.$ (4.33)
+Hence $\operatorname{Re}[A_{R}\mathbf{U},\mathbf{U}]\leqslant 0$ for all $\mathbf{U}\in D(A_{R})$ if and only if $(y\mid y)_{0}\leqslant(x\mid x)_{1}$ for all $(x,y)\in R$, which is exactly (4.28). ∎
+Remark 2.
+We note contractivity of $R$ implies contractivity of $R^{[*]}$ in this finite- dimensional setting.
+In fact, by the polar decomposition of an invertible Hermitian matrix we can write $C_{0}=J_{0}H_{0}$ and $C_{1}=J_{1}H_{1}$ with $H_{0},H_{1}$ positive definite and $J_{0},J_{1}$ fundamental symmetries of the respective Krein spaces.
+The contractivity condition (4.28) for $R$, when expressed through the similarity transforms $H_{0}^{1/2}$ and $H_{1}^{1/2}$, becomes a Hilbert-space contraction condition for an equivalent relation.
+Its adjoint in the Hilbert sense is again a contraction, and undoing the similarity transforms shows that $R^{[*]}$ satisfies (4.28) with the roles of $\mathcal{C}_{0}$ and $\mathcal{C}_{1}$ interchanged.
+Consequently the dissipativity of $A_{R}$ automatically entails the dissipativity of $A_{R}^{*}=-A_{R^{[*]}}$, without any additional hypothesis on $R$.
+Theorem 4.6.
+Let $R\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$ be a linear boundary relation.
+Then $A_{R}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ if and only if $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction.
+Proof.
+Suppose $R$ is contractive.
+By Proposition 4.5, $A_{R}$ is dissipative.
+By Remark 2, $R^{[*]}$ is also contractive, so Proposition 4.5 applied to $R^{[*]}$ gives that $A_{R}^{*}=-A_{R^{[*]}}$ is dissipative.
+The Lumer–Phillips theorem then yields that $A_{R}$ generates a contraction semigroup.
+Conversely, if $A_{R}$ generates a contraction semigroup, the Lumer–Phillips theorem implies $A_{R}$ is dissipative, and Proposition 4.5 gives that $R$ is contractive. ∎
+4.2.3. Examples
+We now present extensions $A_{R}$ generating semigroups of contractions, framed as boundary relations.
+The first example corresponds to an operator-type relation (graph of a matrix), while Remark 1 describes the additional Dirichlet-type extensions that fall outside the operator framework.
+Example 4 ($\delta$-type interactions on a tadpole graph).
+Let $N=1$, $\alpha_{0}=\alpha_{1}=\beta_{0}=\beta_{1}$, and let $m_{1},\dots,m_{4}\in\mathbb{R}$. Define the matrix
+$M:=\begin{pmatrix}1&0&0&0&0&0\\[2.0pt] \dfrac{m_{2}+m_{3}}{2}&1&0&1&0&0\\[6.0pt] m_{1}&m_{2}+m_{3}&2&m_{2}&0&0\\[2.0pt] 0&0&0&1&0&0\\[2.0pt] \dfrac{m_{2}+m_{3}}{2}&1&0&1&0&0\\[6.0pt] m_{3}&2&0&m_{4}&0&2\end{pmatrix}$, (4.34)
+and consider the boundary relation $R:=\operatorname{graph}(M)\subset\mathcal{C}_{1}\times\mathcal{C}_{0}$, i.e.
+$R=\bigl\{(x,Mx):x\in\mathcal{C}_{1}\bigr\}.$ (4.35)
+By Definition 4.5, $R$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if and only if the matrix $M^{*}C_{0}M-C_{1}$ is negative semidefinite.
+A direct computation shows that a sufficient condition is
+$\alpha_{0}<0,\qquad m_{4}\leqslant\tfrac{3}{2},\qquad 4m_{1}-(m_{2}^{2}+m_{3}^{2})\leqslant 2.$ (4.36)
+By Theorem 4.6, under (4.36) the operator $A_{R}$ generates a $C_{0}$-semigroup of contractions.
+The Krein adjoint relation is $R^{[*]}=\operatorname{graph}(M^{\#})$ with $M^{\#}=C_{1}^{-1}M^{*}C_{0}$, given explicitly by
+$M^{\#}=\begin{pmatrix}1&0&0&0&0&0\\[2.0pt] \dfrac{-m_{2}-m_{3}}{2}&\dfrac{1}{2}&0&-1&\dfrac{1}{2}&0\\[8.0pt] \dfrac{m_{1}-1}{2}&\dfrac{-(m_{2}+m_{3})}{4}&\dfrac{1}{2}&\dfrac{m_{3}}{2}&\dfrac{-(m_{2}+m_{3})}{4}&0\\[8.0pt] 0&0&0&1&0&0\\[2.0pt] 0&0&0&0&0&0\\[2.0pt] \dfrac{m_{2}}{2}&-\dfrac{1}{2}&0&\dfrac{m_{4}-1}{2}&-\dfrac{1}{2}&\dfrac{1}{2}\end{pmatrix}$, (4.37)
+which is also a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction under (4.36), in agreement with Remark 2.
+Explicitly, the domain of $A_{R}$ consists of all $(\phi,\psi)\in D(A_{0}^{*})$ satisfying $(\Gamma_{1}(\phi,\psi),\,\Gamma_{0}(\phi,\psi))\in R$, that is,
+$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\\[4.0pt] &\phi^{\prime}(L)-\phi^{\prime}(-L)=\left(\tfrac{m_{2}+m_{3}+2}{2}\right)\phi(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-2\psi^{\prime\prime}(L)=(m_{3}+m_{4})\,\psi(L)+2\,\phi^{\prime}(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-2\phi^{\prime\prime}(-L)=(m_{1}+m_{2})\,\phi(-L)+(m_{2}+m_{3})\,\phi^{\prime}(-L).\end{split}$ (4.38)
+In particular, setting $m_{4}=m_{1}=0$, $m_{3}=-2$, and writing $z=m_{2}/2\in\mathbb{R}$, the conditions (4.38) reduce to
+$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\qquad\phi^{\prime}(L)-\phi^{\prime}(-L)=z\,\phi(-L),\\[4.0pt] &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)-\psi^{\prime\prime}(L)=z\,\phi^{\prime}(-L)+(z-1)\,\phi(-L).\end{split}$ (4.39)
+The adjoint operator $A_{R}^{*}=-A_{R^{[*]}}$ has domain consisting of all $(\phi,\psi)\in D(A_{0}^{*})$ satisfying $(\Gamma_{0}(\phi,\psi),\,\Gamma_{1}(\phi,\psi))\in R^{[*]}$, i.e.
+$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\qquad\psi^{\prime}(L)=0,\\[4.0pt] &-\tfrac{m_{2}+m_{3}}{2}\,\phi(L)+\tfrac{1}{2}\,\phi^{\prime}(L)-\phi(-L)+\tfrac{1}{2}\,\psi^{\prime}(L)=\phi^{\prime}(-L),\\[4.0pt] &\tfrac{m_{1}-1}{2}\,\phi(L)-\tfrac{m_{2}+m_{3}}{4}\,\phi^{\prime}(L)+\tfrac{1}{2}\,\phi^{\prime\prime}(L)+\tfrac{m_{3}}{2}\,\phi(-L)-\tfrac{m_{2}+m_{3}}{4}\,\psi^{\prime}(L)=\phi^{\prime\prime}(-L),\\[4.0pt] &\tfrac{m_{2}}{2}\,\phi(L)-\tfrac{1}{2}\,\phi^{\prime}(L)+\tfrac{m_{4}-1}{2}\,\phi(-L)-\tfrac{1}{2}\,\psi^{\prime}(L)+\tfrac{1}{2}\,\phi^{\prime\prime}(L)=\psi^{\prime\prime}(L).\end{split}$ (4.40)
+Remark 3.
+A canonical example of extensions not reachable by linear operators is the partial Dirichlet relation
+$R_{j}:=\bigl\{(x,y)\in\mathcal{C}_{1}\times\mathcal{C}_{0}:x_{j}=0\bigr\},$ (4.41)
+where $x_{j}\in\mathbb{C}^{3}$ is the $j$-th block of $x$, which imposes $\partial\psi_{j}(L)=0$ (i.e.
+$\psi_{j}(L)=\psi_{j}^{\prime}(L)=\psi_{j}^{\prime\prime}(L)=0$) while leaving $\Gamma_{0}\mathbf{U}$ unconstrained.
+Such relations are contractive whenever the residual quadratic form $(y\mid y)_{0}$ is non-positive on $\{0\}\times\operatorname{mul}(R)$.
+No finite matrix $\mathcal{L}$ can encode these boundary conditions, since that would require $\mathcal{L}\cdot 0=y\neq 0$, which is impossible for a linear map.
+4.3. Boundary system separating derivatives
+A finer description of the boundary behavior is obtained by separating the quadratic form involving first derivatives from those involving the function values and second derivatives.
+The technique of separating the first-derivative component from the boundary system, developed in the present subsection, is not merely a formal device.
+It has already proven to be an effective tool in the qualitative analysis of dispersive equations on metric graphs: in [42], this separation of derivatives was exploited to construct solitary-wave solutions for the Korteweg–de Vries equation on star graphs.
+Furthermore, the class of extensions $A_{Y,R_{1}}$ introduced below, together with the contraction semigroups they generate, constitutes the functional-analytic backbone of an ongoing program aimed at the stabilization and controllability of the KdV equation on metric graphs.
+We recall the notation $\vec{\mathbf{U}}$, $\vec{\mathbf{U}}^{\prime}$, $\vec{\mathbf{U}}^{\prime\prime}$ introduced in Subsection 2.1.
+Lemma 4.7.
+For any $\mathbf{U}=(\phi,\psi_{1},\dots,\psi_{N})$ and $\mathbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$,
+$A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]=\bigl(D_{\beta}\vec{\mathbf{U}}\mid\vec{\mathbf{V}}\bigr)+\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\mid\vec{\mathbf{V}}\bigr)+\bigl(D_{\alpha}\vec{\mathbf{U}}\mid\vec{\mathbf{V}}^{\prime\prime}\bigr)-\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime}\mid\vec{\mathbf{V}}^{\prime}\bigr),$ (4.42)
+where
+$D_{\alpha}:=\operatorname{diag}(\alpha_{0},-\alpha_{0},\alpha_{1},\dots,\alpha_{N}),\qquad D_{\beta}:=\operatorname{diag}(\beta_{0},-\beta_{0},\beta_{1},\dots,\beta_{N}).$ (4.43)
+Definition 4.7.
+Consider the spaces $\mathcal{K}_{+}\cong\mathcal{K}_{-}\cong\mathbb{C}^{N+1}$ equipped with the weighted inner products
+$(x\mid y)_{\mathcal{K}_{+}}:=(D_{\alpha}^{+}\,x\mid y),\qquad(x\mid y)_{\mathcal{K}_{-}}:=(D_{\alpha}^{-}\,x\mid y),$ (4.44)
+where
+$D_{\alpha}^{+}:=\operatorname{diag}(\alpha_{0},\alpha_{1},\dots,\alpha_{N}),\qquad D_{\alpha}^{-}:=\frac{1}{N+1}\operatorname{diag}(\alpha_{0},\alpha_{0},\dots,\alpha_{0}).$ (4.45)
+Both $(\mathcal{K}_{\pm},(\cdot\mid\cdot)_{\mathcal{K}_{\pm}})$ are Krein spaces.
+The matrices $D_{\alpha}^{-}$ and $D_{\alpha}^{+}$ split $D_{\alpha}$: the former collects the coefficient of $\phi(L)$ and the latter collects all remaining coefficients.
+Specifically, (4.42) and the identity
+$\bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime}\mid\vec{\mathbf{V}}^{\prime}\bigr)=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}$ (4.46)
+yield the split Green identity
+$\begin{split}[A_{0}^{*}\mathbf{U},\mathbf{V}]+[\mathbf{U},A_{0}^{*}\mathbf{V}]&=\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)+\Bigl(\vec{\mathbf{U}}\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)\\ &\quad-\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}+\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\;\middle|\;\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.47)
+4.3.1. Boundary relations for the first-derivative part
+We introduce derivative trace maps $\Lambda_{+},\Lambda_{-}:D(A_{0}^{*})\to\mathcal{K}_{\pm}$ by
+$\Lambda_{+}\mathbf{U}:=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix},\qquad\Lambda_{-}\mathbf{U}:=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}$.
+(4.48)
+Definition 4.8.
+A derivative boundary relation is a linear subspace $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$. Its Krein adjoint $R_{1}^{[*]}\subset\mathcal{K}_{-}\times\mathcal{K}_{+}$ is defined by
+[…]
+The relation $R_{1}$ is called a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction if
+[…]
+When $R_{1}=\operatorname{graph}(\mathcal{L})$ for a linear map $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$, the Krein adjoint satisfies $R_{1}^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}:=(D_{\alpha}^{+})^{-1}\mathcal{L}^{*}D_{\alpha}^{-}$, and (4.50) reduces to $(D_{\alpha}^{+})^{-1}\mathcal{L}^{*}D_{\alpha}^{-}\mathcal{L}-I\preceq 0$.
+4.3.2. Extensions generating contractive dynamics via splitting derivatives
+Definition 4.9.
+Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace. Define
+$\begin{split}D(A_{Y,R_{1}}):=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\vec{\mathbf{U}}\in Y,\quad D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp},\\ &\text{and}\quad(\Lambda_{+}\mathbf{U},\,\Lambda_{-}\mathbf{U})\in R_{1}\Bigr\},\end{split}$ (4.51)
+with $A_{Y,R_{1}}\mathbf{U}:=-A_{0}^{*}\mathbf{U}$.
+The subspace condition $\vec{\mathbf{U}}\in Y$ and the orthogonality condition $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$ together govern how function values and second derivatives interact at the vertex.
+The relation condition $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$ governs the first-derivative coupling independently.
 Proposition 4.8.
-Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined. $A_{\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction.
+Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and $Y\subset\mathbb{C}^{N+2}$ a closed subspace.
+Then $A_{Y,R_{1}}^{*}=-A_{Y,R_{1}^{[*]}}$, where $A_{Y,R_{1}^{[*]}}$ acts as $A_{0}^{*}$ with domain
+$\begin{split}D\!\left(A_{Y,R_{1}^{[*]}}\right)=\Bigl\{\mathbf{V}\in D(A_{0}^{*})\;\Big|\;&\vec{\mathbf{V}}\in Y,\quad\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\in Y^{\perp},\\ &\text{and}\quad(\Lambda_{-}\mathbf{V},\,\Lambda_{+}\mathbf{V})\in R_{1}^{[*]}\Bigr\}.\end{split}$ (4.52)
 Proof.
-Let $\textbf{V}\in D(A_{\mathcal{L}}^{*})$. From (4.7) we have
-$\begin{split}2\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]\\ &=\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}-\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}.\end{split}$ (4.38)
-Hence, $\mbox{Re}[A_{\mathcal{L}}^{*}\textbf{V},\textbf{V}]\leqslant 0$ if and only if
-(#$\left(\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{1}\leqslant\left(\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)_{0}\ \mbox{ for all }\textbf{V}\in D(A_{\mathcal{L}}^{*}).$ (4.39)
-From the definition of $D(A_{\mathcal{L}}^{*})$ we conclude (4.39) is equivalent to $\mathcal{L}^{\#}$ being a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction. ∎
-From the previous two propositions above and the Lummer-Phillips Theorem (see Theorem 2.2) we can obtain the following characterization of extensions generating $C_{0}$-semigroups of contractions.
+Let $\mathbf{U}\in D(A_{Y,R_{1}})$ and $\mathbf{V}\in D(A_{0}^{*})$. Using (4.47) and $A_{Y,R_{1}}\mathbf{U}=-A_{0}^{*}\mathbf{U}$,
+$\begin{split}[A_{Y,R_{1}}\mathbf{U},\mathbf{V}]&=[\mathbf{U},A_{0}^{*}\mathbf{V}]-\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)-\Bigl(\vec{\mathbf{U}}\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)\\ &\quad+(\Lambda_{+}\mathbf{U}\mid\Lambda_{+}\mathbf{V})_{\mathcal{K}_{+}}-(\Lambda_{-}\mathbf{U}\mid\Lambda_{-}\mathbf{V})_{\mathcal{K}_{-}}.\end{split}$ (4.53)
+We need $A_{Y,R_{1}}\mathbf{U},\mathbf{V}]=[\mathbf{U},A_{Y,R_{1}}^{*}\mathbf{V}]$ for all $\mathbf{U}\in D(A_{Y,R_{1}})$.
+Taking $\mathbf{U}$ with $\vec{\mathbf{U}}=\vec{\mathbf{U}}^{\prime}=0$ gives $(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\mid\vec{\mathbf{V}})=0$ for all $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}\in Y^{\perp}$, hence $\vec{\mathbf{V}}\in Y$.
+For arbitrary $x\in Y$, taking $\mathbf{U}$ with $\vec{\mathbf{U}}=x$ and $\vec{\mathbf{U}}^{\prime}=0$, the condition $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$ and (4.53) yield
+$\Bigl(x\;\Big|\;\tfrac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\Bigr)=\Bigl(D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\tfrac{1}{2}D_{\beta}\vec{\mathbf{U}}\;\Big|\;\vec{\mathbf{V}}\Bigr)=0,$
+hence $\frac{1}{2}D_{\beta}\vec{\mathbf{V}}+D_{\alpha}\vec{\mathbf{V}}^{\prime\prime}\in Y^{\perp}$.
+At this point (4.53) reduces to the requirement that for all $\mathbf{U}\in D(A_{Y,R_{1}})$,
+$(\Lambda_{+}\mathbf{U}\mid\Lambda_{+}\mathbf{V})_{\mathcal{K}_{+}}=(\Lambda_{-}\mathbf{U}\mid\Lambda_{-}\mathbf{V})_{\mathcal{K}_{-}}.$
+Since $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$, the definition (4.49) of the Krein adjoint shows this is equivalent to $(\Lambda_{-}\mathbf{V},\Lambda_{+}\mathbf{V})\in R_{1}^{[*]}$. ∎
+We now obtain, once more, dissipativity via contractivity.
+Proposition 4.9.
+$A_{Y,R_{1}}$ is dissipative if and only if $R_{1}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
+Proof.
+For $\mathbf{U}\in D(A_{Y,R_{1}})$, set $(x,y):=(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$.
+The split Green identity (4.47) applied to $(\mathbf{U},\mathbf{U})$, together with the conditions $\vec{\mathbf{U}}\in Y$ and $D_{\alpha}\vec{\mathbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\mathbf{U}}\in Y^{\perp}$, gives
+$2\operatorname{Re}[A_{Y,R_{1}}\mathbf{U},\mathbf{U}]=(x\mid x)_{\mathcal{K}_{+}}-(y\mid y)_{\mathcal{K}_{-}}.$ (4.54)
+Hence $\operatorname{Re}[A_{Y,R_{1}}\mathbf{U},\mathbf{U}]\leqslant 0$ for all $\mathbf{U}\in D(A_{Y,R_{1}})$ if and only if $(y\mid y)_{\mathcal{K}_{-}}\leqslant(x\mid x)_{\mathcal{K}_{+}}$ for all $(x,y)\in R_{1}$, i.e.
+$R_{1}$ is contractive.
+∎
+Remark 4.
+The same Hilbert-space similarity argument as in Remark 2 shows that contractivity of $R_{1}$ automatically implies contractivity of $R_{1}^{[*]}$.
+Hence dissipativity of $A_{Y,R_{1}}$ entails dissipativity of $A_{Y,R_{1}}^{*}$.
 Theorem 4.10.
-Let $\mathcal{L}:\mathcal{C}_{1}\to\mathcal{C}_{0}$ be densely defined.
-$A_{\mathcal{L}}$ is the generator of a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39).
+Let $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a linear boundary relation and $Y\subset\mathbb{C}^{N+2}$ a closed subspace.
+Then $A_{Y,R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ if and only if $R_{1}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
 Proof.
-Suppose $A_{\mathcal{L}}$ generates a semigroup $\{T(t)\}_{t\geqslant 0}$ of contractions.
-By Proposition 2.1, $A_{\mathcal{L}}$ is dissipative.
-Since $A_{\mathcal{L}}^{*}$ generates the semigroup $\{T(t)^{*}\}_{t\geqslant 0}$ which is also a contraction semigroup, from Proposition 2.1 we conclude $A_{\mathcal{L}}^{*}$ is dissipative.
-From Propositions 4.5 and 4.6 we conclude that $\mathcal{L}$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction and $\mathcal{L}^{\#}$ satisfies (4.39).
-The converse follows form Propositions 4.5 and 4.6 and Theorem 2.2.
+If $R_{1}$ is contractive, Proposition 4.9 gives dissipativity of $A_{Y,R_{1}}$, and Remark 4 gives dissipativity of $A_{Y,R_{1}}^{*}$.
+The Lumer–Phillips theorem yields the conclusion.
+Conversely, if $A_{Y,R_{1}}$ generates a contraction semigroup it is dissipative, and Proposition 4.9 gives contractivity of $R_{1}$.
 ∎
 4.3.3. Examples
-Let us show an example an extension $A_{\mathcal{L}}$ generating semigroups of contractions.
-We intent to present an operator whose first row preserves the looping-edge structure ($\phi(L)=\phi(-L)$) with continuity at the vertex and still keeping several parameters.
-For simplicity we consider a tadpole graph ($N=1$) with $\alpha_{0}=\alpha_{1}=\beta_{0}=\beta_{1}$.
-Example 4 ($\delta$ type interactions on a tadpole graph).
-Let $m_{1},\dots,m_{4}\in\mathbb{R}$ and let
-$M:=\begin{pmatrix}1&0&0&0&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{1}&m_{2}+m_{3}&2&m_{2}&0&0\\ 0&0&0&1&0&0\\ \frac{m_{2}+m_{3}}{2}&1&0&1&0&0\\ m_{3}&2&0&m_{4}&0&2\end{pmatrix}.$ (4.40)
-According to (2.4), $M$ is a $(\mathcal{C}_{1},\mathcal{C}_{0})$-contraction if for all $x\in\mathbb{C}^{6}$
-$\left(C_{0}Mx\mid Mx\right)-\left(C_{1}x\mid x\right)\leqslant 0,$ (4.41)
-i.e., if the matrix $M^{*}C_{0}M-C_{1}$ is negative semidefinite.
-It can be seen that it is enough to consider
-$\alpha_{0}<0,\ m_{4}\leqslant\frac{3}{2}\mbox{ and }4m_{1}-(m_{2}^{2}+m_{3}^{2})\leqslant 2.$ (4.42)
-Now, according to (4.32) we have
-$M^{\#}=\begin{pmatrix}1&0&0&0&0&0\\ \frac{-m_{2}-m_{3}}{2}&\frac{1}{2}&0&-1&\frac{1}{2}&0\\ \frac{m_{1}-1}{2}&\frac{-m_{2}-m_{3}}{4}&\frac{1}{2}&\frac{m_{3}}{2}&\frac{-m_{2}-m_{3}}{4}&0\\ 0&0&0&1&0&0\\ 0&0&0&0&0&0\\ \frac{m_{2}}{2}&\frac{-1}{2}&0&\frac{m_{4}-1}{2}&\frac{-1}{2}&\frac{1}{2}\end{pmatrix}.$ (4.43)
-In order to $M_{z}^{\#}$ be a $(\mathcal{C}_{0},\mathcal{C}_{1})$-contraction we require for all $y\in\mathbb{C}^{6}$ that
-$\left(C_{1}M^{\#}y\mid M^{\#}y\right)\leqslant\left(C_{0}y\mid y\right),$ (4.44)
-i.e., we require the matrix $(M^{\#})^{*}C_{1}M^{\#}-C_{0}$ to be negative semidefinite The latter holds under the exact same condition (4.42).
-The operator $M$ defines an operator $A_{M}$ acting as $A_{M}\textbf{U}=-A_{0}^{*}\textbf{U}$ with domain consisting of functions $(\phi,\psi)\in D(A_{0}^{*})$ such that
-$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=\left(\frac{m_{2}+m_{3}+2}{2}\right)\phi(-L),\\ &\phi^{\prime\prime}(L)-2\psi^{\prime\prime}(L)=(m_{3}+m_{4})\psi(L)+2\phi^{\prime}(-L)\ \mbox{ and }\\ &\phi^{\prime\prime}(L)-2\phi^{\prime\prime}(-L)=(m_{1}+m_{2})\phi(-L)+(m_{2}+m_{3})\phi^{\prime}(L).\end{split}$ (4.45)
-Note if you take the parameters satisfying $m_{4}=m_{1}=0$, $m_{3}=-2$ and denoting $z=m_{2}/2\in\mathbb{R}$ it follows from (4.45)
-$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-\phi^{\prime}(-L)=z\phi(-L)\mbox{ and }\\ &\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)-\psi^{\prime\prime}(-L)=z\phi^{\prime}(-L)+(z-1)\phi(-L).\end{split}$ (4.46)
-We conclude from Proposition 4.5, Proposition 4.6 and Theorem 4.7 that $A_{M}$ generates a semigroup of contractions.
-4.3. Boundary system separating derivatives
-A better description of the boundary behavior can be performed by separating the relation on the first derivatives exhibited by the matrices $B_{j}$ in (4.2).
-It can be attained by rewriting Proposition 4.1 in the following way.
-We recall the notation associated to U, $\vec{\textbf{{U}}}$, $\vec{\textbf{{U}}}^{\prime}$ and $\vec{\textbf{{U}}}^{\prime\prime}$ given in subsection 2.1.
-Lemma 4.8.
-For any $\textbf{U}=(\phi,\psi_{1}\dots,\psi_{N})$ and $\textbf{V}=(\tilde{\phi},\tilde{\psi}_{1},\dots,\tilde{\psi}_{N})$ in $D(A_{0}^{*})$ we have
-$\begin{split}&[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]=\left(D_{\beta}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime\prime}\mid\vec{\textbf{V}}\right)+\left(D_{\alpha}\vec{\textbf{{U}}}\mid\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{{U}}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right),\end{split}$ (4.47)
-where
-$\begin{split}&D_{\alpha}:=\begin{pmatrix}\alpha_{0}&&&&\\ &-\alpha_{0}&&&&\\ &&\alpha_{1}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\beta}:=\begin{pmatrix}\beta_{0}&&&&\\ &-\beta_{0}&&&&\\ &&\beta_{1}&&\\ &&&\ddots&\\ &&&&\beta_{N}\end{pmatrix}.\end{split}$ (4.48)
-Definition 4.6.
-Consider the spaces $\mathcal{K}_{+}\cong\mathcal{K}_{-}\cong\mathbb{C}^{N+1}$ and endow them with the weighted products
-$\left(x\mid y\right)_{\mathcal{K}_{+}}:=\left(D^{+}_{\alpha}x\mid y\right)\mbox{ and }\left(x\mid y\right)_{\mathcal{K}_{-}}:=\left(D_{\alpha}^{-}x\mid y\right),$ (4.49)
-where
-$\begin{split}&D_{\alpha}^{+}:=\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{1}&&&&\\ &&\alpha_{2}&&\\ &&&\ddots&\\ &&&&\alpha_{N}\end{pmatrix}\mbox{ and }\ D_{\alpha}^{-}:=\frac{1}{N+1}\begin{pmatrix}\alpha_{0}&&&&\\ &\alpha_{0}&&&&\\ &&\alpha_{0}&&\\ &&&\ddots&\\ &&&&\alpha_{0}\end{pmatrix}.\end{split}$
-Then $(\mathcal{K}_{\pm},(\cdot\mid\cdot)_{\mathcal{K}_{\pm}})$ are Krein spaces.
-Note the matrices $D_{\alpha}^{-}$ and $D_{\alpha}^{+}$ are simply splitting the matrix $D_{\alpha}$ into two matrices, one dealing with the coefficient of $\phi(L)$ and another dealing with the rest of the coefficients.
-In fact,
-$\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.50)
-From (4.47) and (4.50) we can rewrite
-$\begin{split}[A_{0}^{*}\textbf{U},\textbf{V}]+[\textbf{U},A_{0}^{*}\textbf{V}]&=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)-\left(D_{\alpha}\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}^{\prime}\right)\\ &=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)+\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\hskip 28.45274pt-\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}+\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.51)
-4.3.1. Extensions generating contractile dynamics via splitting derivatives
-Let us introduce extensions of $A_{0}$ that are parameterized by two items: a linear subspace $Y$ that characterize the interaction between the function and its second derivatives and a linear operation $\mathcal{L}$ that governs how the first derivatives relate each other.
-Different than in the case of subsection 4.2.1, we cannot obtain mixed expression involving the function and its first derivatives with these extensions $A_{Y,\mathcal{L}}$ defined next.
-One possible application of the results below, can be to investigate the existence and stability of solitary waves for the KdV model (1.6) on looping edge graphs in terms of the coefficients of the equations on each edge (see [39]-section 3 and Theorem 4.2).
-Definition 4.7.
-Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$. We define $A_{Y,\mathcal{L}}$ on the set
-$\begin{split}D(A_{Y,\mathcal{L}}):=\left\{\textbf{U}\in D(A_{0}^{*})\mid\right.&\vec{\textbf{U}}\in Y,\ D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right\},\end{split}$ (4.52)
-acting as $A_{Y,\mathcal{L}}\textbf{U}=-A_{0}^{*}\textbf{U}$.
-The following proposition describe explicitly the form of the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$.
-Proposition 4.9.
-Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and consider a closed subspace $Y\subset\mathbb{C}^{N+2}$.
-We have that the adjoint operator $A_{Y,\mathcal{L}}^{*}$ of $A_{Y,\mathcal{L}}$ is defined on the domain
-$\begin{split}D(A_{Y,\mathcal{L}}^{*})=\{\textbf{V}\in D(A_{0}^{*})\mid&\vec{\textbf{V}}\in Y,\ \frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}\mbox{ and }\\ &\hskip 8.53581pt\left.\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\right\}\end{split}$ (4.53)
-acting as $A_{Y,\mathcal{L}}^{*}\textbf{V}=A_{0}^{*}\textbf{V}$.
-Proof.
-Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ and $\textbf{V}\in D(A_{0}^{*})$. We have
-$\begin{split}[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]&=[-A_{0}^{*}\textbf{U},\textbf{V}]\\ &=[\textbf{U},A_{0}^{*}\textbf{V}]-\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)\\ &\ +\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.54)
-In order to V be in $D(A_{Y,\mathcal{L}}^{*})$ it must satisfy $[A_{Y,\mathcal{L}}\textbf{U},\textbf{V}]=[\textbf{U},A_{Y,\mathcal{L}}^{*}\textbf{V}]$ for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$.
-Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that ${\textbf{U}}=c{\textbf{U}}^{\prime}=\vec{\textbf{0}}$.
-It follows from (4.54) that $\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\mid\vec{\textbf{V}}\right)=0$.
-Since by definition of $A_{Y,\mathcal{L}}$ we have $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}\in Y^{\bot}$, we conclude
-$\vec{\textbf{V}}\in Y.$ (4.55)
-Now, Let $x\in Y$ be arbitrary.
-There exists $\textbf{U}\in D(A_{Y,\mathcal{L}})$ such that $\vec{\textbf{U}}=x$ with $\vec{\textbf{U}}^{\prime}=0$.
-Since $D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}}$ is in $Y^{\bot}$, from (4.54) we get
-$\left(x,\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(\vec{\textbf{U}},\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\right)=\left(D_{\alpha}\vec{\textbf{U}}^{\prime\prime}+\frac{1}{2}D_{\beta}\vec{\textbf{U}},\vec{\textbf{V}}\right)=0.$ (4.56)
-We conclude
-$\frac{1}{2}D_{\beta}\vec{\textbf{V}}+D_{\alpha}\vec{\textbf{V}}^{\prime\prime}\in Y^{\bot}$ (4.57)
-From (4.54), (4.55) and (4.57) we obtain that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$ we must have
-$\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}=\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(L)\\ \tilde{\phi}^{\prime}(L)\\ \vdots\\ \tilde{\phi}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.$ (4.58)
-From definition of $D(A_{Y,\mathcal{L}})$, (4.58) is equivalent to have that for all $\textbf{U}\in D(A_{Y,\mathcal{L}})$
-$\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}_{1}^{\prime}(L)\\ \vdots\\ \tilde{\psi}_{N}^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.$ (4.59)
-It is clear that
-$\left\{\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}})\right\}\subset D(\mathcal{L}).$
-Moreover, the equality holds since for any $x\in D(\mathcal{L})$ there exists $\textbf{V}\in D(A_{Y,\mathcal{L}})$ such that $x=\begin{pmatrix}\tilde{\phi}^{\prime}(-L)\\ \tilde{\psi}^{\prime}_{1}(L)\\ \vdots\\ \tilde{\psi}^{\prime}_{N}(L)\end{pmatrix}$.
-Therefore, (4.59) is equivalent to say
-$\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L}^{\#})\mbox{ with }\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(-L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}.$ (4.60)
-Proposition 4.10.
-Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace.
-$A_{Y,\mathcal{L}}$ is dissipative if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
-Proof.
-Let $\textbf{U}\in D(A_{Y,\mathcal{L}})$. We have
-$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]&=[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]+[\textbf{U},A_{Y,\mathcal{L}}\textbf{U}]=-[A_{0}^{*}\textbf{U},\textbf{U}]-[\textbf{U},A_{0}^{*}\textbf{U}]\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}-\left(\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\mid\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}_{1}(L)\\ \vdots\\ \psi^{\prime}_{N}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}.\end{split}$ (4.61)
-Since $\textbf{U}\in D(A_{Y,\mathcal{L}})$ is arbitrary we conclude (arguing as in the proof of Proposition 4.9) that $\mbox{Re}[A_{Y,\mathcal{L}}\textbf{U},\textbf{U}]\leqslant 0$ if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction.
-∎
-The following two propositions characterize when the extensions $A_{Y,\mathcal{L}}$ are dissipative.
-Proposition 4.11.
-Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace.
-$A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
-Proof.
-Let $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$. We have
-$\begin{split}2\mbox{Re}[A_{Y,\mathcal{L}}^{*}\textbf{V},\textbf{V}]&=[A_{0}^{*}\textbf{V},\textbf{V}]+[\textbf{V},A_{0}^{*}\textbf{V}]=\\ &=\left(\begin{pmatrix}\phi^{\prime}(-L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{-}}-\left(\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\mathcal{L}^{\#}\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\right)_{\mathcal{K}_{+}}.\end{split}$ (4.62)
-It is clear that
-$\left\{\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}\mid\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})\right\}\subset D(\mathcal{L}^{\#}).$
-Moreover, the equality holds since for any $y\in D(\mathcal{L}^{\#})$ there is $\textbf{V}\in D(A_{Y,\mathcal{L}}^{*})$ such that
-$y=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\\ \vdots\\ \phi^{\prime}(L)\end{pmatrix}.$
-From (4.62) we conclude $A_{Y,\mathcal{L}}^{*}$ is dissipative if and only if $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
-The following theorem is a direct consequence of the previous two propositions and can be proved as done in the proof of Theorem 4.7.
-Theorem 4.12.
-Let $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a densely defined operator and let $Y\subset\mathbb{C}^{N+2}$ be a closed subspace.
-$A_{Y,\mathcal{L}}$ generates a contraction semigroup if and only if $\mathcal{L}$ is a $(\mathcal{K}_{+},\mathcal{K}_{-})$-contraction and $\mathcal{L}^{\#}$ is a $(\mathcal{K}_{-},\mathcal{K}_{+})$-contraction.
-4.3.2. Examples
-Let us first describe the operators $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ that define contractions so that its respective adjoint $\mathcal{L}^{\#}$ is also a contraction.
-Let us assume for simplicity $N=1$ (i.e.
-a tadpole graph) and $\alpha_{0}=\alpha_{1}$.
+For simplicity we take $N=1$ (tadpole graph) and $\alpha_{0}=\alpha_{1}$.
 Example 5.
 For $m_{1},m_{2},m_{3}\in\mathbb{R}$ with $m_{3}\neq 0$, consider the operator
-$\mathcal{L}:=\begin{pmatrix}m_{1}&m_{2}\\ m_{3}&\frac{-m_{1}m_{2}}{m_{3}}\end{pmatrix}.$
-Note we have
-#$\mathcal{L}^{\#}=\begin{pmatrix}\frac{m_{1}}{2}&\frac{m_{3}}{2}\\ \frac{m_{2}}{2}&\frac{-m_{1}m_{2}}{2m_{3}}\end{pmatrix}.$
-It can be seen that both, $\mathcal{L}$ and $\mathcal{L}^{\#}$ are contractions under the conditions
-$m_{3}^{2}+m_{1}^{2}\leqslant 2,\ m_{2}^{2}(m_{1}^{2}+1)\leqslant 2,\ m_{1}(m_{3}^{2}-m_{2}^{2})=0,\ m_{1}^{2}+m_{2}^{2}\leqslant 1\ \mbox{and }\frac{m_{1}^{2}m_{2}^{2}}{m_{3}^{2}}+m_{3}^{2}\leqslant 2.$ (4.63)
-In case $m_{3}=m_{2}\neq 0$, (4.63) reduces to
-$m_{2}^{2}+m_{1}^{2}\leqslant 2\mbox{ and }m_{2}^{2}(m_{1}^{2}+1)\leqslant 2.$
-In particular, if we take $m_{2}=m_{1}=1$ you get that
-$\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$
-defines a contraction with $(\mathcal{K}_{-},\mathcal{K}_{+})$-adjoint also a contraction.
-The next example pretends to characterize the extensions obtained via separating derivatives that maintain the looping-edge structure ($\phi(-L)=\phi(L)$) and include continuity at the vertex (when $z=1$ below).
+$\mathcal{L}:=\begin{pmatrix}m_{1}&m_{2}\\ m_{3}&-\tfrac{m_{1}m_{2}}{m_{3}}\end{pmatrix}$
+and the operator-type relation $R_{1}:=\operatorname{graph}(\mathcal{L})\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$. Its Krein adjoint is $R_{1}^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with
+$\mathcal{L}^{\#}=\begin{pmatrix}\tfrac{m_{1}}{2}&\tfrac{m_{3}}{2}\\[4.0pt] \tfrac{m_{2}}{2}&-\tfrac{m_{1}m_{2}}{2m_{3}}\end{pmatrix}$.
+Both $R_{1}$ and $R_{1}^{[*]}$ are contractive (hence, by Theorem 4.10, $A_{Y,R_{1}}$ generates a contraction semigroup) under the conditions
+$m_{3}^{2}+m_{1}^{2}\leqslant 2,\quad m_{2}^{2}(m_{1}^{2}+1)\leqslant 2,\quad m_{1}(m_{3}^{2}-m_{2}^{2})=0,\quad m_{1}^{2}+m_{2}^{2}\leqslant 1,\quad\frac{m_{1}^{2}m_{2}^{2}}{m_{3}^{2}}+m_{3}^{2}\leqslant 2.$ (4.55)
+In the case $m_{3}=m_{2}\neq 0$, (4.55) reduces to $m_{2}^{2}+m_{1}^{2}\leqslant 2$ and $m_{2}^{2}(m_{1}^{2}+1)\leqslant 2$.
+In particular, $m_{1}=m_{2}=1$ gives the contractive relation $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$.
+The next example characterizes extensions that maintain the looping-edge structure ($\phi(-L)=\phi(L)$) and include continuity at the vertex.
 Example 6.
-Let $z\in\mathbb{R}$ be fixed and $\mathcal{L}:\mathcal{K}_{+}\to\mathcal{K}_{-}$ be a linear operator.
-Consider the subspace $Y_{z}:=\mbox{span}\{(1,1,z)^{T}\}$.
-Let us first assume $z\neq 0$.
-We have that
-$Y_{z}^{\perp}=\left\{(x,y,w)^{T}\in\mathbb{C}^{3}\mid w=\frac{-x-y}{z}\right\}.$
-From the Definition 4.7 we have that the domain of $A_{Y_{z},\mathcal{L}}$ is made of functions $u\in D(A_{0}^{*})$ such that
-$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=z\phi(L),\ z\left(\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right)\\ &\mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$ (4.64)
-For instance, if we take $z=1$ and $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we have that $A_{Y_{1},\mathcal{L}}$ with domain
-$\begin{split}D(A_{Y_{1},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L)=\psi(L),\ \phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\ \psi^{\prime}(L)=0\mbox{ and }\alpha_{1}\psi^{\prime\prime}(L)+\frac{\beta_{1}}{2}\psi(L)=\alpha_{0}(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L))\}.\end{split}$ (4.65)
-defines a contractile dynamic on the tadpole graph.
-$-L$
-$L$
-$\psi$
-$\phi$
-Figure 3. Example of profile in (4.66)
-On the other hand, assume $z=0$. In such case
-$Y_{0}^{\perp}=\{(x,y,w)\in\mathbb{C}^{3}\mid x=-y\}.$
-Therefore the domain of $A_{Y_{0},\mathcal{L}}$ is made of functions $\textbf{U}\in D(A_{0}^{*})$ such that
-$\begin{split}&\phi(-L)=\phi(L),\ \psi(L)=0,\ \phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\ \mbox{and }\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}\in D(\mathcal{L})\ \mbox{with }\mathcal{L}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}.\end{split}$
-In particular if we take $\mathcal{L}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ we get that $A_{Y_{0},\mathcal{L}}$ with domain
-$\begin{split}D(A_{Y_{0},\mathcal{L}})=\{\textbf{U}\in D(A_{0}^{*})\mid&\ \phi(L)=\phi(-L),\ \psi(L)=0,\ \phi^{\prime}(L)=\phi^{\prime}(-L)\mbox{ and }\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\}.\end{split}$ (4.66)
-defines an extension of $A_{0}$ generating a contractive dynamic on the tadpole.
+Let $z\in\mathbb{R}$ and $R_{1}\subset\mathcal{K}_{+}\times\mathcal{K}_{-}$ be a contractive boundary relation. Consider the subspace $Y_{z}:=\operatorname{span}\{(1,1,z)^{\top}\}$.
+Case $z\neq 0$$Y_{z}^{\perp}=\bigl\{(x,y,w)^{\top}\in\mathbb{C}^{3}:w=-(x+y)/z\bigr\}$. By Definition 4.9, the domain of $A_{Y_{z},R_{1}}$ consists of all $\mathbf{U}\in D(A_{0}^{*})$ satisfying
+$\phi(-L)=\phi(L),\quad\psi(L)=z\,\phi(L),\quad z\!\left(\alpha_{1}\psi^{\prime\prime}(L)+\tfrac{\beta_{1}}{2}\psi(L)\right)=\alpha_{0}\!\left(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\right),$ (4.56)
+together with $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$.
+Taking $z=1$ and $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$ (which is contractive by Example 5), the domain specializes to
+$\begin{split}D(A_{Y_{1},R_{1}})=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\phi(L)=\phi(-L)=\psi(L),\quad\phi^{\prime}(L)=\phi^{\prime}(-L),\\ &\psi^{\prime}(L)=0,\quad\alpha_{1}\psi^{\prime\prime}(L)+\tfrac{\beta_{1}}{2}\psi(L)=\alpha_{0}\bigl(\phi^{\prime\prime}(L)-\phi^{\prime\prime}(-L)\bigr)\Bigr\},\end{split}$ (4.57)
+and by Theorem 4.10, $A_{Y_{1},R_{1}}$ generates a contraction semigroup on the tadpole graph.
+[…]
+$\phi(-L)=\phi(L),\quad\psi(L)=0,\quad\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L),$ (4.58)
+together with $(\Lambda_{+}\mathbf{U},\Lambda_{-}\mathbf{U})\in R_{1}$.
+Taking $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, the domain specializes to
+$\begin{split}D(A_{Y_{0},R_{1}})=\Bigl\{\mathbf{U}\in D(A_{0}^{*})\;\Big|\;&\phi(L)=\phi(-L),\quad\psi(L)=0,\\ &\phi^{\prime}(L)=\phi^{\prime}(-L),\quad\phi^{\prime\prime}(L)=\phi^{\prime\prime}(-L)\Bigr\},\end{split}$ (4.59)
+and $A_{Y_{0},R_{1}}$ generates a contraction semigroup on the tadpole graph.
+Example 7.
+Assume $\alpha_{j}\equiv\alpha$ and $\beta_{j}\equiv\beta$ for all $j=0,\dots,N$. Let
+$Y:=\operatorname{span}\{(0,0,1,\dots,1)^{T}\}\subset\mathbb{R}^{N+2},$
+so that
+$Y^{\perp}=\left\{(y_{1},y_{2},x_{1},\dots,x_{N})\in\mathbb{R}^{N+2}\;\Bigm|\;\sum_{j=1}^{N}x_{j}=0\right\}.$
+For $N\in\mathbb{N}$ arbitrary and parameters $m_{1},\dots,m_{N}\in\mathbb{R}$ satisfying $m_{j}^{2}\leqslant N+1$ for all $j=1,\dots,N$, consider the operator-type boundary relation
+$R_{1}:=\operatorname{graph}(\mathcal{L}),\qquad\mathcal{L}:=\operatorname{diag}(1,m_{1},\dots,m_{N}):\mathcal{K}_{+}\to\mathcal{K}_{-}.$ (4.60)
+Its Krein adjoint relation is $R_{1}^{[*]}=\operatorname{graph}(\mathcal{L}^{\#})$ with $\mathcal{L}^{\#}=\frac{1}{N+1}\mathcal{L}$.
+The condition $m_{j}^{2}\leqslant N+1$ ensures that both $R_{1}$ and $R_{1}^{[*]}$ are $(\mathcal{K}_{+},\mathcal{K}_{-})$-contractions in the boundary system of Section 4.3, so that Theorem 4.10 applies.
+The corresponding extension $A_{Y,R_{1}}$ is defined on
+$\begin{split}D(A_{Y,R_{1}}):=\Big\{\mathbf{U}\in H^{3}(\mathcal{G})\;\Bigm|\;&\phi(-L)=\phi(L)=0,\quad\psi_{1}(L)=\psi_{j}(L)\text{ for all }j=1,\dots,N,\\ &\phi^{\prime}(-L)=\phi^{\prime}(L)=m_{j}\psi_{j}^{\prime}(L)\text{ for all }j=1,\dots,N,\\ &\sum_{j=1}^{N}\psi_{j}^{\prime\prime}(L)=\frac{-\beta}{2\alpha}\,N\psi_{1}(L)\Big\},\end{split}$ (4.61)
+and is the infinitesimal generator of a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$.
+The special case $m_{j}=0$ for all $j=1,\dots,N$ corresponds to the boundary relation
+$R_{1}^{0}:=\operatorname{graph}(\operatorname{diag}(1,0,\dots,0)),$
+whose multivalued part satisfies $\operatorname{mul}(R_{1}^{0})=\{(0,y_{2},\dots,y_{N}):y_{j}\in\mathbb{C}\}$, reflecting the fact that the first-derivative traces $\psi_{1}^{\prime}(L),\dots,\psi_{N}^{\prime}(L)$ are forced to zero without any constraint on $\Lambda_{-}\mathbf{U}$.
+This is a genuine relation, not representable by any invertible operator, and the domain reduces to
+$\begin{split}D(A_{Y,R_{1}^{0}})=\Big\{\mathbf{U}\in H^{3}(\mathcal{G})\;\Bigm|\;&\phi(L)=\phi(-L)=\phi^{\prime}(L)=\phi^{\prime}(-L)=0,\\ &\psi_{1}(L)=\psi_{j}(L)\text{ for all }j=1,\dots,N,\\ &\psi_{j}^{\prime}(L)=0\text{ for all }j=1,\dots,N,\\ &\sum_{j=1}^{N}\psi_{j}^{\prime\prime}(L)=\frac{-\beta}{2\alpha}\,N\psi_{1}(L)\Big\}.\end{split}$ (4.62)
+The operator $A_{Y,R_{1}^{0}}$ generates a $C_{0}$-semigroup of contractions by Theorem 4.10, and its domain coincides with the looping-edge analogous of the class of vertex conditions employed in the study of stabilization and controllability of dispersive equations on star-shaped networks [6, 48].
+4.4. A simple application to exponential stabilization
+For simplicity we work within the setting of Example 6 (case $z=0$) on the tadpole graph $\mathcal{G}$ with $N=1$, taking $\alpha_{0}=\alpha_{1}=\alpha>0$ and $\beta_{0}=\beta_{1}=\beta\in\mathbb{R}$.
+The operator $A=A_{Y_{0},R_{1}}$ from (4.59), together with the relation $R_{1}=\operatorname{graph}\!\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}$, acts on all $\mathbf{U}=(\phi,\psi)\in H^{3}(\mathcal{G})$ satisfying
+$\phi(-L)=\phi(L),\qquad\psi(L)=0,\qquad\phi^{\prime}(-L)=\phi^{\prime}(L),\qquad\phi^{\prime\prime}(-L)=\phi^{\prime\prime}(L),$
+together with $\psi^{\prime}(L)=0$. This last condition follows directly from $R_{1}$: the equation $\Lambda_{-}\mathbf{U}=\mathcal{L}\,\Lambda_{+}\mathbf{U}$ reads
+$\begin{pmatrix}\phi^{\prime}(L)\\ \phi^{\prime}(L)\end{pmatrix}=\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\begin{pmatrix}\phi^{\prime}(-L)\\ \psi^{\prime}(L)\end{pmatrix}$,
+and subtracting the two rows forces $\psi^{\prime}(L)=0$, after which either row gives $\phi^{\prime}(L)=\phi^{\prime}(-L)$. Fix $\gamma>0$ and consider the damped problem
+$\partial_{t}\mathbf{U}(t)=A\,\mathbf{U}(t)-\gamma\,\mathbf{U}(t),\qquad t>0,\qquad\mathbf{U}(0)=\mathbf{U}_{0}\in D(A).$ (4.63)
+Theorem 4.11.
+Let $\mathbf{U}(t)$ be a classical solution of (4.63). Then the energy $E(t):=\tfrac{1}{2}\|\mathbf{U}(t)\|^{2}_{L^{2}(\mathcal{G})}$ satisfies
+$E(t)\leqslant e^{-2\gamma t}\,E(0),\qquad t\geqslant 0,$
+or equivalently $\|\mathbf{U}(t)\|_{L^{2}(\mathcal{G})}\leqslant e^{-\gamma t}\|\mathbf{U}_{0}\|_{L^{2}(\mathcal{G})}$ for all $t\geqslant 0$.
+In particular, the solution decays to zero exponentially at rate $\gamma$, independently of the dispersive coefficients $\alpha$ and $\beta$.
+Proof.
+We first address the existence of classical solutions to (4.63).
+Since $A_{Y_{0},R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$ by Theorem 4.10, the operator $A-\gamma I$ is dissipative for every $\gamma>0$.
+Moreover, $(A-\gamma I)^{*}=A^{*}-\gamma I$, and since $A^{*}$ is also dissipative (as noted in Remark 4), so is $A^{*}-\gamma I$.
+The Lumer–Phillips theorem therefore guarantees that $A-\gamma I$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$, and in particular that (4.63) admits a unique classical solution $\mathbf{U}\in C^{1}([0,\infty);L^{2}(\mathcal{G}))\cap C([0,\infty);D(A))$ for every $\mathbf{U}_{0}\in D(A)$.
+Writing (4.63) componentwise gives $\phi_{t}=-\alpha\phi^{\prime\prime\prime}-\beta\phi^{\prime}-\gamma\phi$ on (L,L)(-L,L) and $\psi_{t}=-\alpha\psi^{\prime\prime\prime}-\beta\psi^{\prime}-\gamma\psi$ on $(L,\infty)$.
+Differentiating the energy and substituting the equations of motion yields
+$\frac{d}{dt}E(t)$ $=-\alpha\!\int_{-L}^{L}\!\phi\phi^{\prime\prime\prime}\,dx-\beta\!\int_{-L}^{L}\!\phi\phi^{\prime}\,dx-\gamma\!\int_{-L}^{L}\!\phi^{2}\,dx-\alpha\!\int_{L}^{\infty}\!\psi\psi^{\prime\prime\prime}\,dx-\beta\!\int_{L}^{\infty}\!\psi\psi^{\prime}\,dx-\gamma\!\int_{L}^{\infty}\!\psi^{2}\,dx.$
+We claim that all terms except those with factor $\gamma$ vanish identically. Indeed, integrating by parts twice on the loop edge,
+$\int_{-L}^{L}\phi\,\phi^{\prime\prime\prime}\,dx=\Bigl[\phi^{\prime\prime}\phi\Bigr]_{-L}^{L}-\int_{-L}^{L}\phi^{\prime\prime}\phi^{\prime}\,dx=\bigl(\phi^{\prime\prime}(L)\phi(L)-\phi^{\prime\prime}(-L)\phi(-L)\bigr)-\frac{1}{2}\bigl(\phi^{\prime}(L)^{2}-\phi^{\prime}(-L)^{2}\bigr),$
+and every boundary term cancels upon applying $\phi(-L)=\phi(L)$, $\phi^{\prime}(-L)=\phi^{\prime}(L)$, and $\phi^{\prime\prime}(-L)=\phi^{\prime\prime}(L)$, giving $\int_{-L}^{L}\phi\,\phi^{\prime\prime\prime}\,dx=0$.
+Besides, since $\phi\phi^{\prime}=\tfrac{1}{2}(\phi^{2})^{\prime}$,
+$\int_{-L}^{L}\phi\,\phi^{\prime}\,dx=\frac{1}{2}\bigl(\phi(L)^{2}-\phi(-L)^{2}\bigr)=0,$
+again by $\phi(-L)=\phi(L)$.
+On the other hand, functions in $D(A)$ belong to $H^{3}(L,\infty)$, so $\psi$, $\psi^{\prime}$, and $\psi^{\prime\prime}$ all decay to zero as $x\to\infty$.
+Integrating by parts on the half-line therefore gives
+$\int_{L}^{\infty}\psi\,\psi^{\prime\prime\prime}\,dx=\Bigl[\psi^{\prime\prime}\psi\Bigr]_{L}^{\infty}-\int_{L}^{\infty}\psi^{\prime\prime}\psi^{\prime}\,dx=-\psi^{\prime\prime}(L)\psi(L)+\frac{1}{2}\psi^{\prime}(L)^{2}=0,$
+where the last equality uses $\psi(L)=0$ and $\psi^{\prime}(L)=0$.
+Similarly, $\int_{L}^{\infty}\psi\,\psi^{\prime}\,dx=\tfrac{1}{2}[\psi^{2}]_{L}^{\infty}=-\tfrac{1}{2}\psi(L)^{2}=0$.
+Consequently all boundary contributions vanish and the energy equation reduces to
+$\frac{d}{dt}E(t)=-\gamma\|\mathbf{U}(t)\|^{2}_{L^{2}(\mathcal{G})}=-2\gamma\,E(t).$
+An application of Grönwall’s lemma to the differential inequality $E^{\prime}(t)\leqslant-2\gamma E(t)$ gives the conclusion. ∎
+Remark 5.
+The vanishing of all boundary terms is not a coincidence.
+It is a direct manifestation of the fact that $A_{Y_{0},R_{1}}$ generates a $C_{0}$-semigroup of contractions on $L^{2}(\mathcal{G})$: by Theorem 4.10, this is equivalent to the contractivity of $R_{1}$, which in turn forces the boundary terms arising in the Green identity to be non-positive.
+In the particular case at hand they vanish altogether, reflecting the precise balance between the subspace $Y_{0}$ and the relation $R_{1}$ in the definition of the domain.
+The systematic exploitation of this connection between the contractivity of the boundary relation, the structure of the domain, and the resulting energy dissipation is part of an ongoing work for the KdV equation on general metric graphs.
 5. The Schrödinger operator on looping-edge and $\mathcal{T}$-shaped graphs
-In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering $-L$ as a terminal vertex.
+In this section we employ the same techniques used in the case of the free Airy operator to characterize the extensions of the Schrödinger operator on a looping-edge and a graph determined by the edges $e_{0}=[-L,L]$ and $e_{j}=[L,\infty)$ considering L-L as a terminal vertex.
 This strategy will allow us in the future to study the dynamics of NLS model (1.1) specially on $\mathcal{T}$-graphs.
@@
 5.1. Deficiency indices of the free Schrödinger operator
-For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $-\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
+For the symmetric operator $\mathcal{H}_{0}$ in (5.1) we have that free solutions of $\partial_{x}^{2}u_{e}\pm iu_{e}=0$ are given as linear combinations of complex exponentials of the form
 $u_{e}=c_{1}e^{\frac{(1\pm i)}{\sqrt{2}}x_{e}}+c_{2}e^{-\frac{(1\pm i)}{\sqrt{2}}x_{e}}.$
 We note that $e^{\pm\frac{(1\pm i)}{\sqrt{2}}x}$ belong to $L^{2}(-L,L)$.
-On the other hand, on a half-line $[L,\infty)$, it would be square integrable only if the real part of ±$\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative.
+On the other hand, on a half-line $L,\infty)$, it would be square integrable only if the real part of $\pm\frac{(1\pm i)}{\sqrt{2}}$ is negative.
 Along each edge $e=[L,\infty)$ we would therefore always have that the deficiency indices are ${d_{+}}_{e}={d_{-}}_{e}=1$.
@@
 Definition 5.1.
-We define in the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by
+We define on the graph of $\mathcal{H}_{0}^{*}$, $G(\mathcal{H}_{0}^{*})=\{(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mid\textbf{U}\in D(\mathcal{H}_{0}^{*})\}$, the standard skew-symmetric form $\Lambda:G(\mathcal{H}_{0}^{*})\times G(\mathcal{H}_{0}^{*})\to\mathbb{C},$ defined by
+$\Lambda((\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),(\textbf{V},\mathcal{H}_{0}^{*}\textbf{V})):=[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}].$ (5.2)
+Definition 5.2.
+Let $\mathcal{J}=\mathbb{C}^{N+2}$. Define the standard skew-symmetric sesquilinear form on $\mathcal{J}\oplus\mathcal{J}$ by
+$\rho\bigl((x_{1},y_{1}),(x_{2},y_{2})\bigr):=\langle y_{1},x_{2}\rangle_{\mathcal{J}}-\langle x_{1},y_{2}\rangle_{\mathcal{J}},\qquad(x_{i},y_{i})\in\mathcal{J}\oplus\mathcal{J},$ (5.3)
+where $\langle\,\cdot\,,\,\cdot\,\rangle_{\mathcal{J}}$ is the standard inner product on $\mathbb{C}^{N+2}$.
+Definition 5.3.
+We define $\mathbb{F}$ to be the surjective boundary map
+$\begin{split}\mathbb{F}:&\ G(\mathcal{H}_{0}^{*})\to\mathcal{J}\oplus\mathcal{J}\\ &(\mathbf{U},\mathcal{H}_{0}^{*}\mathbf{U})\mapsto\bigl(\Gamma_{0}\mathbf{U},\,\Gamma_{1}\mathbf{U}\bigr),\end{split}$ (5.4)
+where
+$\Gamma_{0}\mathbf{U}=\begin{pmatrix}\phi(-L)\\ \phi(L)\\ \psi_{1}(L)\\ \vdots\\ \psi_{N}(L)\end{pmatrix},\qquad\Gamma_{1}\mathbf{U}=\begin{pmatrix}\phi^{\prime}(-L)\\ -\phi^{\prime}(L)\\ \psi_{1}^{\prime}(L)\\ \vdots\\ \psi_{N}^{\prime}(L)\end{pmatrix}$.
+Proposition 5.1.
+The quintuple $(\Lambda,\mathcal{J},\mathcal{J},\mathbb{F},\rho)$ is a boundary system for $\mathcal{H}_{0}$, that is,
 […]
+Proof.
+The proof follows by integration by parts. ∎
+5.3. Self-adjoint extensions via boundary systems
+The following theorem, taken from [49, Theorem 3.3], gives a complete parametrization of all self-adjoint extensions of $\mathcal{H}_{0}$ in terms of the boundary system constructed in the previous subsection.
+Theorem 5.2.
+An operator $\mathcal{H}$ is a self-adjoint extension of $\mathcal{H}_{0}$ if and only if there exist a subspace $X\subset\mathcal{J}=\mathbb{C}^{N+2}$ and a self-adjoint operator $\mathcal{L}:X\to X$ such that $\mathcal{H}=\mathcal{H}_{X,\mathcal{L}}$, where
+$D(\mathcal{H}_{X,\mathcal{L}})=\Bigl\{\mathbf{U}\in H^{2}(\mathcal{G})\;\Big|\;\Gamma_{0}\mathbf{U}\in X,\quad\mathcal{L}\Gamma_{0}\mathbf{U}=P_{X}\Gamma_{1}\mathbf{U}\Bigr\}.$ (5.5)
+Here $P_{X}:\mathcal{J}\to X$ is the orthogonal projection onto $X$.
+Theorem 5.2 suggests a practical procedure for constructing self-adjoint extensions by prescribing boundary conditions in two stages.
+First, one translates the desired $k$ conditions (with $k<N+2$) into conditions on $\Gamma_{0}$ and $\Gamma_{1}$, which can then be directly incorporated into the definition of a matrix $M:X\to X$.
+The second stage consists of completing the remaining (unconstrained) entries of $M$ so that it becomes self-adjoint, and of determining the resulting conditions imposed by the relation $M\Gamma_{0}\mathbf{U}=P_{X}\Gamma_{1}\mathbf{U}$.
+We exemplify the procedure next.
+5.4. Examples
+Example 8 (Self-adjoint extensions on a looping-edge graph).
+We seek all self-adjoint extensions of $\mathcal{H}_{0}$ whose domain includes only the condition looping condition $\phi(-L)=\phi(L)$.
+This condition constrains $\Gamma_{0}\mathbf{U}$ in a way that the condition $\Gamma_{0}\mathbf{U}\in X$ translates into the choice
+$X:=\bigl\{x\in\mathcal{J}\;\big|\;x_{1}=x_{2}\bigr\},\qquad X^{\perp}=\operatorname{span}\{e_{1}-e_{2}\},\qquad\dim X=N+1,$
+with orthogonal projection $P_{X}v=v-\tfrac{v_{1}-v_{2}}{2}(e_{1}-e_{2})$. Since no further constraint is imposed, $\mathcal{L}:X\to X$ is a free self-adjoint operator on $X$.
+Consider the orthonormal basis for $X$
+$g_{1}:=\frac{e_{1}+e_{2}}{\sqrt{2}},\qquad g_{k}:=e_{k+1},\quad k=2,\ldots,N+1.$
+In this basis the coordinate vectors of $\Gamma_{0}\mathbf{U}$ and $P_{X}\Gamma_{1}\mathbf{U}$ are, respectively,
+$\widetilde{\Gamma}_{0}\mathbf{U}=\bigl(\sqrt{2}\,\phi(-L),\;\psi_{1}(L),\;\ldots,\;\psi_{N}(L)\bigr)^{\!\top},\qquad\widetilde{\Gamma}_{1}\mathbf{U}=\Bigl(\tfrac{\phi^{\prime}(-L)-\phi^{\prime}(L)}{\sqrt{2}},\;\psi_{1}^{\prime}(L),\;\ldots,\;\psi_{N}^{\prime}(L)\Bigr)^{\!\top}.$
+Write the matrix of $\mathcal{L}$ in the given basis in the following block form
+$M=\begin{pmatrix}a&\mathbf{b}^{*}\\ \mathbf{b}&M^{\prime}\end{pmatrix},\qquad a\in\mathbb{R},\quad\mathbf{b}\in\mathbb{C}^{N},\quad M^{\prime}\in M_{N}(\mathbb{C})\text{ Hermitian},$ (5.6)
+where $\mathbf{b}^{*}$ denotes the conjugate transpose of $\mathbf{b}$. This constitutes an $(N+1)^{2}$-real-parameter family.
+Expanding $M\,\widetilde{\Gamma}_{0}\mathbf{U}=\widetilde{\Gamma}_{1}\mathbf{U}$ row by row one gets
+$\phi^{\prime}(-L)-\phi^{\prime}(L)$ $=2a\,\phi(-L)+\sqrt{2}\,\mathbf{b}^{*}\vec{\psi}(L),$ (5.7)
+$\vec{\psi}\,^{\prime}(L)$ $=\sqrt{2}\,\mathbf{b}\,\phi(-L)+M^{\prime}\vec{\psi}(L),$ (5.8)
+where $\vec{\psi}(L)=(\psi_{1}(L),\ldots,\psi_{N}(L))^{\!\top}$ and $\vec{\psi}\,^{\prime}(L)=(\psi_{1}^{\prime}(L),\ldots,\psi_{N}^{\prime}(L))^{\!\top}$.
+Theorem 5.3.
+The self-adjoint extensions of $\mathcal{H}_{0}$ whose domain satisfies the looping condition $\phi(L)=\phi(-L)$ form an $(N+1)^{2}$-real-parameter family $\{\mathcal{H}_{a,\mathbf{b},M^{\prime}}\}$, parametrised by $a\in\mathbb{R}$, $\mathbf{b}\in\mathbb{C}^{N}$, and a Hermitian $M^{\prime}\in M_{N}(\mathbb{C})$ with domain
+[…]
+Remark 6.
+The $2N+1$ real parameters $(a,\operatorname{Re}\mathbf{b},\operatorname{Im}\mathbf{b})$ account for two effects.
+The vector $\mathbf{b}\in\mathbb{C}^{N}$ couples the loop to the half-lines while the scalar $a\in\mathbb{R}$ introduces a ’Robin-type’ condition on the difference $\phi^{\prime}(-L)-\phi^{\prime}(L)$.
+Note in particular, setting $a=0$ and $\mathbf{b}=\mathbf{0}$ in (5.7)–(5.8) induces the pure periodic condition $\phi^{\prime}(-L)=\phi^{\prime}(L)$ together with $\vec{\psi}\,^{\prime}(L)=M^{\prime}\vec{\psi}(L)$ obtaining the fully decoupled self-adjoint extensions $(\mathcal{H}_{M^{\prime}},D(\mathcal{H}_{M^{\prime}}))$ with
+$D(\mathcal{H}_{M^{\prime}})=\Bigl\{\mathbf{U}\in H^{2}(\mathcal{G})\;\Big|\;\phi(-L)=\phi(L),\quad\phi^{\prime}(-L)=\phi^{\prime}(L),\quad M^{\prime}\vec{\psi}(L)=\vec{\psi}\,^{\prime}(L)\Bigr\}.$ (5.9)
+Several relevant extensions arise as special cases of $\mathcal{H}_{M^{\prime}}$:
+  • $M^{\prime}=0$: the loop satisfies periodic boundary conditions and every half-line satisfies a Neumann condition $\psi_{j}^{\prime}(L)=0$ at the vertex.
+  • $M^{\prime}=\alpha I_{N}$, $\alpha\in\mathbb{R}$: uniform Robin condition on the half-lines, $\psi_{j}^{\prime}(L)=\alpha\,\psi_{j}(L)$ for all $j$.
+  • For a real non-zero parameter $Z$, the particular choice $M^{\prime}=\tfrac{1}{Z}\mathbf{1}\mathbf{1}^{T}$ (all entries equal to $\tfrac{1}{Z}$) is of particular interest for future works.
+Opening $\tfrac{1}{Z}\mathbf{1}\mathbf{1}^{T}\vec{\psi}(L)=\vec{\psi}^{\prime}(L)$ we get
+    $\psi_{1}^{\prime}(L)=\cdots=\psi_{N}^{\prime}(L),\qquad\sum_{j=1}^{N}\psi_{j}(L)=Z\psi_{1}^{\prime}(L).$ (5.10)
+    We obtain then the one-parameter family of decoupled self-adjoint extensions $\mathcal{H}_{Z}$ with domain
+[…]
+5.5. Extensions generating unitary dynamics parameterized by linear subspaces
+We now emulate the strategy of separating derivatives used for the Airy operator.
+We note in this case there is no actual separation of derivatives rather a simple reference frame to build some of the self-adjoint realizations.
 Proposition 5.4.
 For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
-$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(P_{+}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(-L)\\ \partial\tilde{\psi}(L)\end{pmatrix}\right)-\left(P_{-}\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\mid\begin{pmatrix}\partial\tilde{\phi}(L)\\ \vdots\\ \partial\tilde{\phi}(L)\end{pmatrix}\right)$ (5.3)
+$\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(Q\vec{\textbf{U}}\mid\vec{\textbf{V}}^{\prime}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right),$ (5.12)
 where
-$P_{+}:=\begin{pmatrix}P_{0}\\ &P_{0}\\ &&\ddots\\ &&&P_{0}\end{pmatrix}\ \mbox{with }P_{0}:=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\ \mbox{and }P_{-}:=\frac{1}{N+1}P_{+}.$
-Proof.
-The proof follows integrating by parts. ∎
-Definition 5.2.
-Consider the spaces $\mathcal{J}_{+}\cong\mathcal{J}_{-}\cong\mathbb{C}^{2(N+1)}$ edowed with the indefinite non-degenerated products
-$\langle x,y\rangle_{+}:=\left(P_{+}x\mid y\right)\ \mbox{and }\langle x,y\rangle_{-}=\left(P_{-}x\mid y\right).$
-$(\mathcal{J}_{\pm},\langle\cdot,\cdot\rangle_{\pm})$ are Krein spaces.
-Definition 5.3.
-We define the function
-$\begin{split}\mathbb{F}:&\ G(\mathcal{H}_{0}^{*})\to\mathcal{J}_{+}\oplus\mathcal{J}_{-}\\ &(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})\mapsto\left(\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(L)\end{pmatrix},\begin{pmatrix}\partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}\right)\end{split}$ (5.4)
-and the sesquilinear form
-$\begin{split}\rho:&\ (\mathcal{J}_{+}\oplus\mathcal{J}_{-})\times(\mathcal{J}_{+}\oplus\mathcal{J}_{-})\to\mathbb{C}\\ &((x,y),(\tilde{x},\tilde{y}))\mapsto\langle x,\tilde{x}\rangle_{+}-\langle y,\tilde{y}\rangle_{-}.\end{split}$ (5.5)
-From (5.2), (5.3) and the previous definition we conclude for $u\in D(\mathcal{H}_{0}^{*})$ that
-$\Lambda((\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}))=\rho(\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U}),\mathbb{F}(\textbf{U},\mathcal{H}_{0}^{*}\textbf{U})).$ (5.6)
-It follows directly from (5.6) and [44] the following theorem.
-Theorem 5.2.
-Let $\mathcal{H}$ be an extension of $\mathcal{H}_{0}$.
-$\mathcal{H}$ is self-adjoint if and only if there exists a subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$, $X$ $\rho$-self-orthogonal with $\mathbb{F}(G(\mathcal{H}))=X$.
-5.3. Extensions generating unitary dynamics
-In a similar fashion as done to obtain Theorem 4.3 in the case of the Airy operator, we can characterize the $\rho$-self-orthogonal subspaces in terms of the graph of a linear operator that, according to Theorem 5.2, characterize the self-adjoint extensions of $\mathcal{H}_{0}$.
-Theorem 5.3.
-A subspace $X\subset\mathcal{J}_{+}\oplus\mathcal{J}_{-}$ is $\rho$-self-orthogonal if and only if there is $\mathcal{L}:\mathcal{J}_{+}\to\mathcal{J}_{-}$ unitary such that $X=G(\mathcal{L})$.
-Proof.
-The proof follows in a similar fashion as in the proof of Theorem 4.3. ∎
-With our first example we show that the extensions inducing the so-called $\delta$ interactions can be seen as a member of a $\frac{(N+2)(N+1)}{2}$-parameters family of extensions generating unitary dynamics.
-Example 7 ($\delta$ type interactions on looping edge graphs).
-Let us first set $N=1$ and let $\mathcal{L}_{1}$ be defined by
-$\mathcal{L}_{1}=\begin{pmatrix}1&0&0&0\\ m_{1}&2&m_{2}&0\\ 0&0&1&0\\ m_{2}&0&m_{3}&2\end{pmatrix}.$ (5.7)
-It is not hard to see that $\mathcal{L}_{1}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator.
-From Theorems 5.2 and 5.3 $\mathcal{L}_{1}$ defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}_{1}}$ of $\mathcal{H}_{0}$ via the relation $\mathbb{F}(G(\mathcal{H}_{\mathcal{L}_{1}}))=G(\mathcal{L}_{1})$.
-That is, an extension $\mathcal{H}_{\mathcal{L}_{1}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
-$\begin{split}&\phi(-L)=\phi(L)=\psi(L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=(m_{1}+m_{2})\phi(-L),\ \mbox{and }\\ &\phi^{\prime}(L)-2\psi^{\prime}(L)=(m_{2}+m_{3})\phi(-L).\end{split}$ (5.8)
-Adding the equations in (5.8) we get
-$\phi^{\prime}(L)-\phi^{\prime}(-L)=\psi^{\prime}(L)+\frac{m_{1}+2m_{2}+m_{3}}{2}\phi(-L).$ (5.9)
-Of course by taking $m_{1}=-m_{3}$ we recover $D_{Z,1}$ defined in (1.5).
-More generally, let us consider $N\in\mathbb{Z}^{+}$ arbitrary and parameters $m_{ij}\in\mathbb{R}$ such that $m_{ij}=m_{ji}$ for $i,j\in\{1,\dots,N+1\}$.
-We define $\mathcal{L}_{N}$ as a $2(N+1)\times 2(N+1)$ matrix given in terms of $2\times 2$ blocks $(A_{ij})_{ij}$ obeying the rule
-$A_{ii}=\begin{pmatrix}1&0\\ m_{ii}&N+1\end{pmatrix}\mbox{ and }A_{ij}=\begin{pmatrix}0&0\\ m_{ij}&0\end{pmatrix}\mbox{ if }i\neq j.$ (5.10)
-We abuse notation to denote with $\psi_{0}(L):=\phi(-L)$ and $\psi^{\prime}_{0}(L)=\phi^{\prime}(-L)$. Note that if $i=j$ we have that
-$A_{ii}\partial\psi_{i-1}(L)=\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi^{\prime}_{i-1}(L)\end{pmatrix}$ (5.11)
-and if $i\neq j$ we have
-$A_{ij}\partial\psi_{j-1}(L)=\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}.$ (5.12)
-From the structure of the rows of $\mathcal{L}_{N}$ it can be seen that it is a full-rank invertible matrix.
-In particular we conclude $\mathcal{L}_{N}$ is injecitve.
-Moreover, a tedious but elementary computation shows that $\mathcal{L}_{N}^{*}P_{-}\mathcal{L}_{N}=P_{+}$, which implies that for all $x,y\in\mathbb{C}^{2(N+1)}$
-$\langle\mathcal{L}_{N}x,\mathcal{L}_{N}y\rangle_{-}=\langle x,y\rangle_{+}.$
-The latter implies $\mathcal{L}_{N}^{\#}=\mathcal{L}_{N}^{-1}$, showing $\mathcal{L}_{N}$ defines a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator.
-From Theorems 5.2 and 5.3, $\mathcal{L}_{N}$ defines a self-adjoint extension $\mathcal{H_{\mathcal{L}_{N}}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions in $D(\mathcal{H}_{0}^{*})$ such that
-$\mathcal{L}_{N}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi_{1}(L)\\ \vdots\\ \partial\psi_{N}(L)\end{pmatrix}=\begin{pmatrix}\partial\phi(L)\\ \partial\phi(L)\\ \vdots\\ \partial\phi(L)\end{pmatrix}.$ (5.13)
-Note from the even rows of $\mathcal{L}_{N}$ that we are a priori fixing $\phi(-L)=\phi(L)=\psi_{1}(L)=\cdots=\psi_{N}(L)$. Taking this into account, let us open the expression (5.13). We have
-$\begin{split}&A_{11}\partial\psi_{0}(L)+A_{12}\partial\psi_{1}(L)+\cdots+A_{1(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &A_{21}\partial\psi_{0}(L)+A_{22}\partial\psi_{1}(L)+\cdots+A_{2(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ &\hskip 142.26378pt\vdots\\ &A_{(N+1)1}\partial\psi_{0}(L)+A_{(N+1)2}\partial\psi_{1}(L)+\cdots+A_{(N+1)(N+1)}\partial\psi_{N}(L)=\partial\phi(L),\ \\ \end{split}$ (5.14)
-Adding the equations in (5.14) we get
-$\sum_{i=1}^{N+1}A_{i1}\partial\psi_{0}(L)+A_{i2}\partial\psi_{1}(L)+\cdots+A_{i(N+1)}\partial\psi_{N}(L)=(N+1)\partial\phi(L).$ (5.15)
-According to (5.11) and (5.12), we can rewrite (5.15) as
-$\frac{1}{N+1}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{N+1}\begin{pmatrix}0\\ m_{ij}\psi_{j-1}(L)\end{pmatrix}+\frac{1}{N+1}\sum_{i=1}^{N+1}\begin{pmatrix}\psi_{i-1}(L)\\ m_{ii}\psi_{i-1}(L)+(N+1)\psi_{i-1}^{\prime}(L)\end{pmatrix}=\begin{pmatrix}\phi(L)\\ \phi^{\prime}(L)\end{pmatrix}.$ (5.16)
-From the continuity at the vertex and the second entries in (5.16) we conclude
-$\phi^{\prime}(L)-\phi^{\prime}(-L)=\sum_{i=1}^{N}\psi^{\prime}_{i}(L)+\phi(-L)\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}.$ (5.17)
-Any choice of the parameters such that $\frac{1}{N+1}\sum_{i,j=1}^{N+1}m_{ij}=Z\in\mathbb{R}$ makes $D(\mathcal{H}_{\mathcal{L}_{N}})=D_{Z,N}$, where $D_{Z,N}$ is defined on (1.5).
-Remark 1.
-It is worth mentioning that the operator $\mathcal{L}_{N}$ was built by fixing continuity at the vertex (see the even rows of $\mathcal{L}_{N}$) and removing the problematic entries that won’t allow $\mathcal{L}_{N}$ to be a real unitary matrix in its boundary system.
-Therefore, we can infer from the previous example that all the self-adjoint extensions of the Schrödinger operator coming from a real matrix that include continuity at the vertex and are posed on a looping-edge graph are of the form $(\mathcal{H},D(\mathcal{H}))=(\mathcal{H}_{0},D_{Z,N})$ with $Z$ varying on $\mathbb{R}$.
-The following example pretend to explore families of operators that induce unitary dynamics and are compatible with continuity of the derivatives at the vertex.
-Example 8 ($\delta^{\prime}$ type interactions on looping edge and $\mathcal{T}$-graphs).
-Consider $N=1$ (tadpole case).
-At first, we begin considering $-L$ as an independent vertex, i.e., we won’t fix the condition $\phi(-L)=\phi(L)$ directly from the operator.
-Let $m_{1},\dots,m_{4}\in\mathbb{R}$ be parameters.
-Define
-$\mathcal{L}=\begin{pmatrix}m_{1}&m_{2}&0&0\\ 0&\frac{2}{m_{1}}&0&m_{3}\\ -m_{1}m_{3}&-m_{2}m_{3}&2&m_{4}\\ 0&0&0&1\end{pmatrix}.$ (5.18)
-It can be seen that $\mathcal{L}$ is a $(\mathcal{J}_{+},\mathcal{J}_{-})$-unitary operator.
-From theorems 5.2 and 5.3 we have a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ with domain made of the functions in $D(\mathcal{H}_{0}^{*})$ such that
-$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)-m_{1}\phi(-L)=m_{2}\phi^{\prime}(-L),\ \phi^{\prime}(L)-2m_{1}^{-1}\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{1}m_{3}\phi(-L)=-m_{2}m_{3}\phi^{\prime}(-L)+2\psi(L)+m_{4}\psi^{\prime}(L).\end{split}$ (5.19)
-In particular if we take $m_{1}=1$ and $m_{2}=0$ we recover the looping structure $\phi(-L)=\phi(L)$ since (5.19) can be rewritten as
-$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L),\ \phi(L)=\phi(-L),\ \phi^{\prime}(L)-2\phi^{\prime}(-L)=m_{3}\psi^{\prime}(L)\\ &\mbox{ and }\phi(L)+m_{3}\phi(-L)-2\psi(L)=m_{4}\psi^{\prime}(L).\end{split}$ (5.20)
-Moreover, if we additionally let $m_{3}=-1$ in (5.20) we get
-$\begin{split}&\psi^{\prime}(L)=\phi^{\prime}(L)=\phi^{\prime}(-L),\ \phi(L)=\phi(-L),\\ &\mbox{ and }\psi(L)=\frac{-m_{4}}{2}\psi^{\prime}(L).\end{split}$ (5.21)
-Remark 2.
-The existence and stability of standing waves profiles for the NLS model (1.1) on a tadpole graph, ${\bf{U}}(x,t)=e^{-i\omega t}(\phi,\psi)$, with $(\phi,\psi)$ satisfying the $\delta^{\prime}$-type interactions in (5.21) is the aim of a work in progress by the authors.
-5.4. Extensions generating unitary dynamics parameterized by linear subspaces
-We now emulate the strategy of separating derivatives used for the Airy operator.
-We note in this case there is no actual separation of derivatives but another reference frame to build the extensions.
-We begin by following the same approach as done in Section 4.3 to rewrite (5.3) in a simpler way.
-Proposition 5.4.
-For any $\textbf{U},\textbf{V}\in D(\mathcal{H}_{0}^{*})$ we have
-$[\mathcal{H}_{0}^{*}\textbf{U},\textbf{V}]-[\textbf{U},\mathcal{H}_{0}^{*}\textbf{V}]=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(Q\vec{\textbf{U}}\mid\vec{\textbf{V}}^{\prime}\right)=\left(Q\vec{\textbf{U}}^{\prime}\mid\vec{\textbf{V}}\right)-\left(\vec{\textbf{U}}\mid Q\vec{\textbf{V}}^{\prime}\right),$ (5.22)
-where
-$Q=\begin{pmatrix}1\\ &-1\\ &&1\\ &&&\ddots\\ &&&&1\end{pmatrix}.$
+$Q=\operatorname{diag}(1,-1,1,\dots,1).$
 Definition 5.4.
 Given a subspace $Y\subset\mathbb{C}^{2+N}$ consider the operator $\mathcal{H}_{Y}$, acting as $\mathcal{H}_{Y}\textbf{U}=\mathcal{H}_{0}\textbf{U}$, with domain
-$D(\mathcal{H}_{Y}):=\left\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{U}}\in Y\ \mbox{and }Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}\right\}.$ (5.23)
+[…]
 The next proposition determine explicitly the adjoint operator of $\mathcal{H}_{Y}$.
 Proposition 5.5.
-For any closed subspace $Y\subset\mathbb{C}^{2+N}$ the operator $\mathcal{H}_{Y}$ is a self-adjoint extension of $\mathcal{H}_{0}$.
+For any subspace $Y\subset\mathbb{C}^{2+N}$ the operator $\mathcal{H}_{Y}$ is a self-adjoint extension of $\mathcal{H}_{0}$.
 Proof.
 It is enough to prove that the adjoint of $\mathcal{H}_{Y}$ is given by $\mathcal{H}_{Y}^{*}\textbf{V}=\mathcal{H}_{0}^{*}\textbf{V}$ with domain
-$D(\mathcal{H}_{Y}^{*}):=\left\{\textbf{V}\in D(\mathcal{H}_{0}^{*})\mid\ \vec{\textbf{V}}\in Y\ \mbox{and }Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}\right\}=D(\mathcal{H}_{Y}).$ (5.24)
+[…]
 It is obvious from the definition of $D(\mathcal{H}_{Y})$ that if $Q\vec{\textbf{V}}^{\prime}\in Y^{\perp}$ and $\vec{\textbf{V}}\in Y$ then $\textbf{V}\in D(\mathcal{H}_{Y}^{*})$.
@@
 Let $\textbf{V}\in D(\mathcal{H}_{0}^{*})$.
-By definition V is in $D(\mathcal{H}_{Y}^{*})$ if and only if $[\mathcal{H}_{Y}\textbf{U},\textbf{V}]=[\textbf{U},\mathcal{H}_{Y}^{*}\textbf{V}]$ for all $\textbf{U}\in D(\mathcal{H}_{Y})$.
+By definition V is in $D(\mathcal{H}_{Y}^{*})$ if and only if $\mathcal{H}_{Y}\textbf{U},\textbf{V}]=[\textbf{U},\mathcal{H}_{Y}^{*}\textbf{V}]$ for all $\textbf{U}\in D(\mathcal{H}_{Y})$.
 Let $x\in Y^{\perp}$ be arbitrary. There exists $\textbf{U}\in D(\mathcal{H}_{Y})$ such that $\vec{\textbf{U}}=0$ and $Q\vec{\textbf{U}}^{\prime}=x$. From (5.12), we must have
@@
 Consider the space $Y_{0}=\mbox{span}\{(1,1,0)^{T}\}$.
-Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$.
+Note $\vec{\textbf{U}}\in Y_{0}$ if and only if $\phi(-L)=\phi(L)$ and $\psi(L)=0$.
 Also, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\phi^{\prime}(L)$.
 We therefore have that
-$D(\mathcal{H}_{Y_{0}})=\{u\in D(\mathcal{H}_{0}^{*})\mid\ \phi(-L)=\phi(L)\ \mbox{and }\phi^{\prime}(L)=\phi^{\prime}(-L)\}.$
-Example 10.
-Consider the space $Y=\mbox{span}\{(1,0,-1)\}$.
-Consequently, $Y^{\perp}=\{(x,y,w)\mid x=w\}$.
-Note $\vec{\textbf{U}}\in Y$ if and only if $\phi(L)=0$ and $\phi(-L)=-\psi(L)$.
-Similarly, $Q\vec{\textbf{U}}^{\prime}\in Y^{\perp}$ if and only if $\phi^{\prime}(-L)=\psi^{\prime}(L)$.
-According to Proposition 5.5, ${\mathcal{H}}_{Y}$ defines a self-adjoint extension of $\mathcal{H}_{0}$ with domain
-$D(\mathcal{H}_{Y})=\{\textbf{U}\in D(\mathcal{H}_{0}^{*})\mid\phi(-L)=-\psi(L),\ \phi(L)=0,\ \mbox{and}\ \phi^{\prime}(-L)=\psi^{\prime}(L)\}.$ (5.25)
-5.5. The Schrödinger operator on $\mathcal{T}$-shaped metric graphs
-As highlighted in the introduction of the section, the analysis performed in the previous subsections for the Schrödinger equation on looping-edge graphs extends to $\mathcal{T}$-shaped graphs, i.e., to graphs $\mathcal{G}$ that consists on an internal edge that can be identified with $[-L,0]$ and a finite amount of half-lines $[0,\infty)$ attached to a common vertex identified with 00 (see Figure 4).
-A similar frame-work can be employed to obtain self-adjoint extensions of $\mathcal{H}_{0}$ posed on these kind of metric graphs.
-We proceed to exhibit some examples of them.
-$L$
-00
-Figure 4. $\mathcal{T}$-shaped metric graph.
-Example 11.
-Let $N=1$ (exactly one half-line attached to 00) and consider again the operator $\mathcal{L}$ defined in (5.18) with $m_{1}=1$ and $m_{2}=0$ (see (5.20)).
-Instead of $m_{3}=-1$, let us now consider $m_{3}=1$.
-It can be obtained analogously to (5.20) an extension of $\mathcal{H}_{0}$ with domain of functions in $H^{2}(\mathcal{G})$ such that
-$\begin{split}&\psi^{\prime}(0+)=\phi^{\prime}(0-),\ \phi^{\prime}(-L)=0,\ \phi(0-)=\phi(-L)\ \\ &\mbox{ and }\phi(0-)-\psi(0+)=\frac{m_{4}}{2}\psi^{\prime}(0+),\end{split}$ (5.26)
-where $0+$ and $0-$ refer to the right and left-handed limits respectively.
-Thus, (5.26) is an example of the kind of domain of extensions that might be related to interactions of $\delta^{\prime}$-type in a $\mathcal{T}$-graph.
-With the following example we intend to present, under this frame of reference, all the extensions generating unitary dynamics with continuity in the derivatives at the vertex.
-This family includes the so-called $\delta^{\prime}$ interactions.
-For simplicity we work with $N=1$ and $N=2$.
-Example 12.
-Let us first set $N=1$ and let $\mathcal{L}$ be defined by
-$\mathcal{L}=\begin{pmatrix}2&m_{1}&0&m_{2}\\ 0&1&0&0\\ 0&m_{2}&2&m_{3}\\ 0&0&0&1\end{pmatrix}.$ (5.27)
-It can be seen that in a respective boundary system $\mathcal{L}$ is a unitary operator which defines a self-adjoint extension $\mathcal{H}_{\mathcal{L}}$ of $\mathcal{H}_{0}$ via the relation $\mathcal{L}\begin{pmatrix}\partial\phi(-L)\\ \partial\psi(0+)\end{pmatrix}=\begin{pmatrix}\partial\phi(0-)\\ \partial\phi(0-)\end{pmatrix}$.
-That is, an extension $\mathcal{H}_{\mathcal{L}}$ with domain made of functions $(\phi,\psi)\in H^{2}(\mathcal{G})$ such that
-$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}(0+),\ \phi(0+)-2\phi(-L)=(m_{1}+m_{2})\psi^{\prime}(0+)\mbox{ and }\\ &\phi(0-)-2\psi(0+)=(m_{2}+m_{3})\psi^{\prime}(0+).\end{split}$ (5.28)
-Adding the last two equations in (5.28) we get
-$\phi(0-)-\phi(-L)=\psi(0+)+\frac{m_{1}+2m_{2}+m_{3}}{2}\psi^{\prime}(0+).$ (5.29)
-In particular, if $m_{1}=-m_{3}$ it follows
-$\phi(0-)-\phi(-L)=\psi(0+)+m_{2}\psi^{\prime}(0+).$ (5.30)
-Similarly, let us assume $N=2$ and consider the operator
-$\mathcal{L}_{2}=\begin{pmatrix}3&m_{1}&0&m_{2}&0&m_{3}\\ 0&1&0&0&0&0\\ 0&m_{2}&3&m_{4}&0&m_{5}\\ 0&0&0&1&0&0\\ 0&m_{3}&0&m_{5}&3&m_{6}\\ 0&0&0&0&0&1\end{pmatrix}$ (5.31)
-It can be seen that $\mathcal{L}_{2}$ defines a unitary extension $H_{\mathcal{L}_{2}}$ of $\mathcal{H}_{0}$ with domain consisting of the functions $(\phi,\psi_{1},\psi_{2})\in H^{2}(\mathcal{G})$ such that
-$\begin{split}&\phi^{\prime}(-L)=\phi^{\prime}(0-)=\psi^{\prime}_{1}(0+)=\psi^{\prime}_{2}(0+),\ \phi(0-)-3\phi(-L)=(m_{1}+m_{2}+m_{3})\psi^{\prime}(0+)\\ &\phi(0-)-3\psi_{1}(0+)=(m_{2}+m_{4}+m_{5})\psi^{\prime}(0+)\mbox{ and }\phi(0-)-3\psi_{2}(0+)=(m_{3}+m_{5}+m_{6})\psi^{\prime}(0+).\end{split}$ (5.32)
-Adding the last three equations we get
-$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+(m_{1}+m_{4}+m_{6}+2m_{2}+2m_{3}+2m_{5})\psi^{\prime}(0+).$ (5.33)
-Of course there is a choice of the parameters that turn (5.33) into
-$\phi(0-)-\phi(-L)=\sum_{j=1}^{2}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.34)
-Moreover, the operator $\mathcal{L}_{2}$ can be generalized to any amount of half-lines $N\in\mathbb{Z}^{+}$ by following a similar procedure to the one done in Example 7.
-That way it can be obtained an extension with domain made of functions satisfying continuity in the derivatives and
-$\phi(0-)-\phi(-L)=\sum_{j=1}^{N}\psi_{j}(0+)+Z\psi^{\prime}(0+),\ \ Z\in\mathbb{R}.$ (5.35)
+[…]
+5.6. The $\mathcal{T}$-shaped graph case
+The abstract theory developed for the looping-edge graph $\mathcal{G}$ can be extended to address $\mathcal{T}$-shaped graphs (see Figure 2).
+These graphs are likewise characterized by a metric graph structure with $V=\{-L,L\}$ and $E=\{[-L,L],[L,\infty),\dots,[L,\infty)\}$.
+In this setting, the vertex L-L is treated as an independent terminal node, and the looping condition $\phi(-L)=\phi(L)$ is not imposed.
+Figure 2. A $\mathcal{T}$-shaped graph
+Via the theory of boundary systems developed in Section 5.2, we characterize in the following example extension with a derivative coupling at the common vertex.
+Example 10 (Extensions with prescribed continuity on the derivatives).
+We seek all self-adjoint extensions of $\mathcal{H}_{0}$ whose domain includes the condition
+$\phi^{\prime}(L)=\psi_{1}^{\prime}(L)=\cdots=\psi_{N}^{\prime}(L).$ (5.15)
+From the boundary map definition and because of the sign in the second component, the conditions $\phi^{\prime}(L)=\psi_{j}^{\prime}(L)$ are equivalent to
+$(\Gamma_{1}\mathbf{U})_{2}+(\Gamma_{1}\mathbf{U})_{j+2}=0,\qquad j=1,\ldots,N,$
+or, setting $v_{j}:=e_{2}+e_{j+2}\in\mathcal{J}$,
+$v_{j}^{\,*}\,\Gamma_{1}\mathbf{U}=0,\qquad j=1,\ldots,N.$ (5.16)
+Since the prescribed conditions place no restriction on $\Gamma_{0}\mathbf{U}$, we take the full space $X=\mathcal{J}$ in Theorem 5.2.
+The domain condition then reads $\mathcal{L}\,\Gamma_{0}\mathbf{U}=\Gamma_{1}\mathbf{U}$ for all $\mathbf{U}\in D(\mathcal{H}_{X,\mathcal{L}})$.
+Substituting into (5.16),
+$v_{j}^{\,*}\,\mathcal{L}\,\Gamma_{0}\mathbf{U}=0\qquad\text{for all }\Gamma_{0}\mathbf{U}\in\mathcal{J},$
+which must hold structurally, i.e. as a condition on the rows of the matrix $\mathcal{L}\in M_{N+2}(\mathbb{C})$:
+$\mathrm{row}_{j+2}(\mathcal{L})=-\,\mathrm{row}_{2}(\mathcal{L}),\qquad j=1,\ldots,N.$ (5.17)
+We require $\mathcal{L}=\mathcal{L}^{*}$.
+The row constraints (5.17) propagate to column constraints via Hermiticity: $\mathrm{col}_{j+2}(\mathcal{L})=-\,\mathrm{col}_{2}(\mathcal{L})$ for $j=1,\ldots,N$.
+Enforcing both simultaneously and reading off the diagonal entries (which must be real), the most general Hermitian matrix satisfying (5.17) is
+$\mathcal{L}=\begin{pmatrix}a&\beta&-\beta&\cdots&-\beta\\[4.0pt] \bar{\beta}&c&-c&\cdots&-c\\[2.0pt] -\bar{\beta}&-c&c&\cdots&c\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ -\bar{\beta}&-c&c&\cdots&c\end{pmatrix},\qquad a,\,c\in\mathbb{R},\quad\beta\in\mathbb{C},$ (5.18)
+where the lower-right $N\times N$ block is the constant matrix with every entry equal to $c$. This constitutes a four-real-parameter family $(a,c,\operatorname{Re}\beta,\operatorname{Im}\beta)$.
+Expanding $\mathcal{L}\,\Gamma_{0}\mathbf{U}=\Gamma_{1}\mathbf{U}$ row by row and introducing the shorthand
+$\sigma:=\phi(L)-\sum_{j=1}^{N}\psi_{j}(L),$
+one obtains the following two independent conditions:
+$\phi^{\prime}(-L)$ $=a\,\phi(-L)+\beta\,\sigma,$ (5.19)
+$-\phi^{\prime}(L)$ $=\bar{\beta}\,\phi(-L)+c\,\sigma.$ (5.20)
+The remaining $N$ equations (rows $3$ through $N+2$) each read $\psi_{j}^{\prime}(L)=\bar{\beta}\,\phi(-L)+c\,\sigma$, which is exactly -(5.20), thereby reproducing the prescribed condition (5.15) automatically.
+Theorem 5.6.
+The self-adjoint extensions of $\mathcal{H}_{0}$ whose domain satisfies (5.15) form the four-real-parameter family $\{\mathcal{H}_{a,c,\beta}\}_{a,c\in\mathbb{R},\,\beta\in\mathbb{C}}$, where
+[…]
+Remark 7.
+Setting $\beta=0$ decouples the left endpoint from the vertex: (5.19) becomes a standard Robin condition $\phi^{\prime}(-L)=a\,\phi(-L)$, and (5.20) becomes a Robin condition at the vertex controlled by $c$ and the combination $\sigma$.
+More precisely, for parameters $Z_{1}\in\mathbb{R}$, $0\neq Z_{2}\in\mathbb{R}$ we have the self-adjoint extension $(H_{Z_{1},Z_{2}},D(H_{Z_{1},Z_{2}}))$ with
+[…]
+The choice $a=c=\beta=0$ yields the pure Neumann extension $\phi^{\prime}(-L)=0$, $\phi^{\prime}(L)=\psi_{j}^{\prime}(L)=0$ for all $j$.
+5.7. A simple application to instability of standing waves of the cubic NLS
+Let $D_{\mathrm{ext}}=D(\mathcal{H})$ where $D(\mathcal{H})$ is defined in (5.9) with $M^{\prime}=0$.
+That is, $D_{\mathrm{ext}}$ defines a self-adjoint extension of the Laplacian inducing pure periodic conditions on the loop and independent pure Neumann conditions on each half-line.
+We seek real-valued standing waves $\mathbf{U}(t,x)=e^{i\omega t}\boldsymbol{\Phi}_{\omega}(x)$ with $\boldsymbol{\Phi}_{\omega}=(\phi_{\omega},\psi_{1,\omega},\dots,\psi_{N,\omega})\in D_{\mathrm{ext}}$.
+Substituting into (1.1) with $p=1$, each component of $\boldsymbol{\Phi}_{\omega}$ satisfies an independent stationary NLS equation $f^{\prime\prime}+f^{3}-\omega f=0$ on its respective edge.
+For the loop component we consider a positive, real, $2L$-periodic solution constructed via the Jacobi dnoidal function.
+By Theorem 2.1 of [9], for $\omega>\omega^{*}:=\tfrac{\pi^{2}}{2L^{2}}$, the function
+$\phi_{\omega}(x)=\eta_{1}\,\mathrm{dn}\!\left(\frac{\eta_{1}}{\sqrt{2}}\,x;\,k\right)$ (5.21)
+with elliptic modulus $k\in(0,1)$ determined by
+$k^{2}(\eta,\omega)=\frac{2\omega-2\eta^{2}_{2}}{2\omega-\eta^{2}_{2}},\qquad\eta^{2}_{1}+\eta^{2}_{2}=2\omega,\qquad 0<\eta_{2}<\eta_{1},$ (5.22)
+solves the stationary loop equation. The half-line stationary equation is solved by the half-soliton
+$\psi_{\omega}(x)=\sqrt{2\omega}\,\operatorname{sech}\!\bigl(\sqrt{\omega}(x-L)\bigr),\qquad x\in[L,+\infty),$
+which is the restriction to $L,+\infty)$ of the full soliton centred at $x=L$.
+The Neumann condition $\psi_{\omega}^{\prime}(L)=0$ holds automatically because the full soliton is even about $L$.
+For $j=1,\dots,n\leqslant N$ we therefore set $\psi_{j,\omega}\equiv\psi_{\omega}$, and for $j=n+1,\dots,N$ we set $\psi_{j,\omega}\equiv 0$.
+The standing wave profile under study is thus (see Figure 3)
+$\boldsymbol{\Phi}_{\omega,n}\equiv\boldsymbol{\Phi}_{\omega}=\bigl(\phi_{\omega},\,\underbrace{\psi_{1,\omega},\dots,\psi_{n,\omega}}_{n\text{ half-solitons}},\,\underbrace{0,\dots,0}_{N-n\text{ zero-solitons}}\bigr)\in D_{\mathrm{ext}},\qquad\omega>\omega^{*}.$ (5.23)
+Figure 3. Schematic of the standing wave profile $\Phi_{\omega,n}$ with $N=3$ and $n=2$.
+Theorem 5.7.
+Let $L>0$ and $N\geqslant 1$. For every $n\in\{1,\dots,N\}$ and every $\omega>\omega_{*}$, the orbit
+$\mathcal{O}(\mathbf{\Phi}_{\omega})=\bigl\{e^{i\theta}\bigl(\phi_{\omega}(\cdot+s),\,\psi_{1,\omega},\dots,\psi_{n,\omega},0,\dots,0\bigr):\theta\in\mathbb{R},\;s\in\mathbb{R}/(2L\mathbb{Z})\bigr\}$
+is orbitally unstable in $H^{1}(\mathcal{G})$.
+Proof.
+The decoupled structure of $D_{\mathrm{ext}}$ reduces global well-posedness on $H^{1}(\mathcal{G})$ to independent problems on each edge.
+On the loop [L,L][-L,L] with periodic boundary conditions, global well-posedness in $H^{1}$ is classical and follows from semigroup theory together with conservation of energy and charge.
+On each half-line $L,+\infty)$ with Neumann boundary condition $\psi^{\prime}(L)=0$, well-posedness is reduced to the full-line case via the reflection method: extending $\psi$ to an even function on $\mathbb{R}$ embeds the problem into the cubic NLS on $\mathbb{R}$ restricted to even initial data, which is invariant under the flow.
+Global solutions for all $t\in\mathbb{R}$ are then obtained from the conservation of energy and charge via the Gagliardo–Nirenberg inequality.
+Fix $\varepsilon>0$ and define the initial datum
+$\mathbf{U}_{\varepsilon}(0):=\bigl(\phi_{\omega},\;\psi_{1,\omega+\varepsilon},\;\psi_{2,\omega},\dots,\psi_{n,\omega},\;0,\dots,0\bigr),$
+where $\psi_{1,\omega+\varepsilon}(x)=\sqrt{2(\omega+\varepsilon)}\,\operatorname{sech}\!\bigl(\sqrt{\omega+\varepsilon}(x-L)\bigr)$ is the half-soliton at the shifted frequency $\omega+\varepsilon$.
+Since $\omega\mapsto\psi_{1,\omega}$ is $C^{1}$ as a map into $H^{1}([L,+\infty))$, the mean value inequality in Banach spaces gives
+$\|\mathbf{U}_{\varepsilon}(0)-\mathbf{\Phi}_{\omega}\|_{H^{1}(\mathcal{G})}=\|\psi_{1,\omega+\varepsilon}-\psi_{1,\omega}\|_{H^{1}([L,+\infty))}\leqslant C_{\omega}\,\varepsilon,$ (5.24)
+where $C_{\omega}=\sup_{\omega^{\prime}\in(\omega,\,\omega+1)}\|\partial_{\omega}\psi_{1,\omega^{\prime}}\|_{H^{1}([L,+\infty))}<\infty$.
+In particular, $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(0),\mathcal{O}(\mathbf{\Phi}_{\omega}))\to 0$ as $\varepsilon\to 0$.
+Since every component of $\mathbf{U}_{\varepsilon}(0)$ is an exact stationary profile at its respective frequency, each component satisfies its own NLS equation with its boundary conditions.
+By the uniqueness of $H^{1}$ solutions, the global solution is therefore given by
+$\mathbf{U}_{\varepsilon}(t)=\bigl(e^{i\omega t}\phi_{\omega},\;e^{i(\omega+\varepsilon)t}\psi_{1,\omega+\varepsilon},\;e^{i\omega t}\psi_{2,\omega},\;\dots,\;e^{i\omega t}\psi_{n,\omega},\;0,\dots,0\bigr).$ (5.25)
+At the time $t_{\varepsilon}:=\pi/\varepsilon$ one has $e^{i(\omega+\varepsilon)t_{\varepsilon}}=e^{i\omega t_{\varepsilon}}\cdot e^{i\pi}=-e^{i\omega t_{\varepsilon}}$, so the solution becomes
+$\mathbf{U}_{\varepsilon}(t_{\varepsilon})=e^{i\omega t_{\varepsilon}}\bigl(\phi_{\omega},\;-\psi_{1,\omega+\varepsilon},\;\psi_{2,\omega},\dots,\psi_{n,\omega},\;0,\dots\bigr),$ (5.26)
+that is, the first half-line component has flipped sign relative to all others.
+We now derive a lower bound on the distance from the orbit at time $t_{\varepsilon}$.
+Take any orbit element $\mathbf{V}=e^{i\theta}(\phi_{\omega}(\cdot+s),\psi_{1,\omega},\dots)\in\mathcal{O}(\mathbf{\Phi}_{\omega})$ and set $\mu:=\theta-\omega t_{\varepsilon}$.
+Since the components are supported on disjoint edges and $\|(1-e^{i\mu})\psi_{j,\omega}\|_{L^{2}}^{2}\geqslant 0$ for $j=2,\dots,N$, we obtain
+$\begin{split}\|\mathbf{U}_{\varepsilon}(t_{\varepsilon})-\mathbf{V}\|_{L^{2}(\mathcal{G})}^{2}&\geqslant\|\phi_{\omega}-e^{i\mu}\phi_{\omega}(\cdot+s)\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}+e^{i\mu}\psi_{1,\omega}\|_{L^{2}}^{2}\\ &=A+2\cos(\mu)\bigl(P-q(s)\bigr),\end{split}$ (5.27)
+where
+$\begin{split}A&:=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2},\\ P&:=\langle\psi_{1,\omega+\varepsilon},\psi_{1,\omega}\rangle_{L^{2}([L,+\infty))},\\ q(s)&:=\langle\phi_{\omega},\phi_{\omega}(\cdot+s)\rangle_{L^{2}([-L,L])}.\end{split}$
+Since $\cos\mu\in[-1,1]$, minimising the right-hand side of (5.27) over $\mu$ and then taking the infimum over $(\theta,s)$ gives
+$\inf_{(\theta,s)}\|\mathbf{U}_{\varepsilon}(t_{\varepsilon})-\mathbf{V}\|_{L^{2}}^{2}\geqslant A-2\sup_{s\in\mathbb{R}/(2L\mathbb{Z})}|P-q(s)|.$ (5.28)
+It is shown in Appendix A that
+$A-2\sup_{s}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
+where $q_{\min}:=\min_{s}q(s)>0$ and $c_{0}$ depends only on $\omega$, $L$, and $k$, but not on $\varepsilon$. Combining with (5.28) we conclude
+$\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\,\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant\mathrm{dist}_{L^{2}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\,\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant c_{0}>0.$ (5.29)
+Finally, given $\delta>0$ arbitrary, one can choose $\varepsilon>0$ small enough that $C_{\omega}\varepsilon<\delta$.
+Then (5.24) gives $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(0),\mathcal{O}(\mathbf{\Phi}_{\omega}))<\delta$, while (5.29) gives $\mathrm{dist}_{H^{1}}(\mathbf{U}_{\varepsilon}(t_{\varepsilon}),\mathcal{O}(\mathbf{\Phi}_{\omega}))\geqslant c_{0}$.
+This is precisely the opposite of orbital stability, and the proof is complete.
+∎
 Acknowledgments
+The authors express their gratitude to Prof.
+Nataliia Goloshchapova (IME-USP) for her invaluable comments on boundary systems, which significantly improved the results presented in the second part of this manuscript.
+This study is financed, in part, by the São Paulo Research Foundation (FAPESP), Brazil. Process number 2024/20623-7.
 A.
-Muñoz is partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil.
+Muñoz was partially supported by CNPq, Conselho Nacional de Desenvolvimento Científico e Tecnológico - Brazil grant 170246/2023-0 and INCTMat, Instituto Nacional de Ciência e Tecnologia de Matemática - Brazil.
 J.
 Angulo was partially funded by CNPq/Brazil Grant.
-Data availability statement. The data that support the findings of this study are available upon request from the authors.
+Appendix A Lower bound computation
+In this appendix we establish the inequality
+$A-2\sup_{s\in\mathbb{R}/2L\mathbb{Z}}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
+used in Section 5.7.
+Since $\phi_{\omega}$ and $\psi_{\omega}$ are strictly positive, both $P$ and $q(s)$ are strictly positive.
+The map $s\mapsto q(s)$ is continuous on the compact set $\mathbb{R}/2L\mathbb{Z}$, so it attains strictly positive extrema $0<q_{\min}<q_{\max}$.
+The maximum is attained at $s=0$, with $q_{\max}=q(0)=\|\phi_{\omega}\|^{2}_{L^{2}}$, as a consequence of the Cauchy–Schwarz inequality and the invariance of the $L^{2}$ norm under periodic translation.
+Since $P$ is independent of $s$, the convexity of the absolute value gives
+$\sup_{s}|P-q(s)|\leqslant\max\bigl(|P-q_{\min}|,\;|P-q_{\max}|\bigr).$
+We argue in three cases according to the position of $P$ relative to $q_{\min},q_{\max}]$.
+Case 1: $P\leqslant q_{\min}$. Since $q(s)\geqslant q_{\min}\geqslant P$ for all $s$, one has |Pq(s)|=q(s)P|P-q(s)|=q(s)-P and the supremum is attained at $s=0$:
+$\sup_{s}|P-q(s)|=q_{\max}-P=\|\phi_{\omega}\|^{2}_{L^{2}}-P.$
+Substituting back and cancelling the loop norm terms:
+$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2\|\phi_{\omega}\|_{L^{2}}^{2}+2P\\ &=\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2P.\end{split}$
+All three terms are non-negative and $\|\psi_{1,\omega}\|_{L^{2}}^{2}=2\sqrt{\omega}$, so
+$A-2\sup_{s}|P-q(s)|\geqslant 2\|\psi_{1,\omega}\|_{L^{2}}^{2}=4\sqrt{\omega}>0.$
+Case 2: $q_{\min}\leqslant P\leqslant q_{\max}$. Here $|P-q_{\min}|=P-q_{\min}$ and $|P-q_{\max}|=q_{\max}-P$, so
+$\sup_{s}|P-q(s)|=\max(P-q_{\min},\;q_{\max}-P).$
+Subcase 2a: $q_{\max}-P\geqslant P-q_{\min}$. The supremum equals $q_{\max}-P=\|\phi_{\omega}\|_{L^{2}}^{2}-P$, and the same computation as in Case 1 yields
+$A-2\sup_{s}|P-q(s)|=\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2P\geqslant 4\sqrt{\omega}>0.$
+Subcase 2b: $P-q_{\min}>q_{\max}-P$. The supremum equals $P-q_{\min}$. Since $P\leqslant q_{\max}=\|\phi_{\omega}\|_{L^{2}}^{2}$, we have $2\|\phi_{\omega}\|_{L^{2}}^{2}-2P\geqslant 0$, and therefore
+$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P+2q_{\min}\\ &\geqslant\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}+2q_{\min}\geqslant 2q_{\min}>0.\end{split}$
+Combining both subcases, in Case 2:
+$A-2\sup_{s}|P-q(s)|\geqslant\min\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)>0.$
+Case 3: $P>q_{\max}$. Since $P>q(s)$ for all $s$, one has |Pq(s)|=Pq(s)|P-q(s)|=P-q(s) and the supremum is attained at the minimizer of $q$:
+$\sup_{s}|P-q(s)|=P-q_{\min}.$
+By the Cauchy–Schwarz inequality, $P\leqslant\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}\|\psi_{1,\omega}\|_{L^{2}}$, hence
+$\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P\geqslant\bigl(\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}-\|\psi_{1,\omega}\|_{L^{2}}\bigr)^{2}\geqslant 0.$
+Therefore:
+$\begin{split}A-2\sup_{s}|P-q(s)|&=2\|\phi_{\omega}\|_{L^{2}}^{2}+2q_{\min}+\bigl(\|\psi_{1,\omega+\varepsilon}\|_{L^{2}}^{2}+\|\psi_{1,\omega}\|_{L^{2}}^{2}-2P\bigr)\\ &\geqslant 2\|\phi_{\omega}\|_{L^{2}}^{2}+2q_{\min}\geqslant 2q_{\min}>0.\end{split}$
+Collecting all cases, we conclude that
+$A-2\sup_{s\in\mathbb{R}/2L\mathbb{Z}}|P-q(s)|\geqslant\min\!\bigl(4\sqrt{\omega},\;2q_{\min}\bigr)=:c_{0}^{2}>0,$
+where $c_{0}^{2}$ depends only on $\omega$, $L$, and the elliptic modulus $k=k(\omega)$, but not on $\varepsilon$.